### v1

# Evaluating LLMs at Evaluating Temporal Generalization

Abstract
The rapid advancement of Large Language Models (LLMs) highlights the urgent need for evolving evaluation methodologies that keep pace with improvements in language comprehension and information processing. However, traditional benchmarks, which are often static, fail to capture the continually changing information landscape, leading to a disparity between the perceived and actual effectiveness of LLMs in ever-changing real-world scenarios. Furthermore, these benchmarks do not adequately measure the models’ capabilities over a broader temporal range or their adaptability over time. We examine current LLMs in terms of temporal generalization and bias, revealing that various temporal biases emerge in both language likelihood and prognostic prediction. This serves as a caution for LLM practitioners to pay closer attention to mitigating temporal biases. Also, we propose an evaluation framework Freshbench  for dynamically generating benchmarks from the most recent real-world prognostication prediction. Our code is available at https://github.com/FreedomIntelligence/FreshBench. The dataset will be released soon.
1 Introduction
Proprietary large language models (LLMs) such as GPT-4 (OpenAI, 2023), Bard (GoogleAI, 2023), Claude (Anthropic, 2023), and PaLM (Anil et al., 2023) has been paralleled by efforts in the open-source community to democratize LLMs. Meanwhile, the rapid push for LLM democratization (Touvron et al., 2023) emphasizes the necessity for dynamic and robust benchmarks that accurately reflecting their evolving capabilities. Traditional benchmarks, broadly classified into knowledge-based assessments and open-dialogue evaluations, each present unique challenges. Knowledge-based assessments, exemplified by MMLU (Hendrycks et al., 2020) and C-Eval (Huang et al., 2023), are prone to data manipulation, raising concerns about their real-world applicability and the risk of data leakage. Open-dialogue evaluations like MT-Bench (Zheng et al., 2023) and Alpaca-Eval (Li et al., 2023), rely on subjective human or model-based judgments, making their outcomes susceptible to the influence of dialogue structure, and thereby, potentially compromising assessment validity.
We argue that the assessment of LLMs should be conducted in scenarios that are resistant to hacking manipulation, with results that are objective and free from evaluator biases (Chen et al., 2024). Additionally, time as a dynamic dimension, can also ensure robust and relevant evaluations. This paper explores an alternative way to evaluate large language models, i.e., on temporal generalization where time is considered as an important aspect. By doing so, the inaccessibility of future content would largely reduce the data contamination issue. Moreover, a retrospective perspective on time could beneficial for temporal bias analysis for LLMs.
To achieve this objective, we explore two specific scenarios designed to measure model capabilities without data leakage and to assess model performance on tasks relevant to both past and future contexts. The first scenario involved generating text from recent sources, such as arxiv papers, news articles, or new books, and Wikipedia. We expected the models to generalize to these diverse, newly generated texts with high language likelihood, thereby supporting the concept of ”intelligence as compression” in a manner free from data leakage. The second scenario focused on predicting future events, suggesting that a superior model should be able to understand the present context, incorporate world knowledge, and generalize to future scenarios. In summary, the evaluation at temporal generalization offers two main benefits: it eliminates the possibility of data leakage through temporal boundaries, and its results can be objectively assessed without reliance on subjective interpretations.
Our analysis indicates that using validation loss of new text to evaluate models is mostly effective. However, we observe that while performance on language likelihood is a strong indicator of a model’s linguistic capabilities, it does not always align with specific abilities measured by existing benchmarks, which may represent discrete knowledge or other capacities that are important but not directly related to language modeling.
Our contributions are as follows:
  • We thoroughly define and quantify temporal generalization and bias, establishing a basis for understanding and assessing the capabilities of LLMs over time.
  • Through targeted experimental research, we assess the performance of existing large language models (LLMs) across various timeframes. We evaluate how these models handle tasks from different periods, providing insights into their historical and ongoing performance stability and adaptability.
  • We proposed an evaluation framework Freshbench to test the capabilities of LLMs in predicting future events. This framework is dynamically updated to reflect the latest data, ensuring that our evaluations remain relevant and accurate in the face of rapidly evolving data environments.
2 Conceptualization of Temporal Generalization
In this section, we will conceptualize Temporal Generalization in Sec. 2.1 and further measures it in 2.2. From a retrospective perspective in the Temporal Generalization, we also conceptualize Temporal Biases in Sec. 2.3 and quantify them in 2.4.
2.1 Temporal Generalization
LLMs face the ongoing challenge of remaining applicability as new text evolves. Consequently, it is crucial for these models to not only understand and generate text based on past and present data but also to anticipate and adapt to future linguistic trends. This leads us to an important concept below.
Definition 2.1.
Temporal Generalization refers to the LLMs’ ability to accurately understand, predict, and generate text relevant to past, present, and future contexts.
This requires the models to integrate and apply knowledge from historical and current data to adapt to the evolution of language and emerging trends. Temporal generalization highlights the challenge of designing models that remain effective and accurate over time, without the need for frequent re-training as new data is introduced. LLMs must maintain a delicate balance between leveraging historical data to understand long-standing language patterns and adapting to new expressions, terminology, and topics as they emerge.
2.2 Assessing Temporal Generalization
Having defined temporal generalization, we now turn to a dual-case study to illustrate the dynamic capabilities and limitations of LLMs while adapting to time-evolving linguistic and predictive challenges.
2.2.1 Case I: Language Likelihood
Predicting the likelihood of a sequence of words is a foundational aspect of natural language processing involved in language modeling. Perplexity, a measure of how uncertain a language model is about its predictions, has been critically evaluated in various studies. Saunshi et al. (2021) demonstrates how perplexity can significantly enhance zero-shot inference capabilities, highlighting the importance of these models in understanding and generating language. Wettig et al. (2023) and Tay et al. (2022) have scrutinized the direct correlation between low perplexity during pre-training and improved fine-tuning outcomes, while Liang et al. (2023) discusses the potential issues of data contamination when relying solely on historical text data.
Language Likelihood
In the assessment of language models, metrics such as perplexity (PPL) and bits per character (BPC) offer insights into model performance. Perplexity measures the model’s ability to predict a sequence of words by calculating the exponential of the negative average log likelihood. However, we refine our focus to BPC for a more normalized comparison.
Unlike traditional BPC which is specific to character-level language models, our adapted Bits Per Character (BPC) metric offers a comparison between models by normalizing the log likelihood based on character count, irrespective of the tokenization method, allowing for a more equitable assessment of model performance across different languages and tokenization schemes, the BPC is calculated as follows:
$BPC(T)=\frac{-\sum_{i=1}^{N}\log p(w_{i}|w_{1},\ldots,w_{i-1})}{\text{len-utf-8}(T)}$ (1)
[…]
2.2.2 Case II: Prognostication Prediction
Predictive modeling in the realm of LLMs is increasingly becoming a focal point of research due to its implications for understanding and forecasting dynamic real-world events. Bonde (2022) employed GPT-3 in a few-shot setting to tackle binary questions from the Metaculus platform, demonstrating the capabilities of LLMs in deriving informed predictions. Further, Zou et al. (2022) enriched the discourse by introducing a comprehensive dataset that includes forecasting questions alongside a news corpus. This approach enhances model training by incorporating varied and relevant data contexts. Moreover, Halawi et al. (2024) showcased how integrating news content can significantly improve the accuracy of predictions in real-time scenarios.
Building on this foundation, our research seeks to explore the temporal generalization capabilities of LLMs—how well they adapt and remain accurate as new data emerges and contexts evolve. We aim to deepen the understanding of how these models integrate and apply learned knowledge from past data to predict future events effectively through the case below.
Prognostication Prediction
As an example of using LLMs to predict political outcomes, we gather new factual data about future events, such as "What will President Biden’s approval rating be as of 7 June 2024, according to FiveThirtyEight?". To ensure objectivity in our predictions, answers are typically formatted as ranges within a prediction market scenario. For instance, President Biden’s approval rating might be categorized into intervals such as lower than 34%, at least 34.0%, but less than 36.0%, at least 36.0%, but less than 38.0%, up to 46.0% or higher.
However, since it is impossible to collect actual future data, we adopt a retrospective study approach. We use data available after the event’s question was initially posed and before the prediction market’s closure, treating this as ”future” data for model evaluation. This evaluation is an ongoing, online process, allowing to evaluate the model’s predictive capability in an objective and structured manner. Given the assessment’s nature, we define the accuracy of predictions, denoted by $Acc$, as follows:
$Acc=\frac{N_{\text{correct}}}{N_{\text{total}}}$ (2)
where $N_{\text{correct}}$ represents the number of correct predictions and $N_{\text{total}}$ is the total number of predictions made. This metric evaluates the effectiveness of LLMs in predicting future events, serving as a straightforward measure of their performance in real-world scenarios.
This approach allows us to continuously evaluate the model with the latest data, ensuring that the assessment is reflective of the model’s current capabilities and the evolving nature of the information landscape. By evaluating the generalization capabilities of LLMs across those scenarios, we seek to better understand how these models incorporate and reflect perceived world knowledge when confronted with dynamic, real-world conditions.
2.3 Temporal Bias
In the context of LLMs, there exists two biases, namely Nostalgia Bias (see Definition 2.2) and Neophilia Bias (see Definition 2.3), which reflect two contrasting tendencies in how these models might predict or generate text about the past, present, or future. Understanding the two biases is crucial for interpreting the models’ outputs, especially when attempting to predict future text. A visual representation of these biases is provided in Figure 1, where the trends associated with nostalgia and neophilia biases are plotted over time.
Figure 1: Schematic figure of nostalgia bias and neophilia bias. Prior to the cutoff time, the performance of the model exhibiting nostalgia bias is skewed toward past dates, showing a preference for historical data. In contrast, a model with neophilia bias is skewed toward future dates, indicating a tendency to favor more recent information. However, after the cutoff time, the model’s understanding of future trends declines, leading to a decrease in performance.
Definition 2.2.
Nostalgia Bias Nostalgia Bias in LLMs refers to a tendency to over-rely on or give undue preference to historical data or past events in generating text. This means that when predicting text about the past or the present, the model might disproportionately reflect the perspectives, styles, or information prevalent in its training data, which stops at a certain point in the past. This bias can lead to outputs that are more reflective of past conditions and less attuned to recent developments or changes that have occurred since the model’s last training data update. In essence, Nostalgia Bias can make a model seem ‘stuck in the past’, limiting its ability to accurately represent the present or predict the future.
Definition 2.3.
Neophilia Bias Neophilia Bias, on the other hand, would imply a model’s overemphasis on novelty, future trends, or speculative ideas, potentially at the expense of well-established information from the past or present. This bias could manifest in a model’s predictions or text generation leaning towards overly optimistic or speculative futures without sufficient grounding in historical data or current realities. While less likely in practice due to the nature of LLM training on historical datasets, a perceived Neophilia Bias could arise from the way prompts are structured or the interpretation of the model’s outputs, especially if they seem to prioritize futuristic concepts over historical accuracy.
In a word, Nostalgia bias refers to the model’s tendency to reflect knowledge from earlier periods before the training cutoff, whereas neophilia bias indicates a model’s inclination towards the most recent information near the cutoff time.
Challenges in Temporal Generalization Predicting future text with LLMs while navigating Nostalgia Bias and Neophilia Bias presents distinct challenges. On one hand, Nostalgia Bias can cause models to generate predictions that might not fully account for recent trends or emerging phenomena, making the predictions potentially outdated or irrelevant. On the other hand, an overcorrection towards Neophilia Bias (or an attempt to counteract Nostalgia Bias) could lead to predictions that are too speculative or disconnected from past (especially long-term) trajectories. The challenge lies in balancing these biases to leverage the historical and contemporary data effectively, ensuring that predictions about the future are both informed by past patterns and open to new possibilities. Achieving this balance requires careful selection and adjustment of the model’s training data, and possibly the integration of up-to-date information alongside the model’s outputs.
2.4 Qualitative Measurement of Temporal Bias
We aim to assess temporal bias qualitatively by introducing the Temporal Bias Index (TBI).
Definition 2.4.
Temporal Bias Index (TBI) qualitatively determines whether the model’s performance is biased towards past or future dates. It could be calculated by analyzing the slope of the performance over time, which indicates changes in model performance. A negative slope suggests a decline in performance, indicating a bias towards historical data, whereas a positive slope indicates an improvement, suggesting a bias towards more recent data.
The methodology for calculating TBI varies based on the metric used to measure model performance. Specifically, when performance is measured by language likelihood, we calculate the slope of BPC over time directly: This slope indicates whether BPC is increasing (a positive slope, suggesting degradation) or decreasing (a negative slope, indicating improvement).
Conversely, when performance is measured by the correctness of answers to questions, we define several key symbols: $n$ represents the total number of attempts at questions within the month, $P$ is the observed frequency of correct answers, $E(P_{\text{random}})$ is the expected frequency of correct answers under random guessing calculated based on the average number of options, and $\Phi$ is the cumulative distribution function (CDF) of the standard normal distribution.
Using these symbols, the performance evaluation formula TBI is presented as follows:
$\text{TBI}=\Phi\left(\frac{\sqrt{n}\times(P-E(P_{\text{random}}))}{\sqrt{P\times(1-P)}}\right)$ (3)
Equation 3 normalizes the deviation between the observed frequency of correct answers and the expected frequency under random guessing, producing a performance score between 0 and 1. A higher score indicates better performance relative to the expected level of a standard normal distribution, effectively adjusting for the biases introduced by random guessing and the variability in the number of options.
Figure 2: A case study of temporal bias, assessing the accuracy of responses to queries from Good Judgment Open dataset (smoothed), spanning from September 1, 2015, to the models’ cutoff time. The blue line for ”Yi-6B” is right-skewed with a bias of -0.006, signifying a tendency to favor past dates, which infers a ’Nostalgia’ bias within the model. In contrast, the green line, which depicts the ”Phi-1.5” model, shows a minimal bias of 0.001, signifying a slight preference for more recent information. This can be interpreted as a ’Neophilia’ bias.
In our case study of temporal bias, as depicted in Figure 2, we demonstrate the validity of our defined measurement of temporal bias. The bias metric effectively gauges whether the model’s performance exhibits a tendency to favor either past or future dates.
3 Temporal Generalization in terms of Language Likelihood
Temporal generalization in language models evaluates their ability to maintain performance over time, even as data evolves. This section discusses how various models handle data from different periods, showcasing their performance stability or volatility.
3.1 Experimental Settings
Our system aggregates data from various platforms, including Yahoo for financial news, BBC for political insights, Reddit for discussions, Wattpad for literature, Wikipedia for encyclopedic updates, arXiv for academic research, GitHub for software trends, and Quora for diverse questions and answers. Utilizing a Python-based crawling framework, detailed in our public repository, it’s designed for adaptability and efficient data collection, with Playwright addressing navigation and specific site challenges. We also did some basic data pre-processing, see details in Appendix A. An overview of the data volume and the average entry length from a single crawl is provided in Appendix B.
Data sources of Fresh Corpora We selected a set of models due to their diversity and significance in advancing the capabilities of NLP systems. They are not only benchmarks of their respective sizes, ranging from 3 billion to 13 billion parameters, but also exemplify the rapid progression in NLP technology. The models originate from diverse sources, showcasing the collaborative and competitive spirit driving innovation in the industry.
To further delve into the model’s performance dynamics over varying text lengths, we truncate the input text to the first 5000 characters and compute the sum of loss. The loss values are then accumulated to derive their sum, providing insight into the model’s efficiency at different stages of text comprehension.
Additionally, we calculate the average loss over token length milestones—specifically at 10, 50, 100, 500, 1000, 1500, and 2000 tokens. This step allows us to observe the impact of length on the model’s understanding and predictive capabilities, highlighting how performance metrics vary with increasing input size.
3.2 Temporal Bias
The slope is obtained by performing a linear regression analysis where the independent variable $x$ is the month and the dependent variable $y$ is the model’s performance score, here is BPC, the lower the better, a negative slope indicates the presence of Neophilia Bias. Conversely, a positive slope would suggest that the model’s performance is decreasing, which shows Nostalgia Bias.
To delve deeper into the temporal generalization capabilities of models, we plot the BPC curve of fresh text over time. We sampled arXiv papers from 2015 to 2024 and applied smoothing techniques to the data because the number of samples was not sufficient to completely stabilize the average language likelihood. Nonetheless, a slight downward trend in model BPC can be observed, which is noteworthy. What is more, models with initially higher BPC showed a larger percentage decrease in BPC (as seen in models like OPT-2.7B and PHI-2 in Figure  4). For models with lower initial BPC, the decrease trend is less pronounced. Taking arXiv text as an example, nearly most models do not show clear flaws in absolute generalization, as BPC decreases over time. Moreover, in terms of arXiv data, LLMs exhibit a neophilia bias rather than a nostalgia bias, capturing the present more effectively than the past.
Turning to BBC news text, the observed trends in BPC from 2020 to 2023, depicted in Figure 4, provide insight into how well these models process and understand news content over a span of several years. As delineated in the figure, we note a pronounced increase in BPC across the observed period, which suggests a diminishing adaptability of the models to newer data, indicative of a nostalgia bias. This trend may stem from the dynamic and evolving nature of news narratives. Detailed data for arXiv, BBC, and Wiki datasets are listed in Table 1.
Figure 3: BPC on arXiv text from 2015.9 to 2024.1 (smoothed)
Figure 4: BPC on BBC news text from 2020 to 2023, with three-month intervals (smoothed)
Model Name Released Date TBI*1000 (arXiv) TBI*1000 (BBC) TBI*1000 (Wiki)
OPT-13B May 2022 -19.4 1689.0 -30.2
OPT-2.7B May 2022 -18.8 1080.7 -46.9
LLaMA-7B Feb 2023 -15.3 713.0 52.0
Pythia-12B Mar 2023 -8.1 802.7 -11.1
Falcon-rw-1B Apr 2023 -26.8 607.4 -47.7
Baichuan-13B-Chat Jun 2023 -11.7 685.7 4.8
Baichuan-7B-Chat Jun 2023 -14.6 726.6 -16.5
LLaMA-2-13B Jul 2023 -11.1 769.6 65.9
LLaMA-2-7B Jul 2023 -12.6 684.6 23.2
Baichuan-13B-Chat Jul 2023 -7.0 717.2 12.1
Zhongjing-Base Jul 2023 -15.0 759.7 56.9
InternLM-Chat-7B Jul 2023 -19.3 574.0 -22.0
Baichuan2-7B-Base Aug 2023 -9.5 687.5 3.7
Baichuan2-7B-Chat Aug 2023 -8.4 781.8 -11.6
Mistral-7B-v0.1 Sep 2023 -15.1 756.8 10.4
Phi-1.5 Sep 2023 -26.0 1212.9 -60.7
Baichuan2-13B-Base Sep 2023 -9.1 747.9 15.8
Baichuan2-13B-Chat Sep 2023 -7.0 717.2 12.1
Colossal-LLaMA-2-7B-Base Sep 2023 -21.2 696.3 -36.2
Qwen-14B-Chat Sep 2023 -7.9 762.2 51.5
Qwen-7B Sep 2023 -11.7 662.4 7.4
Qwen-7B-Chat Sep 2023 -13.2 705.7 -3.6
Skywork-13B-Base Oct 2023 -16.3 549.1 30.8
ChatGLM3-6B Oct 2023 -25.5 412.6 -70.4
Zephyr-7B-beta Oct 2023 -21.1 506.9 3.7
Yi-6B Nov 2023 -9.4 451.7 24.6
Yi-6B-Chat Nov 2023 -9.4 506.5 20.2
Qwen-1.8B Nov 2023 -22.7 374.0 -46.2
Qwen-1.8B-Chat Nov 2023 -25.3 426.8 -50.3
RWKV-v5-Eagle-7B Nov 2023 -12.8 511.3 4.2
TinyLLaMA-1.1B-Chat-v0.6 Dec 2023 -25.0 300.1 -45.3
Phi-2 Dec 2023 -21.2 669.3 -44.7
Table 1: Comprehensive list of models with their release dates and metrics for TBI (multiplied by 1000) across the arXiv, BBC, and Wiki datasets. Negative slope values indicate a trend of increasing BPC, which refers to decreasing performance.
3.3 Temporal Generalization
To quantify model temporal generalization, we computed the BPC for various models on the Wiki and BBC datasets. We established a base BPC using data from the six months prior to each model’s release. Subsequently, we measured the changes in BPC at 3, 6, 9 and 12 months post-release relative to the base BPC. The specific results are detailed in Appendix C
The generalization of BPC in Wikipedia dataset shows that the models exhibit a relatively stable performance on Wikipedia data, which suggests a consistent ability to handle well-structured and formal text types like those found in Wikipedia.
The generalization of BPC in BBC dataset shows that, as opposed to Wikipedia data, providing a more significant differentiation between models. This distinction is crucial as it reveals the models’ ability to generalize to future, unseen data. Notably, although there is an overall trend of increasing BPC for BBC text from 2020 to 2024, the immediate post-cutoff performance of individual models varies. For instance, Baichuan2 series appears to adapt better to new texts, showcasing an improved familiarity with them. On the other hand, models like Zephyr exhibit poorer generalization in this context.
In our analysis across both the Wiki and BBC datasets, we have observed significant correlations between model attributes and performance decay, detailed in Table 2. The data shows a consistent pattern where higher benchmark scores correlate with faster performance decay, indicating a trade-off between capability and flexibility. This suggests that models with higher capabilities and sophistication, as reflected in their task scores, may compromise adaptability, becoming so specialized that they struggle to adjust to new or evolving data scenarios without a noticeable decrease in performance.
Specifially, in the Wiki dataset, larger models, which generally have higher capabilities, tend to decay quicker (with a correlation of 0.504), likely due to overfitting on training data. Conversely, in the BBC dataset, where the information is more dynamic and rapidly changing, the correlation between model size and performance decay is almost negligible, suggesting that the impact of model size on decay is significantly less in environments that continuously present new challenges. Additionally, across both datasets, newer models show larger correlations, hinting that despite their advanced capabilities, they may decay faster, potentially as a consequence of these enhanced capabilities.
These insights emphasize the importance of designing models that not only perform well on benchmarks but also maintain robustness over time and across varying data environments. As machine learning applications continue to grow in complexity and scale, the ability to generalize well temporally will become increasingly critical. This analysis serves as a cornerstone for future research aimed at enhancing the longevity and adaptability of machine learning models in dynamic settings.
Task or model feature Correlation with BPC change rate average (Wiki) Correlation with BPC change rate average (BBC)
Model Size 0.504 -0.044
Release Date(month since 2015) 0.267 0.313
MMLU 0.599 0.384
MMLU:Humanities 0.500 0.317
MMLU:SocialSciences 0.594 0.311
MMLU:STEM 0.393 0.186
MMLU:Other 0.528 0.153
GSM8 0.473 0.188
Ar20c 0.380 -0.071
Hellawasg 0.581 0.248
TrustfulQA 0.200 0.485
Winogrande 0.632 0.204
Longbench 0.404 0.549
Table 2: Combined correlation between task scores and performance decay on Wikipedia and BBC
3.4 Correlation with Existing Benchmarks
The Selected Existing Benchmarks
We explored the correlation between BPC and existing benchmarks, particularly focusing on how specific content types, such as GSM, influence model performance. The selected benchmarks encompass common sense reasoning, depth and breadth of knowledge across diverse subjects, truthfulness, natural language understanding, and mathematical reasoning, ensuring a well-rounded assessment of the model’s performance.
Benchmark Description
HellaSwag Tests an AI’s common sense reasoning by requiring it to complete sentences or narratives that reflect everyday logic, highlighting the model’s ability to predict logical continuations. (Zellers et al., 2019)
MMLU Evaluates an AI’s breadth and depth of understanding across a diverse range of subjects, from science to literature, showcasing its versatility and comprehensive grasp of human language. (Hendrycks et al., 2021)
TruthfulQA Focuses on the model’s capacity to provide honest and accurate responses, particularly to questions where the truth may not be intuitive, testing its commitment to truthfulness and its proficiency in avoiding misinformation. (Lin et al., 2022)
Winogrande Assesses common sense reasoning and natural language understanding through sentence completion tasks, measuring an AI’s ability to apply common sense knowledge in language processing. (Sakaguchi et al., 2019)
GSM8K Challenges AI models with elementary-level math problems, testing their mathematical reasoning capabilities and their understanding and application of basic math concepts. (Cobbe et al., 2021)
Table 3: Overview of AI Benchmarks and Their Evaluation Focus
Base vs. Chat Models
Table 4 shows the comparative analysis of base and tuned models. Our findings suggest that: 1) Base models exhibit a stronger correlation with established benchmarks, implying that these models, in their fundamental form, possess a robust capacity for modeling language. This observation advocates for conducting initial evaluations using base models to capture their intrinsic language modeling capabilities. 2) The exploration of tuned models reveals potential deeper patterns in model tuning, suggesting that specific tuning strategies may unlock further enhancements in model performance across different tasks and domains.
Benchmarks Loss of Base Model Loss of Chat Model Delta
MMLU -0.61 -0.22 0.39
MMLU:Humanities -0.59 -0.42 0.17
MMLU:SocialSciences -0.54 0.42 0.96
MMLU:STEM -0.29 0.02 0.31
MMLU:Other -0.47 -0.1 0.37
Arc -0.46 0.34 0.8
Hellawasg -0.66 -0.57 0.09
Wnogrande -0.56 0.57 1.13
Longbench -0.04 -0.65 -0.61
GSM8K -0.4 0.22 0.62
TrustfulQA 0.05 0.39 0.34
Table 4: Correlation with Different Benchmarks, Base vs. Chat Models
Correlation alone Text Length
Table 5 is about a long text understanding benchmark Longbench.
For chat models in this task, the correlation between BPC and scores is logical—higher probabilities of generating the text correlate with higher scores, and this correlation strengthens with longer text lengths. This suggests a consistent alignment between the model’s language likelihood in long text and its scoring, reinforcing the chat model’s robustness in handling extensive text inputs.
However, an inverse trend is observed with base models, where a familiar text paradoxically results in lower scores on related comprehension tasks. This indicates that fine-tuning may alter the base models’ understanding of the texts, perhaps shifting their processing in ways that do not favor traditional comprehension metrics.
This indicates scenarios where a model scores high on language likelihood due to its text-generation skills but may not perform equally well on tasks that require deep semantic understanding or critical thinking. Thus, relying solely on language likelihood as an indicator of overall performance might overlook crucial aspects of cognitive and interpretative abilities that are essential for more complex applications.
Length 10 30 50 100 200 300 400 500 800 1000 1200 1500 2000
Base Models 0.35 0.13 0.06 -0.03 -0.05 -0.06 -0.05 -0.05 -0.04 -0.04 -0.04 -0.04 -0.04
Chat Models 0.01 -0.07 -0.16 -0.39 -0.53 -0.58 -0.61 -0.63 -0.65 -0.65 -0.67 -0.67 -0.68
Table 5: Correlation with Longbench values across different lengths
On the Genres of Corpora
As can be seen in Appendix  D, whether all models are observed or only those of a specific scale does not make a big difference. Comparisons across different model sizes show that the differences are not substantial, indicating consistency in performance correlations across benchmarks.
Further analysis from Table 6 within the MMLU benchmark sections reveals interesting patterns. While we assume models to be generally excel in domains closely aligned with their training data, this correlation is not strong. Specifically, STEM scores do not show strong correlation with STEM-related texts. Moreover, academic papers, even in Social Sciences and Humanities, exhibit stronger correlations than Internet data, which includes forums and news.
In this context, code-related data show similarities in performance correlation with academic texts, suggesting a closer alignment than previously anticipated. Similarly, Internet data, including forums and news, also shows comparative closeness in performance, indicating nuanced relationships across different types of content.
Benchmark Academia STEM Academia non-STEM Code Internet article Internet Chat
MMLU:STEM -0.17 -0.19 -0.11 -0.01 -0.01
MMLU:Humanities -0.54 -0.53 -0.51 -0.29 -0.39
MMLU:Social Sciences -0.49 -0.42 -0.46 -0.25 -0.31
Table 6: Correlation Coefficients between BPC in specific areas and Domain-Specific Performance
Additionally, compared to chat forum texts, more formal academic texts demonstrate a stronger correlation with benchmarks. The genre of text has a minor impact on the sign of the correlation coefficient, which is relatively small compared to the differences across model types.
3.5 The Drawbacks of Evaluating Language Likelihood
Inaccessibility to API-based Models
A fundamental limitation of using language likelihood as a metric lies in its inapplicability to closed-source models, where key indicators such as perplexity, logits, or losses are inaccessible. This restriction excludes a broad spectrum of proprietary models from evaluation and significantly narrows the scope of comparative analysis. The consequent inability to assess these models hampers our ability to gain a comprehensive understanding of the language technology landscape and the capabilities of different models within it.
Susceptibility to Training Focus
The effectiveness of language likelihood as a metric can be compromised by models that are finely tuned or trained with a particular emphasis on certain formats or domains. In such cases, the likelihood measure might disproportionately favor these specialized areas, skewing the overall evaluation and presenting a distorted view of a model’s general performance.
4 Temporal Generalization in terms of Prognostication Prediction
The above Drawbacks in Sec. 3.5 motivated us to research temporal generalization and bias from the perspective of future prediction. In the context of the model release date, events that occurred before this date are considered ”past events”, while those that took place after are termed ”future events”. In this chapter, we have divided the experiments into two categories based on this criterion: Bias and Temporal Generation.
4.1 Experimental Settings
4.1.1 Data Source
We have sourced all questions from  Good Judgment Open and all binary questions from  Metaculus through web scraping and API downloads. The collected data includes initiation and conclusion dates, domains, questions, descriptions, and answers, which are ordered appropriately for further processing.
4.1.2 Extraction of Predictive Responses
Using the lm-eval-harness  (Gao et al., 2023) tool, we predict model choices based on probability associated with each option. This probability-driven method serves as a reliable metric for evaluating the predictive performance of language models across a variety of topics and question formats.
4.2 Temporal Bias
As can be seen in Table 7, most analyzed models exhibit a negative slope. This is an uncommon scenario in a typical statistical analysis, indicating a general tendency across models to perform better or be more aligned with historical data than recent data. We define the following classification thresholds to discern subtle differences in how models handle temporal information, despite an overall inclination towards nostalgia.
  • Positive TBI (Neophilia Bias): TBI values greater than 0.001 suggest the model has a neophilia bias. This implies the model’s performance is better on more recent data, potentially favoring newer over older training data.
  • Near-Zero TBI (Balanced): TBI values between -0.001 and 0.001 indicate a balanced performance over time. A TBI of 0, in particular, suggests that the model’s performance is evenly distributed around the mean, showing no temporal bias.
  • Negative TBI (Nostalgia Bias): TBI values less than -0.001 indicate a nostalgia bias, suggesting the model may favor historical data, possibly overlooking important recent information.
The choice of ±0.001 as the threshold for these categories is designed to fine-tune the sensitivity of our analysis to subtle temporal biases, ensuring that the classification is both meaningful and practically significant.
model TBI y-axis intercept bias category
OPT-13B -1.37 0.70 Nostalgia
OPT-2.7B -2.85 0.97 Nostalgia
LLaMA-7B -3.71 0.96 Nostalgia
Pythia-12B -3.90 0.95 Nostalgia
Falcon-rw-1B -2.47 0.84 Nostalgia
Baichuan-13B-Chat -2.79 0.98 Nostalgia
Baichuan-7B-Chat -1.17 0.80 Nostalgia
GPT-3.5-turbo-0613 -0.81 0.86 Symmetrical
LLaMA-2-7B -1.66 0.75 Nostalgia
LLaMA-2-13B -0.80 0.73 Symmetrical
LLaMA-2-7B-Chat -3.92 0.93 Nostalgia
Zhongjing-Base 0.08 0.63 Symmetrical
InternLM-Chat-7B -1.76 0.94 Nostalgia
Baichuan2-7B-Base -3.50 0.84 Nostalgia
Baichuan2-7B-Chat -3.71 0.94 Nostalgia
Mistral-7B-v0.1 -0.47 0.75 Symmetrical
Phi-1.5 0.24 0.38 Symmetrical
Baichuan2-13B-Base -3.06 0.72 Nostalgia
Baichuan2-13B-Chat -1.53 0.70 Nostalgia
Colossal-LLaMA-2-7B-Base -3.38 0.91 Nostalgia
Qwen-14B-Chat -2.31 0.93 Nostalgia
Qwen-7B -0.35 0.69 Symmetrical
Qwen-7B-Chat -2.98 0.62 Nostalgia
Skywork-13B-Base -1.46 0.65 Nostalgia
ChatGLM3-6B -4.46 0.92 Nostalgia
Zephyr-7B-beta -2.73 0.99 Nostalgia
Yi-6B -5.18 0.98 Nostalgia
Yi-6B-Chat 0.34 0.71 Symmetrical
Qwen-1.8B -3.59 0.80 Nostalgia
Qwen-1.8B-Chat -4.32 0.96 Nostalgia
GPT-4-1106 -0.08 0.96 Symmetrical
TinyLLaMA-1.1B-Chat-v0.6 -1.78 0.53 Nostalgia
Gemini-pro-1213 -0.43 0.76 Symmetrical
Phi-2 -1.54 0.77 Nostalgia
Table 7: Predict performance of models: TBI(slope) and y-axis intercept are linear fit parameters, bias category classifies the nature of the bias based on TBI
Table 8 shows an analysis covering the period from 2015 to 2024, with the year 2019 serving as a pivotal midpoint, a distinct pattern emerges from the performance data of various language models: all models reached their peak accuracy before 2020. This observation points to a notable trend toward ”nostalgia,” where models demonstrate optimal performance on data from the earlier part of the timeline analyzed. This tendency highlights the models’ alignment with the informational characteristics prevalent prior to the latter half of the observed period, suggesting their training data may disproportionately reflect earlier times.
The peak accuracy periods analyzed are based on the average accuracy of all questions within three-month intervals, to provide a stable and robust assessment of each model’s performance.
Peak Period Model Peak Time of Accuracy Cutoff Time Interpolation
Post-2019 TinyLLaMA1.1B 2019-07 2023-11 52
2017 to 2018 GPT4-1106 2018-07 2023-06 59
LLaMA2-13B 2018-07 2023-02 55
Baichuan2-13B-Base 2018-01 2023-06 65
2015 to 2016 Baichuan2-7B-Chat 2016-01 2023-08 91
Colossal-LLaMA-2-7B-Base 2016-01 2023-09 92
LLaMA2-7B 2016-01 2023-02 85
Table 8: Analysis about the time of the best-achieved accuracy and its cutoff time.
4.3 Temporal Generalization
It is crucial to note that for models, any data post-release represents the future, allowing us to assess their generalization capabilities on upcoming events.
In the analysis of temporal generalization, it is important to acknowledge that the performance of models post-release primarily reflects their ability to predict future events based on past training. As evidenced by the data in Table 9, many models show significant fluctuations in accuracy percentages following their release. This variability suggests that immediately after the cutoff event, most models are essentially guessing, leading to high volatility in performance metrics.
Given the substantial fluctuations observed post-release, the temporal generalization of current models may appear to have limited significance. This is largely because these models exhibit weak capabilities in adapting to new, unseen data, often resulting in what can be described as near-random performance at the cutoff point. Such outcomes highlight the models’ limited understanding of the world and their inability to effectively generalize from past information to future contexts.
However, this should not detract from the potential utility of these measures in the future. As language models continue to evolve and become more sophisticated in their world understanding, the metric of temporal generalization will likely become more relevant and insightful. Future models that better comprehend and adapt to the world will provide more meaningful insights into their generalization capabilities over time, thereby enhancing the value of this assessment.
The current limitations underscore the need for further development in language modeling, particularly in improving the models’ ability to generalize across time. As we progress towards models with a deeper and more accurate understanding of complex world dynamics, temporal generalization measures will play a crucial role in evaluating their long-term effectiveness and adaptability.
Base adjusted ACC Adjusted ACC Change at 3 months (%) Adjusted ACC Change at 6 months (%) Adjusted ACC Change at 9 months (%) adjusted ACC Change at 12 months (%) Mean (%)
Model
OPT-2.7B 0.86 0.11 -13.63 -23.41 -18.98 -13.98
LLaMA-7B 0.66 -0.31 -20.59 5.09 9.89 -1.48
Pythia-12B 0.77 -27.68 -24.85 -6.22 -3.68 -15.61
Falcon-rw-1B 0.69 -34.51 -8.30 7.15 -4.52 -10.04
Baichuan-13B-Chat 0.64 0.69 12.79 8.00 0.13 5.40
GPT-3.5-turbo-0613 0.65 -29.81 16.71 26.91 -39.69 -6.47
LLaMA-2-13B 0.57 29.19 38.00 34.54 - 33.91
LLaMA-2-7B 0.54 25.57 41.09 27.09 - 31.25
LLaMA-2-7B-Chat 0.61 25.39 23.81 12.52 - 20.58
Zhongjing-Base 0.59 25.16 37.70 14.65 - 25.83
InternLM-Chat-7B 0.48 3.43 25.13 21.65 - 16.74
Baichuan2-7B-Base 0.60 14.99 32.31 4.81 - 17.37
Baichuan2-7B-Chat 0.60 16.12 22.67 -18.59 - 6.73
Mistral-7B-v0.1 0.61 30.25 34.87 -22.79 - 14.11
Phi-1.5 0.65 19.30 10.78 -17.57 - 4.17
Baichuan2-13B-Base 0.61 10.81 17.67 6.75 - 11.75
Colossal-LLaMA-2-7B-Base 0.58 15.99 3.33 - - 9.66
Qwen-14B-Chat 0.59 14.68 14.30 - - 14.49
Qwen-7B 0.62 7.45 -7.31 - - 0.07
Qwen-7B-Chat 0.48 60.26 74.85 - - 67.55
Skywork-13B-Base 0.61 20.20 5.04 - - 12.62
ChatGLM3-6B 0.60 3.24 6.29 - - 4.76
Zephyr-7B-beta 0.61 14.33 18.22 - - 16.28
Yi-6B 0.63 8.42 -9.47 - - -0.53
Yi-6B-Chat 0.67 13.11 -6.19 - - 3.46
Qwen-1.8B 0.68 -6.93 -33.98 - - -20.46
Qwen-1.8B-Chat 0.73 4.74 -8.91 - - -2.08
RWKV-v5-Eagle-7B 0.70 -0.09 -20.38 - - -10.23
GPT-4-1106 0.42 6.67 -7.11 - - -0.22
TinyLLaMA-1.1B-Chat-v0.6 0.71 13.27 -16.30 - - -1.51
Gemini-pro-1213 0.71 -15.15 -41.28 - - -28.21
Phi-2 0.81 7.11 - - - 7.11
Table 9: Percentage Change in Model Accuracy Following Release
4.4 Correlation with Existing Benchmarks
Following our previous chapter 3.4, we continue to explore the correlation between prediction performance and existing benchmarks, using the same benchmarks as in 3.4. We have categorized the tags of questions into several groups: Societal Issues Items, Political and Government Affairs, Economic and Business Items, and Science and Technology Items. Each group represents a significant area of interest.
In our analysis, we focus on how the correlation between these categories and various benchmarks evolves over time. We use correlation coefficients to measure the strength of association between the performance scores of questions within each category and the benchmarks across different times.
Notably, our analysis revealed that while the Economic and Business Items and Science and Technology Items on base models displayed relatively high initial correlations, there appears to be little to no relationship in Societal issues Items and Political and Government Affairs, we think such topics are influenced by a multitude of unpredictable factors such as social changes, policy adjustments, and public sentiment. These elements make accurate predictions particularly challenging, leading to the observed low correlation with performance metrics. Detailed figures are in Appendix H.
The graphical representations in our analysis plot the changes in correlation coefficients over time. Each line in the graph represents the changing relationship between a category’s average adjusted performance and a specific benchmark, highlighting periods of stronger or weaker association.
For a detailed breakdown of the categories and the number of entries in each, please refer to the Appendix G. This provides a clearer picture of the dataset’s composition and the scope of our analysis.
4.5 Community Call to Action
We urge the community to take note of the above phenomenon: models tend to have a better comprehension of earlier events. This observation supports the findings in  Cheng et al. (2024), which suggest that models are most familiar with knowledge up to 2020. This observation is contrary to our initial hypothesis that models may overfit current events, which is significantly more concerning. It suggests that caution is needed, especially for those hoping to use model judgments directly.
5 Conclusion
We rigorously define and qualify temporal generalization and bias within LLMs, uncovering significant challenges in their ability to predict and adapt to future contexts. The experiments reveal that LLMs often struggle with temporal biases, affecting their reliability in real-world applications, particularly in dynamic scenarios requiring accurate forecasting.
The evaluation framework we provide allows for an in-depth assessment of these models’ performance over time. The findings underscore the urgent need for improvements in LLM training and updating processes to enhance their adaptability and mitigate biases.
Moving forward, it is crucial to refine LLM development and evaluation techniques to better handle evolving informational environments, ensuring their continued relevance and utility in various applications.
Looking ahead, our agenda for future work includes several ambitious goals aimed at enhancing the versatility and depth of our evaluation framework. Key among these objectives are:
  • Integrating a wider array of models, particularly focusing on those based on the Transformer architecture, such as Griffin and Mamba models.
  • Expanding the number of checkpoints accessed for a more comprehensive temporal analysis of model performance.
  • Delving into the relationship between token-level loss metrics and overall scores, which could offer more granular insights into model capabilities.
  • Developing methodologies to assess the obsolescence of benchmarks and models, potentially offering a dynamic approach to evaluating the relevance and efficiency of language models over time.
These initiatives represent our commitment to advancing the state of the art in language model evaluation, ensuring that our tools remain at the forefront of technological progress and innovation.
Limitations
Despite the system’s innovative approach to evaluating LLMs using the latest textual data, it is important to acknowledge certain limitations. One significant limitation is the system’s applicability primarily to open LLMs that provide access to their logits. This dependency arises because our evaluation mechanism relies on analyzing these logits to compute loss metrics. Consequently, our system may not be directly applicable to non-open LLMs, where such detailed internal outputs are not accessible. This constraint highlights the need for adaptable evaluation strategies that can cater to a broader range of LLMs, including those with closed architectures.
Ethics Statement
We fully acknowledge the need to adhere to ethical standards in handling and generating textual data. Hence, we commit to:
  • Handling User-Uploaded Content: We respect the privacy and copyright of content uploaded by users. All data provided by users will be used solely for the purpose of model evaluation and not for any other purposes. Reasonable measures will be taken to protect this data from unauthorized access, use, or disclosure.
  • Responsibility for Content Generated by Open-Source Models: We recognize that content generated by our system or through the use of open-source models involved in the evaluation might reflect or amplify biases. While efforts are made to minimize these biases and improve the fairness of the content, we cannot take responsibility for the generated content. We encourage users and developers to critically engage with these models and remain vigilant of their outputs.
  • Addressing Data Leak Issues: Our system aims to tackle data leak issues by dynamically generating benchmarks and ensuring models generalize to new or future texts. We commit to continuously monitoring and updating our evaluation methods to avoid potential data leakage, ensuring the validity and reliability of our assessment results.
  • Transparency: We are committed to maintaining the highest level of transparency in our work, including our evaluation methods, data sources used, and assessment outcomes. We encourage open critique and suggestions for improvement of our methodology, aiming to foster knowledge sharing and technological advancement.
We believe that through this approach, we can offer a valuable resource to the research and development community, while fostering deep reflection and discussion on ethical issues in the artificial intelligence field.
Appendix A Details of Data Collection
The data for our system is sourced from a diverse set of online platforms, each selected for its relevance to different data types and domains. We utilize a Python-based crawling framework, which is detailed in our publicly available repository. This framework is designed to be adaptable, allowing for the inclusion of additional sources as they become relevant.
The primary data sources include:
  • Financial News: Crawled from Yahoo, providing up-to-date information on global financial trends.
  • Political News: BBC News is used to gather the latest political developments worldwide.
  • Discussion Forums: Reddit is scraped to capture current discussions across a variety of tofigures.
  • Online Literature: Wattpad offers a rich source of contemporary fiction and non-fiction.
  • Encyclopedia: Wikipedia’s latest changes feed is monitored for updates across all domains.
  • Academic Papers: arXiv is used to access the newest research across multiple scientific fields.
  • Code Repositories: GitHub Trending provides insights into the latest developments in software engineering.
  • Question and Answer Forums: Quora, focusing on popular topics such as Technology, Mathematics, Health, and Movies.
To facilitate efficient and effective data collection, our system’s environment is configured with a specific Python package index URL and utilizes Playwright for web navigation and content extraction. This setup addresses potential challenges, such as page refresh requirements on Yahoo and hostname verification issues encountered with certain websites.
Our system employs a combination of tools and libraries, including Requests, Playwright, and various PDF processing utilities, to collect the latest textual data from the internet. This approach enables us to dynamically update our benchmark datasets with fresh information, ensuring that our evaluation reflects current language use and information trends.
Data Pre-processing For some raw content clean the text by removing these tags to ensure that only the content text is analyzed.
Filtering out short texts: Extremely short texts may not provide enough contextual information for analysis or model training. Setting a minimum text length threshold of 100 chars to filter out such instances.
External Dataset For the analysis of text likelihood and bias, we employed not only our own collection of historical arXiv data but also sampled entries from two additional datasets. We incorporated data from the BBC News and Wikitext datasets, which are publicly available through the Hugging Face datasets repository. Specifically, the datasets used were bbc_news_alltime and wikitext_alltime, accessible via the following URLs:
Sampling from these datasets was carried out every three months, with each instance involving the collection of 50 entries per dataset. This structured sampling approach was designed to ensure a consistent and representative analysis of content over time, enabling an effective assessment of textual bias and likelihood trends.
Appendix B Data Overview of One Crawl
Table 10 presented below offers a comprehensive overview of the diverse sources and categories of texts that were analyzed in one data crawl. Classified into distinct groups such as Academic STEM, Academic Non-STEM, Internet QA, and Internet articles, the table delineates the number of texts extracted from each source along with their average length. This includes detailed counts from different academic archives like arXiv, various categories from Internet question and answer platforms like Reddit and Quora, and articles from well-known online platforms including Wiki and BBC.
Category Source Count Length
Academic STEM arXiv Mathematics 50 42850
arXiv Computer Science 50 43336
arXiv Physics 50 42580
arXiv Statistics 50 42619
arXiv Electrical Engineering and Systems Science 50 43027
Academic Non-STEM arXiv Economics 50 42592
arXiv Quantitative Finance 50 43969
arXiv Quantitative Biology 50 42974
Code GitHub 30 21849
Internet QA Reddit 583 643
Quora Health 20 1150
Quora Movies 69 1349
Quora Mathematics 70 1365
Quora Technology 80 1674
Internet article Wiki 220 4091
Wattpad 38 6615
BBC 45 4792
Yahoo 41 5373
Table 10: Overview of text categories with their sources counts, and average lengths in one sample crawl.
Appendix C Detailed results of BPC generalization on Wiki and BBC dataset
Model Base BPC BPC Change at 3 months (%) BPC Change at 6 months (%) BPC Change at 9 months (%) BPC Change at 12 months (%) mean
OPT-13B 0.450 2.273 -2.087 1.265 0.839 0.572
OPT-2.7B 0.475 1.970 -2.228 1.192 1.143 0.519
LLaMA-7B 0.450 1.270 0.177 -7.776 3.266 -0.766
Pythia-12B 0.480 -0.233 -3.888 -2.612 -1.012 -1.936
Falcon-rw-1B 0.510 0.459 -7.814 1.201 -1.474 -1.907
Baichuan-13B-Chat 0.467 -3.645 -1.078 0.413 - -1.436
LLaMA-2-13B 0.416 -6.928 3.502 1.683 - -0.581
LLaMA-2-7B 0.434 -7.314 3.092 1.279 - -0.981
LLaMA-2-7B-Chat 0.505 -8.308 2.695 0.664 - -1.650
Zhongjing-Base 0.465 -8.488 2.662 -1.290 - -2.372
InternLM-Chat-7B 0.515 -2.476 -1.344 0.236 - -1.195
Baichuan2-7B-Base 0.451 -3.059 0.595 -4.649 - -2.371
Baichuan2-7B-Chat 0.517 -3.571 -0.038 -5.315 - -2.975
Mistral-7B-v0.1 0.398 1.110 5.347 -0.661 - 1.932
Phi-1.5 0.602 -1.772 1.131 -1.837 - -0.826
Baichuan2-13B-Base 0.439 -2.943 0.882 -4.390 - -2.150
Baichuan2-13B-Chat 0.495 -3.541 -0.995 0.232 - -1.435
Colossal-LLaMA-2-7B-Base 0.618 2.825 3.715 - - 3.270
Qwen-14B-Chat 0.445 2.682 4.847 - - 3.764
Qwen-7B 0.441 1.892 3.689 - - 2.791
Qwen-7B-Chat 0.466 1.809 3.606 - - 2.708
Skywork-13B-Base 0.429 7.163 3.261 - - 5.212
ChatGLM3-6B 0.832 5.567 1.684 - - 3.626
Zephyr-7B-beta 0.435 7.340 3.928 - - 5.634
Yi-6B 0.427 6.478 3.144 - - 4.811
Yi-6B-Chat 0.446 3.030 -2.798 - - 0.116
Qwen-1.8B 0.520 1.899 -2.082 - - -0.091
Qwen-1.8B-Chat 0.599 1.635 -2.033 - - -0.199
RWKV-v5-Eagle-7B 0.440 4.956 2.278 - - 3.617
TinyLLaMA-1.1B-Chat-v0.6 0.513 2.539 -3.678 - - -0.569
Phi-2 0.487 1.882 - - - 1.882
Table 11: Percentage Change in BPC Following Release, measured by BBC news data from 2020 to 2024.3. Dash (”-”) symbols in the table indicate missing data points due to the later release dates of some models, which means there is no available data for those specific time intervals.
Model Base BPC BPC Change at 3 months (%) BPC Change at 6 months (%) BPC Change at 9 months (%) BPC Change at 12 months (%) mean
OPT-13B 0.075 -0.623 0.105 -0.413 -0.818 -0.437
OPT-2.7B 0.082 -0.658 0.024 -0.528 -0.915 -0.519
LLaMA-7B 0.052 0.214 -0.425 0.379 -0.067 0.025
Pythia-12B 0.071 -0.578 -0.933 -0.859 -0.903 -0.818
Falcon-rw-1B 0.083 -1.272 -1.113 -1.131 -1.207 -1.181
Baichuan-13B-Chat 0.058 0.019 0.124 -0.125 - 0.006
LLaMA-2-13B 0.041 1.700 1.446 1.796 - 1.647
LLaMA-2-7B 0.050 0.774 0.349 0.517 - 0.547
LLaMA-2-7B-Chat 0.060 0.132 -0.186 -0.093 - -0.049
Zhongjing-Base 0.050 0.756 0.269 0.569 - 0.531
InternLM-Chat-7B 0.080 -0.482 -0.301 -0.250 - -0.344
Baichuan2-7B-Base 0.059 0.523 0.157 0.185 - 0.289
Baichuan2-7B-Chat 0.068 0.343 -0.035 -0.073 - 0.078
Mistral-7B-v0.1 0.053 0.381 -0.028 0.018 - 0.124
Phi-1.5 0.098 0.066 -0.125 -0.248 - -0.102
Baichuan2-13B-Base 0.053 1.019 0.797 0.900 - 0.905
Baichuan2-13B-Chat 0.062 0.299 0.628 0.512 - 0.480
Colossal-LLaMA-2-7B-Base 0.065 -0.042 -0.199 - - -0.121
Qwen-14B-Chat 0.049 1.301 1.939 - - 1.620
Qwen-7B 0.061 0.428 0.430 - - 0.429
Qwen-7B-Chat 0.065 0.302 0.226 - - 0.264
Skywork-13B-Base 0.050 0.075 0.276 - - 0.175
ChatGLM3-6B 0.105 -0.187 -0.245 - - -0.216
Zephyr-7B-beta 0.059 -0.049 -0.033 - - -0.041
Yi-6B 0.058 -0.071 0.022 - - -0.024
Yi-6B-Chat 0.060 -0.028 0.023 - - -0.003
Qwen-1.8B 0.080 -0.131 -0.251 - - -0.191
Qwen-1.8B-Chat 0.091 -0.331 -0.496 - - -0.414
RWKV-v5-Eagle-7B 0.065 0.149 0.220 - - 0.184
TinyLLaMA-1.1B-Chat-v0.6 0.070 -0.332 -0.473 - - -0.403
Phi-2 0.072 -0.037 - - - -0.037
Table 12: Percentage Change in Language Likelihood Following Release, measured by wiki data from 2020 to 2024.3. Dash (”-”) symbols in the table indicate missing data points due to the later release dates of some models, which means there is no available data for those specific time intervals.
Appendix D Analysis of Language Likelihood Values and Benchmark Correlation Across Text and Model Groups
(a) STEM: base models
(b) non-STEM: base models
(c) Non-STEM: chat models
Figure 5: The heatmaps display the correlations by loss for a group of models across different text chunk lengths and different benchmark scores. Each row represents a specific truncation length, while the columns represent different benchmark score categories. The values in the cells indicate the corresponding correlation coefficients. (a) presents the correlations by loss for all 18 base models on STEM Academia texts. (b) presents the correlations by loss for all 18 base models on non-STEM Academia texts. (c) presents the correlations by loss for all 15 chat models on non-STEM Academia texts.
(a) STEM: all chat models
(b) STEM: 7B level chat models
Figure 6: (a) presents the correlations by loss for all 15 chat models on STEM Academia texts. (b) presents the correlations by loss for all 9 7B level chat models on STEM Academia texts. We can conclude that the scale of models being observed does not significantly alter outcomes, suggesting that performance correlations across benchmarks are generally consistent regardless of model size.
Appendix E Function Fitting Strategies
We considered several fitting strategies, each with its unique mathematical approach to model decay:
Linear Decay
Linear decay is represented by a straight line, suggesting a constant rate of decay over time. The model is given by the equation $y=ax+b$.
Exponential Decay
Exponential decay is modeled by an exponential function, expressed as $y=y_{0}\cdot e^{-kt}$, where $y_{0}$ is the initial value, $t$ is time or another decaying factor, and $k$ is the decay constant that denotes the rate of decay.
Power Law Decay
Power law decay implies a polynomial relationship and is expressed as $y=at^{-k}$, where $a$ is a positive constant, $t$ is time or another related factor, and $k$ represents the exponent characterizing the rate of decay.
Logarithmic Decay
Logarithmic decay suggests a decrease that becomes more gradual over time, typically modeled as $y=a+b\log(x)$, where $a$ and $b$ are constants, and $x$ represents time or another related factor.
The variance of the residuals for each model was calculated as follows:
Model Variance
Linear 2.346e-3
Exponential 1.71e3
Power Law 3.78e-3
Logarithmic 3.33e-3
Table 13: Variance of Residuals for Different Models
E.0.1 Findings
Our analysis found that the linear model provided the best fit, with the lowest variance among the residuals. This indicates a steady, unchanging rate of knowledge decay over time, as opposed to the more complex models that assume varying rates.
We aimed to evaluate the understanding and recency of knowledge across various language models, including GPT-4 and others. To accomplish this, we devised a methodology to calculate a weighted accuracy score for each model, reflecting not only its performance but also the contemporaneity of its knowledge base.
We began by assigning weights to different quarters, starting from 0 for the first quarter of 2015, up to 1 for the most recent quarter. This linear interpolation of weights was chosen to emphasize the value of more current information, underlining our hypothesis that accuracy concerning more recent knowledge is indicative of a model’s modernity and relevance.
The formula applied was straightforward: for each model, we multiplied its accuracy (Acc) by the corresponding weight based on the time difference since the first quarter of 2015. The resulting weighted accuracies were then summed up to yield an overall score for each model, which we used to rank them accordingly.
GPT-4 and other leading models demonstrated a notably later peak in knowledge, suggesting that their databases include more recent information. This is particularly valuable in fast-evolving fields where the timeliness of information is crucial.
Below is a partial presentation of rankings and scores for selected models, showcasing their weighted accuracy scores. The complete table can be found in Table 14.
Model Name Score
GPT4-1106 794.275
GPT4-0613 699.8
Zephyr-7B-beta 601.35
InternLM-Chat-7B 592.15
GPT3.5-turbo-0613 578.025
Baichuan-13B-Chat 571.8
Qwen-14B-Chat 565.45
OPT-2.7B 547.375
Yi-6B-Chat 541.05
Mistral-7B-v0.1 528.725
Baichuan-7B-Chat 527.825
LLaMA-7B 524.075
Baichuan2-7B-Chat 517.725
Pythia-12B 513.375
Colossal-LLaMA2-7B-Base 513.35
LLaMA2-7B-Chat 508.4
LLaMA2-13B 508.15
Qwen-7B 506.7
Qwen-1.8B-Chat 505.225
Phi-2 497.4
Yi-6B 496.15
Zhongjing-Base 490.1
LLaMA2-7B 486.2
Falcon-rw-1B 481.15
ChatGLM3-6B 479.8
Baichuan2-7B-Base 479.175
Baichuan2-13B-Chat 463.75
OPT-13B 453.075
Baichuan2-13B-Base 446.225
Skywork-13B-Base 445.8
Qwen-1 8B 435.775
RWKV5-Eagle-7B 408.425
Qwen-7B-Chat 394.9
Phi-1.5 385.975
TinyLLaMA-1.1B-Chat-v0.6 366.125
Table 14: Model Scores
Appendix F Details of Model Accuracy Peaks
Table15 captures the peak performance periods of various models on Freshbench. It lists specific models along with their respective dates when they achieved the best performance.
Date Models
2015-01 Skywork-13B-Base, Phi-1.5
2015-07 Skywork-13B-Base, Phi-1.5
2016-01 Baichuan2-7B-Chat, Colossal-LLaMA-2-7B-Base, LLaMA-2-7B, Qwen-14B-Chat, Qwen-1.8B, Qwen-1.8B-Chat, Yi-6B, Yi-6B-Chat, Baichuan-13B-chat, Baichuan-7B-chat, ChatGLM3-6B, falcon-1B, InternLM-chat-7B, OPT-13B, Phi-2
2016-07 RWKV5-Eagle-7B
2017-01 GPT-3.5-turbo-0613
2017-07 Baichuan2-7B-Base
2018-01 Baichuan2-13B-Base
2018-07 Baichuan2-13B-Chat, LLaMA-2-13B-hf, Qwen-7B, GPT4-0613, GPT4-1106, LLaMA2-7B-Chat, LLaMA-7B, mistral-7B-v0.1, OPT-2.7B, Pythia-12B, Zephyr-7B-beta
2019-07 TinyLLaMA-1.1B-Chat
2020-01
2020-07
2021-01
2021-07
2022-01
2022-07
2023-01
2023-07
2024-01
2024-07
Table 15: Detailed Record of Model Accuracy Peaks
Appendix G Question Tag Details
Group Category Number of Entries
Political and Government Affairs Leader Entry/Exit 199
US Politics 338
Elections and Referenda 265
US Policy 223
Foreign Policy 242
Non-US Politics 415
Security and Conflict 413
Total: 2095
Economic and Business Items Business 575
Economic Indicators 253
Finance 284
Economic Policy 147
Total: 1259
Societal issues Items Society 379
Health 237
Environment 75
Sports 120
Entertainment 132
Total: 943
Science and Technology Items Technology 346
Total: 346
Table 16: Detailed breakdown and sum of entries per category within each group of benchmarks.
Appendix H Prediction Correlation over Time
Figure 7: Science and Technology Items - All Model Spearman Smoothed
Figure 8: Science and Technology Items - Base Model Spearman Smoothed
Figure 9: Societal Issues Items - All Model Spearman Smoothed
Figure 10: Political and Government Affairs - All Model Spearman Smoothed
Figure 11: Economic and Business Items - All Model Spearman Smoothed

### v2

@@
-Evaluating LLMs at Evaluating Temporal Generalization
+Is Your LLM Outdated? Evaluating LLMs at Temporal Generalization
 Abstract
@@
 However, traditional benchmarks, which are often static, fail to capture the continually changing information landscape, leading to a disparity between the perceived and actual effectiveness of LLMs in ever-changing real-world scenarios.
-Furthermore, these benchmarks do not adequately measure the models’ capabilities over a broader temporal range or their adaptability over time.
-We examine current LLMs in terms of temporal generalization and bias, revealing that various temporal biases emerge in both language likelihood and prognostic prediction.
-This serves as a caution for LLM practitioners to pay closer attention to mitigating temporal biases.
-Also, we propose an evaluation framework Freshbench  for dynamically generating benchmarks from the most recent real-world prognostication prediction.
-Our code is available at https://github.com/FreedomIntelligence/FreshBench.
-The dataset will be released soon.
+Our study examines temporal generalization, which includes the ability to understand, predict, and generate text relevant to past, present, and future contexts, revealing significant temporal biases in LLMs.
+We propose an evaluation framework, for dynamically generating benchmarks from recent real-world predictions.
+Experiments demonstrate that LLMs struggle with temporal generalization, showing performance decline over time.
+These findings highlight the necessity for improved training and updating processes to enhance adaptability and reduce biases.
+Our code, dataset and benchmark are available at https://github.com/FreedomIntelligence/FreshBench.
 1 Introduction
-Proprietary large language models (LLMs) such as GPT-4 (OpenAI, 2023), Bard (GoogleAI, 2023), Claude (Anthropic, 2023), and PaLM (Anil et al., 2023) has been paralleled by efforts in the open-source community to democratize LLMs.
-Meanwhile, the rapid push for LLM democratization (Touvron et al., 2023) emphasizes the necessity for dynamic and robust benchmarks that accurately reflecting their evolving capabilities.
+Proprietary large language models (LLMs) such as GPT-4 [1], Bard [2], Claude [3], and PaLM [4] has been paralleled by efforts in the open-source community to democratize LLMs.
+Meanwhile, the rapid push for LLM democratization [5] emphasizes the necessity for dynamic and robust benchmarks that accurately reflect their evolving capabilities.
 Traditional benchmarks, broadly classified into knowledge-based assessments and open-dialogue evaluations, each present unique challenges.
-Knowledge-based assessments, exemplified by MMLU (Hendrycks et al., 2020) and C-Eval (Huang et al., 2023), are prone to data manipulation, raising concerns about their real-world applicability and the risk of data leakage.
-Open-dialogue evaluations like MT-Bench (Zheng et al., 2023) and Alpaca-Eval (Li et al., 2023), rely on subjective human or model-based judgments, making their outcomes susceptible to the influence of dialogue structure, and thereby, potentially compromising assessment validity.
-We argue that the assessment of LLMs should be conducted in scenarios that are resistant to hacking manipulation, with results that are objective and free from evaluator biases (Chen et al., 2024).
+Knowledge-based assessments, exemplified by MMLU [6] and C-Eval [7], are prone to data manipulation, raising concerns about their real-world applicability and the risk of data leakage.
+Open-dialogue evaluations like MT-Bench [8] and Alpaca-Eval [9], rely on subjective human or model-based judgments, making their outcomes susceptible to the influence of dialogue structure, and thereby, potentially compromising assessment validity.
+Given the limitations of traditional methods, researchers are exploring new ways to measure model performance more comprehensively.
+Recently, several methods have been proposed to evaluate models at different times and in various contexts[10, 11, 12].
+TimeQA[13] evaluates a model’s ability to handle temporal evolution, while Situated QA[14] focuses on a model’s performance in specific contexts.
+StreamingQA[15] introduces the concept of continuous information flow, while RealTimeQA[11] and FreshQA[12] emphasize real-time and up-to-date information processing capabilities.
+However, existing methods may be prone to data leakage during their evaluation processes and typically focus on current or past data, lacking exploration of future event prediction capabilities.
+We believe that the evaluation of LLMs should be conducted in scenarios that are resistant to hacking manipulation, with results that are objective and free from evaluator biases [16].
 Additionally, time as a dynamic dimension, can also ensure robust and relevant evaluations.
@@
 By doing so, the inaccessibility of future content would largely reduce the data contamination issue.
-Moreover, a retrospective perspective on time could beneficial for temporal bias analysis for LLMs.
+Moreover, a retrospective perspective on time could be eneficial for temporal bias analysis for LLMs.
 To achieve this objective, we explore two specific scenarios designed to measure model capabilities without data leakage and to assess model performance on tasks relevant to both past and future contexts.
 The first scenario involved generating text from recent sources, such as arxiv papers, news articles, or new books, and Wikipedia.
-We expected the models to generalize to these diverse, newly generated texts with high language likelihood, thereby supporting the concept of ”intelligence as compression” in a manner free from data leakage.
+We expected the models to generalize to these diverse, newly generated texts with high language likelihood, thereby supporting the concept of "intelligence as compression" in a manner free from data leakage.
 The second scenario focused on predicting future events, suggesting that a superior model should be able to understand the present context, incorporate world knowledge, and generalize to future scenarios.
@@
 However, we observe that while performance on language likelihood is a strong indicator of a model’s linguistic capabilities, it does not always align with specific abilities measured by existing benchmarks, which may represent discrete knowledge or other capacities that are important but not directly related to language modeling.
-Our contributions are as follows:
-  • We thoroughly define and quantify temporal generalization and bias, establishing a basis for understanding and assessing the capabilities of LLMs over time.
-  • Through targeted experimental research, we assess the performance of existing large language models (LLMs) across various timeframes.
-We evaluate how these models handle tasks from different periods, providing insights into their historical and ongoing performance stability and adaptability.
-  • We proposed an evaluation framework Freshbench to test the capabilities of LLMs in predicting future events.
-This framework is dynamically updated to reflect the latest data, ensuring that our evaluations remain relevant and accurate in the face of rapidly evolving data environments.
+Our contributions are as follows: 1) We thoroughly define and quantify temporal generalization and bias, establishing a basis for understanding and assessing the capabilities of LLMs over time.
+2) We proposed an benchmark Freshbench to test the capabilities of LLMs in predicting future events.
+This benchmark is dynamically updated to reflect the latest data, ensuring that our evaluations remain relevant and accurate in the face of rapidly evolving data environments.
+It experimentally assesses performance of existing LLMs across various timeframes, providing insights into their historical and ongoing performance stability and adaptability.
 2 Conceptualization of Temporal Generalization
+Figure 1: Schematic figure of temporal generalization.
 In this section, we will conceptualize Temporal Generalization in Sec.
-2.1 and further measures it in 2.2.
+2.1 and further measure it in 2.2.
 From a retrospective perspective in the Temporal Generalization, we also conceptualize Temporal Biases in Sec.
-2.3 and quantify them in 2.4.
+2.3.
+In general, prior to the cutoff time, a LLM might exhibit a Nostalgia bias that is skewed toward past dates (i.e., aligning better with historical data) or a Neophilia bias that is skewed toward future dates ( favoring more recent information).
+However, after the cutoff time, the model’s understanding of future trends declines, leading to temporal degeneration.
 2.1 Temporal Generalization
 LLMs face the ongoing challenge of remaining applicability as new text evolves.
-Consequently, it is crucial for these models to not only understand and generate text based on past and present data but also to anticipate and adapt to future linguistic trends.
-This leads us to an important concept below.
+Consequently, it is crucial for these models to not only understand and generate text based on seen (e.g., past) data but also to anticipate and adapt to future data.
+This inspires temporal generalization as below.
 Definition 2.1.
@@
 2.2 Assessing Temporal Generalization
-Having defined temporal generalization, we now turn to a dual-case study to illustrate the dynamic capabilities and limitations of LLMs while adapting to time-evolving linguistic and predictive challenges.
-2.2.1 Case I: Language Likelihood
-Predicting the likelihood of a sequence of words is a foundational aspect of natural language processing involved in language modeling.
-Perplexity, a measure of how uncertain a language model is about its predictions, has been critically evaluated in various studies.
-Saunshi et al.
-(2021) demonstrates how perplexity can significantly enhance zero-shot inference capabilities, highlighting the importance of these models in understanding and generating language.
-Wettig et al.
-(2023) and Tay et al.
-(2022) have scrutinized the direct correlation between low perplexity during pre-training and improved fine-tuning outcomes, while Liang et al.
-(2023) discusses the potential issues of data contamination when relying solely on historical text data.
-Language Likelihood
-In the assessment of language models, metrics such as perplexity (PPL) and bits per character (BPC) offer insights into model performance.
-Perplexity measures the model’s ability to predict a sequence of words by calculating the exponential of the negative average log likelihood.
-However, we refine our focus to BPC for a more normalized comparison.
+Based on definition in Sec. 2.1, we turn to assess temporal generalization through the lens of 1) I: compression intelligence in Sec. 2.2.1 or II): prognostication prediction in Sec. 2.2.2.
+2.2.1 Case I: Compression Intelligence over Time
+In the assessment of compression intelligence, metrics such as perplexity (PPL) and bits per character (BPC) offer insights into model performance.
+We refine our focus to BPC for a more normalized comparison.
 Unlike traditional BPC which is specific to character-level language models, our adapted Bits Per Character (BPC) metric offers a comparison between models by normalizing the log likelihood based on character count, irrespective of the tokenization method, allowing for a more equitable assessment of model performance across different languages and tokenization schemes, the BPC is calculated as follows:
-$BPC(T)=\frac{-\sum_{i=1}^{N}\log p(w_{i}|w_{1},\ldots,w_{i-1})}{\text{len-utf-8}(T)}$ (1)
+$\small BPC(T)=\frac{-\sum_{i=1}^{N}\log p(w_{i}|w_{1},\ldots,w_{i-1})}{\text{len-utf-8}(T)}$ (1)
 […]
 2.2.2 Case II: Prognostication Prediction
-Predictive modeling in the realm of LLMs is increasingly becoming a focal point of research due to its implications for understanding and forecasting dynamic real-world events.
-Bonde (2022) employed GPT-3 in a few-shot setting to tackle binary questions from the Metaculus platform, demonstrating the capabilities of LLMs in deriving informed predictions.
-Further, Zou et al.
-(2022) enriched the discourse by introducing a comprehensive dataset that includes forecasting questions alongside a news corpus.
+LLMs could performs prognostication prediction, as seen in [18] which employed GPT-3 in a few-shot setting to tackle binary questions from the Metaculus platform.
+Further, [19] enriched the discourse by introducing a comprehensive dataset that includes forecasting questions alongside a news corpus.
 This approach enhances model training by incorporating varied and relevant data contexts.
-Moreover, Halawi et al.
-(2024) showcased how integrating news content can significantly improve the accuracy of predictions in real-time scenarios.
-Building on this foundation, our research seeks to explore the temporal generalization capabilities of LLMs—how well they adapt and remain accurate as new data emerges and contexts evolve.
-We aim to deepen the understanding of how these models integrate and apply learned knowledge from past data to predict future events effectively through the case below.
-Prognostication Prediction
+Moreover, [20] showcased how integrating news content can significantly improve the accuracy of predictions in real-time scenarios.
+Building on this foundation, our research seeks to explore the temporal generalization capabilities of LLMs: how well they adapt and remain accurate as new data emerges and contexts evolve?.
 As an example of using LLMs to predict political outcomes, we gather new factual data about future events, such as "What will President Biden’s approval rating be as of 7 June 2024, according to FiveThirtyEight?".
@@
 For instance, President Biden’s approval rating might be categorized into intervals such as lower than 34%, at least 34.0%, but less than 36.0%, at least 36.0%, but less than 38.0%, up to 46.0% or higher.
-However, since it is impossible to collect actual future data, we adopt a retrospective study approach.
-We use data available after the event’s question was initially posed and before the prediction market’s closure, treating this as ”future” data for model evaluation.
-This evaluation is an ongoing, online process, allowing to evaluate the model’s predictive capability in an objective and structured manner.
-Given the assessment’s nature, we define the accuracy of predictions, denoted by $Acc$, as follows:
-$Acc=\frac{N_{\text{correct}}}{N_{\text{total}}}$ (2)
-where $N_{\text{correct}}$ represents the number of correct predictions and $N_{\text{total}}$ is the total number of predictions made.
-This metric evaluates the effectiveness of LLMs in predicting future events, serving as a straightforward measure of their performance in real-world scenarios.
-This approach allows us to continuously evaluate the model with the latest data, ensuring that the assessment is reflective of the model’s current capabilities and the evolving nature of the information landscape.
-By evaluating the generalization capabilities of LLMs across those scenarios, we seek to better understand how these models incorporate and reflect perceived world knowledge when confronted with dynamic, real-world conditions.
-2.3 Temporal Bias
-In the context of LLMs, there exists two biases, namely Nostalgia Bias (see Definition 2.2) and Neophilia Bias (see Definition 2.3), which reflect two contrasting tendencies in how these models might predict or generate text about the past, present, or future.
-Understanding the two biases is crucial for interpreting the models’ outputs, especially when attempting to predict future text.
-A visual representation of these biases is provided in Figure 1, where the trends associated with nostalgia and neophilia biases are plotted over time.
-Figure 1: Schematic figure of nostalgia bias and neophilia bias.
-Prior to the cutoff time, the performance of the model exhibiting nostalgia bias is skewed toward past dates, showing a preference for historical data.
-In contrast, a model with neophilia bias is skewed toward future dates, indicating a tendency to favor more recent information.
-However, after the cutoff time, the model’s understanding of future trends declines, leading to a decrease in performance.
+However, since collecting actual future data is impossible, we adopt a retrospective study approach.
+We use data available after the event’s question was posed and before the prediction market’s closure, treating this as "future" data for model evaluation.
+This ongoing, online evaluation objectively assesses the model’s predictive capability.
+We define prediction accuracy, denoted by $Acc$, as follows:
+$\small Acc=\frac{N_{\text{correct}}}{N_{\text{total}}}$ (2)
+where $N_{\text{correct}}$ is the number of correct predictions among $N_{\text{total}}$ future prediction questions.
+See.
+Sec.
+4.1 for the collected future prediction question set over time.
+This metric evaluates the effectiveness of LLMs in predicting future events.
+This approach also allows continuous evaluation with the latest data, ensuring the assessment reflects the model’s current capabilities and the evolving information landscape.
+2.3 Temporal Bias in a Retrospective View
+Temporal biases are multifaceted in previous literature.
+For instance, temporal bias can refer to the assumption of an incorrect sequence of events that misleads our reasoning about causality [21] or user preferences over time in point-of-interest recommendations [22].
+Recently, biases have raised more challenges in pre-trained language models [23], particularly focusing on under-represented societies.
+[24] examines lookahead bias in pre-trained language models, providing insights into how these biases affect model performance.
+In the context of LLMs, temporal bias refers to the preferences of LLMs over time.
+There exists two biases, namely Nostalgia Bias (see Definition 2.2) and Neophilia Bias (see Definition 2.3), which reflect two contrasting tendencies in how these models might predict or generate text about the past, present, or future.
+A visual representation of these biases is provided in Fig.
+1, where the trends associated with Nostalgia and Neophilia biases are plotted over time.
 Definition 2.2.
-Nostalgia Bias Nostalgia Bias in LLMs refers to a tendency to over-rely on or give undue preference to historical data or past events in generating text.
-This means that when predicting text about the past or the present, the model might disproportionately reflect the perspectives, styles, or information prevalent in its training data, which stops at a certain point in the past.
-This bias can lead to outputs that are more reflective of past conditions and less attuned to recent developments or changes that have occurred since the model’s last training data update.
-In essence, Nostalgia Bias can make a model seem ‘stuck in the past’, limiting its ability to accurately represent the present or predict the future.
+Nostalgia Bias refers to the over-reliance on historical data when generating text.
+This bias causes models to reflect perspectives, styles, or information prevalent in past training data, leading to outputs that may not account for recent developments.
+Essentially, Nostalgia Bias can make a model seem ‘stuck in the past’, limiting its ability to accurately represent the present or predict the future.
 Definition 2.3.
-Neophilia Bias Neophilia Bias, on the other hand, would imply a model’s overemphasis on novelty, future trends, or speculative ideas, potentially at the expense of well-established information from the past or present.
-This bias could manifest in a model’s predictions or text generation leaning towards overly optimistic or speculative futures without sufficient grounding in historical data or current realities.
-While less likely in practice due to the nature of LLM training on historical datasets, a perceived Neophilia Bias could arise from the way prompts are structured or the interpretation of the model’s outputs, especially if they seem to prioritize futuristic concepts over historical accuracy.
-In a word, Nostalgia bias refers to the model’s tendency to reflect knowledge from earlier periods before the training cutoff, whereas neophilia bias indicates a model’s inclination towards the most recent information near the cutoff time.
+Neophilia Bias implies a model’s overemphasis on novelty and future trends, potentially at the expense of established information.
+This bias could result in predictions that are overly speculative and not well-grounded in historical data or current realities.
+Though less common due to LLM training on historical datasets, Neophilia Bias could be perceived from prompt structuring or interpretation of outputs that prioritize futuristic concepts over historical accuracy.
 Challenges in Temporal Generalization Predicting future text with LLMs while navigating Nostalgia Bias and Neophilia Bias presents distinct challenges.
-On one hand, Nostalgia Bias can cause models to generate predictions that might not fully account for recent trends or emerging phenomena, making the predictions potentially outdated or irrelevant.
-On the other hand, an overcorrection towards Neophilia Bias (or an attempt to counteract Nostalgia Bias) could lead to predictions that are too speculative or disconnected from past (especially long-term) trajectories.
-The challenge lies in balancing these biases to leverage the historical and contemporary data effectively, ensuring that predictions about the future are both informed by past patterns and open to new possibilities.
-Achieving this balance requires careful selection and adjustment of the model’s training data, and possibly the integration of up-to-date information alongside the model’s outputs.
-2.4 Qualitative Measurement of Temporal Bias
-We aim to assess temporal bias qualitatively by introducing the Temporal Bias Index (TBI).
-Definition 2.4.
-Temporal Bias Index (TBI) qualitatively determines whether the model’s performance is biased towards past or future dates.
-It could be calculated by analyzing the slope of the performance over time, which indicates changes in model performance.
-A negative slope suggests a decline in performance, indicating a bias towards historical data, whereas a positive slope indicates an improvement, suggesting a bias towards more recent data.
-The methodology for calculating TBI varies based on the metric used to measure model performance.
-Specifically, when performance is measured by language likelihood, we calculate the slope of BPC over time directly: This slope indicates whether BPC is increasing (a positive slope, suggesting degradation) or decreasing (a negative slope, indicating improvement).
-Conversely, when performance is measured by the correctness of answers to questions, we define several key symbols: $n$ represents the total number of attempts at questions within the month, $P$ is the observed frequency of correct answers, $E(P_{\text{random}})$ is the expected frequency of correct answers under random guessing calculated based on the average number of options, and $\Phi$ is the cumulative distribution function (CDF) of the standard normal distribution.
-Using these symbols, the performance evaluation formula TBI is presented as follows:
-$\text{TBI}=\Phi\left(\frac{\sqrt{n}\times(P-E(P_{\text{random}}))}{\sqrt{P\times(1-P)}}\right)$ (3)
-Equation 3 normalizes the deviation between the observed frequency of correct answers and the expected frequency under random guessing, producing a performance score between 0 and 1.
-A higher score indicates better performance relative to the expected level of a standard normal distribution, effectively adjusting for the biases introduced by random guessing and the variability in the number of options.
-Figure 2: A case study of temporal bias, assessing the accuracy of responses to queries from Good Judgment Open dataset (smoothed), spanning from September 1, 2015, to the models’ cutoff time.
-The blue line for ”Yi-6B” is right-skewed with a bias of -0.006, signifying a tendency to favor past dates, which infers a ’Nostalgia’ bias within the model.
-In contrast, the green line, which depicts the ”Phi-1.5” model, shows a minimal bias of 0.001, signifying a slight preference for more recent information.
-This can be interpreted as a ’Neophilia’ bias.
-In our case study of temporal bias, as depicted in Figure 2, we demonstrate the validity of our defined measurement of temporal bias.
-The bias metric effectively gauges whether the model’s performance exhibits a tendency to favor either past or future dates.
-3 Temporal Generalization in terms of Language Likelihood
+Nostalgia Bias can lead to outdated predictions, while overcorrection towards Neophilia Bias might result in speculative predictions disconnected from past trajectories.
+Balancing these biases is crucial to leverage historical and contemporary data effectively, ensuring future predictions are informed by past patterns and open to new possibilities.
+This requires careful adjustment of the model’s training data and integration of up-to-date information.
+3 Temporal Generalization through the Lens of Compression Intelligence
 Temporal generalization in language models evaluates their ability to maintain performance over time, even as data evolves.
 This section discusses how various models handle data from different periods, showcasing their performance stability or volatility.
-3.1 Experimental Settings
+3.1 Temporal Bias
+3.1.1 Protocol for Quantitative Measurement
 Our system aggregates data from various platforms, including Yahoo for financial news, BBC for political insights, Reddit for discussions, Wattpad for literature, Wikipedia for encyclopedic updates, arXiv for academic research, GitHub for software trends, and Quora for diverse questions and answers.
 Utilizing a Python-based crawling framework, detailed in our public repository, it’s designed for adaptability and efficient data collection, with Playwright addressing navigation and specific site challenges.
-We also did some basic data pre-processing, see details in Appendix A.
-An overview of the data volume and the average entry length from a single crawl is provided in Appendix B.
+We also did some basic data pre-processing, due to space limitation, details in Appendix B.
+An overview of the data volume and the average entry length from a single crawl is provided in Appendix C.
 Data sources of Fresh Corpora We selected a set of models due to their diversity and significance in advancing the capabilities of NLP systems.
@@
 The models originate from diverse sources, showcasing the collaborative and competitive spirit driving innovation in the industry.
-To further delve into the model’s performance dynamics over varying text lengths, we truncate the input text to the first 5000 characters and compute the sum of loss.
+Experimental Settings To further delve into the model’s performance dynamics over varying text lengths, we truncate the input text to the first 5000 characters and compute the sum of loss.
 The loss values are then accumulated to derive their sum, providing insight into the model’s efficiency at different stages of text comprehension.
-Additionally, we calculate the average loss over token length milestones—specifically at 10, 50, 100, 500, 1000, 1500, and 2000 tokens.
-This step allows us to observe the impact of length on the model’s understanding and predictive capabilities, highlighting how performance metrics vary with increasing input size.
-3.2 Temporal Bias
-The slope is obtained by performing a linear regression analysis where the independent variable $x$ is the month and the dependent variable $y$ is the model’s performance score, here is BPC, the lower the better, a negative slope indicates the presence of Neophilia Bias.
+Throughout this experiment, approximately 300 GPU hours on NVIDIA A800 were consumed.
+3.1.2 Observation of Temporal Bias in Compression Intelligence
+To comprehensively explore the temporal generalization capabilities of models, we plot the BPC curve of fresh text over time.
+In text data before release, a negative slope indicates the presence of Neophilia Bias.
 Conversely, a positive slope would suggest that the model’s performance is decreasing, which shows Nostalgia Bias.
-To delve deeper into the temporal generalization capabilities of models, we plot the BPC curve of fresh text over time.
 We sampled arXiv papers from 2015 to 2024 and applied smoothing techniques to the data because the number of samples was not sufficient to completely stabilize the average language likelihood.
 Nonetheless, a slight downward trend in model BPC can be observed, which is noteworthy.
-What is more, models with initially higher BPC showed a larger percentage decrease in BPC (as seen in models like OPT-2.7B and PHI-2 in Figure  4).
+What is more, models with initially higher BPC showed a larger percentage decrease in BPC (as seen in models like OPT-2.7B and PHI-2 in Figure  3).
 For models with lower initial BPC, the decrease trend is less pronounced.
-Taking arXiv text as an example, nearly most models do not show clear flaws in absolute generalization, as BPC decreases over time.
-Moreover, in terms of arXiv data, LLMs exhibit a neophilia bias rather than a nostalgia bias, capturing the present more effectively than the past.
-Turning to BBC news text, the observed trends in BPC from 2020 to 2023, depicted in Figure 4, provide insight into how well these models process and understand news content over a span of several years.
-As delineated in the figure, we note a pronounced increase in BPC across the observed period, which suggests a diminishing adaptability of the models to newer data, indicative of a nostalgia bias.
+Taking arXiv text as an example, BPC decreases over time, LLMs exhibit a Neophilia bias rather than a Nostalgia bias, capturing the present more effectively than the past.
+Turning to BBC news text, the observed trends in BPC from 2020 to 2023, depicted in Fig.
+3, provide insight into how well these models process and understand news content over a span of several years.
+We note a pronounced increase in BPC across the observed period, which suggests a diminishing adaptability of the models to newer data, indicative of a Nostalgia bias.
 This trend may stem from the dynamic and evolving nature of news narratives.
-Detailed data for arXiv, BBC, and Wiki datasets are listed in Table 1.
+Detailed result for various sources are in E
 Figure 2: BPC on arXiv text from 2015.9 to 2024.1 (smoothed)
 Figure 3: BPC on BBC news text from 2020 to 2023, with three-month intervals (smoothed)
+3.2 Temporal Generalization
+3.2.1 Protocol for Quantitative Measurement
+Definition 3.1.
+Temporal Bias Index (TBI) For each model in a certain period, we record $BPC_{i}$ at $n$ time points $t_{i}$, forming $n$ data points $(t_{i},p_{i})$, where $i=1,\ldots,n$. We then fit a linear model:
+$\small BPC_{i}=a*t_{i}+b$ (3)
+where Temporal Bias Index (TBI) is defined as the coefficient $a$. TBI is crucial for understanding the direction and rate of change in model performance over time.
+To quantify model temporal generalization, we computed the BPC for various models on the Wiki and BBC datasets and evaluated their performance using this index.
+Specifically, We established a base BPC using data from the six months prior to each model’s release.
+After that, we measured the changes in BPC at 3, 6, 9 and 12 months post-release relative to the base BPC; and assessed the overall trend in model performance using TBI.
+The specific results are detailed in Appendix D.
+Table 1: Correlation between Model Attributes and Their Performance Decay Across Various Tasks on Wikipedia and BBC Benchmarks
+Corre Model Release MMLU Human- Social- STEM MMLU GSM8 Ar20c Hella- Trust- Wino- Long-
+lation Size Date ities Sciences Other wasg fulQA grande bench
+Wiki 0.504 0.267 0.599 0.500 0.594 0.393 0.528 0.473 0.380 0.581 0.200 0.632 0.404
+BBC -0.044 0.313 0.384 0.317 0.311 0.186 0.153 0.188 -0.071 0.248 0.485 0.204 0.549
+3.2.2 Observation
+Models perform stably on Wikipedia but show more variation on BBC.
+In the Wikipedia dataset, models exhibit stable performance, demonstrating their ability to handle well-structured and formal text.
+However, larger models in the Wikipedia dataset deteriorate faster (correlation coefficient of 0.504), likely due to overfitting.
+In contrast, the BBC dataset shows significant variations in model performance, indicating insufficient adaptability to dynamic and informal text.
+From 2020 to 2024, the overall BPC for BBC texts shows an increasing trend, indicating gradual performance deterioration.
+Individual models show considerable differences in performance on new data; for instance, the Yi, LLaMA2, and Baichuan2 series adapt relatively well to new texts, while Phi-1.5 and OPT series exhibit poorer generalization (see Appendix E).
+These findings indicate that in dynamic environments, different models have significantly varying adaptation capabilities.
+Positive correlation between model attributes and performance decay.
+(see Tab.
+1).
+There is a positive correlation between model attributes (such as model size and release date) and performance decay.
+A positive BPC slope indicates an increase in BPC over time, reflecting performance deterioration.
+Larger model sizes may lead to rapid performance decay, suggesting a trade-off between capability and flexibility.
+In dynamic environments (such as the BBC dataset), the impact of model size on performance decay is smaller.
+Additionally, newer models show higher correlations, meaning that despite their advanced capabilities, they may deteriorate faster.
+These results emphasize the importance of designing models that can maintain robustness over time and across various data environments.
+As machine learning applications grow in complexity and scale, temporal generalization will become increasingly critical.
+4 Temporal Generalization through the Lens of Prognostication Prediction
+Advantages of Prognostication Prediction over Compression Intelligence Compression Intelligence is limited by their inapplicability to closed-source models, where key indicators like logits or losses are inaccessible.
+This exclusion narrows comparative analysis and hinders a comprehensive understanding of the language technology landscape.
+Additionally, language likelihood can be biased by models finely tuned to specific formats or domains, skewing overall evaluation and presenting a distorted view of a model’s general performance.
+This motivated us to research temporal generalization and bias from the perspective of future prediction.
+In this section, we investigate tempera generalization (in Sec.
+4.1) and biases (Sec.
+4.2) through the Lens of Prognostication Prediction.
+4.1 Temporal Biases
+4.1.1 Protocol for Quantitative Measurement: a Hypothesis Testing Method
+To distinguish the past and the future, we take the border as the release date of the model.
+For example, if a LLM is released at Jan.
+2024 (also denoted as ‘cutoff time’ in this work), the time after Jan.
+2024 would be the future.
+We could collect question prediction tasks in different periods.
+Based on different periods, We compute the prediction accuracy in these different periods, denoted as $acc_{\text{past}}$ and $acc_{\text{future}}$, where accuracy of the present (somewhere between the past and future) the is denoted as $acc_{\text{present}}$.
+To investigate the two mentioned temporal biases in Sec.
+2.3, we propose two hypothesis testings as follows.
+These tests help practitioners make informed decisions despite the randomness inherent in the question datasets.
+Hypothesis Testing 1.
+Neophilia Bias:
+$\small T_{1}:\ H_{0}:\mathbf{E}[acc_{\text{past}}]=\mathbf{E}[acc_{\text{present}}]\ \text{versus}\ H_{1}:\mathbf{E}[acc_{\text{past}}]<\mathbf{E}[acc_{\text{present}}],\$ (4)
+This hypothesis tests if there is a significant increase in model accuracy for current questions compared to past questions, suggesting a bias towards newer content.
+The p-value $\widehat{p}$ quantifies the strength of evidence against the null hypothesis.
+Hypothesis Testing 2.
+Nostalgia Bias:
+$\small T_{2}:\ H_{0}^{\prime}:\mathbf{E}[acc_{\textrm{past}}]=\mathbf{E}[acc_{\text{present}}]\ \text{versus}\ H_{1}^{\prime}:\mathbf{E}[acc_{\text{past}}]>\mathbf{E}[acc_{\text{present}}],$ (5)
+Conversely, this hypothesis tests if there is a significant increase in model accuracy for past questions, indicating a preference for older or historical content.
+The p-value $\widehat{p}^{\prime}$ quantifies the strength of evidence against the null hypothesis.
+P-Value Computation
+The p-values for assessing the shifts in accuracy are computed using the standard normal cumulative distribution function $\Phi$ as follows:
+$\small\widehat{p}=1-\Phi\left(\frac{acc_{\text{1}}-acc_{\text{2}}}{\sqrt{\frac{acc_{\text{1}}(1-acc_{\text{1}})}{n_{1}}+\frac{acc_{\text{2}}(1-acc_{\text{2}})}{n_{2}}}}\right),$ (6)
+$acc_{\text{1}}$ and $acc_{\text{1}}$ could be either $acc_{\textrm{past}}$ or $acc_{\textrm{present}}$, which is based on which Hypothesis Testing is used.
+Note that the two Hypothesis Testings are symmetrical.
+Experimental setting We have sourced all questions from GoodJudgmentOpen [25] through web scraping and API downloads.
+The collected data includes initiation and conclusion dates, domains, questions, descriptions, and answers, which are ordered appropriately for further processing.
+Throughout this experiment, approximately 50 GPU hours on NVIDIA A800 were consumed.
+We collected 2532 questions from GoodJudgmentOpen  [25], earliest one was proposed on 2015-09-01 and the most recent one was closed on 2024-05-25.
+It is "Will the adopted EU Artificial Intelligence (AI) Act be published in the Official Journal of the European Union before 25 May 2024?" Most common tags are in Appendix F.
+Table 2: Comparison of Model Performance in Different Pre-Release Periods (Compared to 0-20 months before Release), uses asterisks to denote significance levels: (p < 0.05), ††(p < 0.01), and ††† (p < 0.001), $\boldsymbol{\boldsymbol{\boldsymbol{\boldsymbol{\rightarrow}}}}$ refers to Neophilia Bias, $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$ refers to Nostalgia Bias.
+Models 20-40 Months Before Release 40-60 Months Before Release 60-80 Months Before Release
+OPT-13B -- $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$†† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$†
+OPT-2.7B -- $\boldsymbol{\boldsymbol{\boldsymbol{\boldsymbol{\rightarrow}}}}$  --
+LLaMA-7B -- $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$†† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$†
+Pythia-12B $\boldsymbol{\boldsymbol{\boldsymbol{\boldsymbol{\rightarrow}}}}$† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$†† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$††
+Falcon-rw-1B -- -- --
+Baichuan-7B-Chat -- -- $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$${}^{\dagger}$
+LLaMA-2-13B -- $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$†† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$††
+LLaMA-2-7B $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$†† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$  --
+LLaMA-2-7B-Chat -- $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$†† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$††
+Zhongjing-Base -- -- $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$††
+InternLM-Chat-7B $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$† -- --
+Baichuan2-7B-Chat $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$† -- --
+Mistral-7B-v0.1 $\boldsymbol{\boldsymbol{\boldsymbol{\boldsymbol{\boldsymbol{\rightarrow}}}}}$  $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$††
+Phi-1.5 $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$  $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$††
+Baichuan2-13B-Base $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$†† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$††
+Baichuan2-13B-Chat $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$  $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$†† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$††
+Colossal-LLaMA-2-7B-Base $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$† -- $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$
+Qwen-14B-Chat $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$††
+Qwen-7B -- -- --
+Skywork-13B-Base -- -- --
+Zephyr-7B-beta $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$  $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$†† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$††
+Yi-6B -- -- --
+Yi-6B-Chat $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$  $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$  $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$
+Qwen-1.8B -- -- $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$†
+RWKV-v5-Eagle-7B -- -- --
+Phi-2 $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$† -- $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$
+Command R+ $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$†† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$††
+Mixtral-8x22B-Instruct-v0.1 -- $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$†† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$††
+Phi-3-mini-4k-instruct -- $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$†† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$†
+Qwen1.5-110B-Chat $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$  $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$†† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$††
+4.1.2 Observation
+Neophilia Bias is prevalent closer to release, whereas Nostalgia Bias occurs earlier.
+As shown in Tab.
+2, among the biases, Nostalgia is the most common, followed by periods where neither bias is significant (shown as -), with Neophilia Bias being the least frequent.
+Phi-1.5 and Baichuan2-13B series show Nostalgia Bias in all three periods before release.
+Pythia-12B and Mistral-7B-v0.1 show Neophilia Bias in 20 to 40 months before release but show more significant Nostalgia in another two periods, which we attribute to random fluctuations rather than a consistent trend.
+OPT-2.7B is Neophilia overall, but this observation is likely due to its consistently near-random performance across various periods, rather than a true preference for novelty.
+Models tend to be stronger in earlier periods before release, with performance gradually declining over time.
+Using the -20 to 0 months period as an anchor, if we assign the three periods to represent the near past (20 to 40 months before release), the middle past (40 to 60 months before release), and the distant past (60 to 80 months before release), we observe that the further back in time we go, the more pronounced the Nostalgia Bias becomes, and it is also more significant.
+This suggests that the models gradually degrade during the pre-release period, rather than experiencing a rapid decline around the cutoff period.
+This observation reveals that models generally favor older knowledge and prefer information from more distant past periods.
+The distribution of knowledge is not uniform, nor does it tend to be closer to the present.
+This poses a potential risk in our utilization of these models because there is an expectation that the knowledge utilized should be as up-to-date as possible, ideally close to the model’s release date.
+4.2 Temporal Generalization
+4.2.1 Protocol
+In the prospective view when especially using LLMs for prognostication prediction, there might be a temporal degeneration issue defined as below:
+Definition 4.1.
+Temporal Degeneration The capability of prognostication prediction might be degraded over time due to the larger generalization gap for longer time.
+To qualitatively measure Temporal Degeneration, we propose a new hypothesis testing.
+Hypothesis Testing 3.
+Temporal Degeneration
+$\small T_{3}:$ $\ H_{0}:\mathbf{E}[acc_{\text{future}}]=\mathbf{E}[acc_{\text{present}}]\ \text{versus}\ H_{1}:\mathbf{E}[acc_{\text{future}}]<\mathbf{E}[acc_{\text{present}}],$ (7)
+Similar to Sec 4.1.1, $acc_{\text{future}}$ represents the model accuracy on questions proposed after the model’s release.
+This hypothesis assesses whether there is a significant decline in model accuracy, indicating a degeneration over time.
+The hypothetical p-value is similar to that of Eq.
+6.
+Table 3: Accuracy performance of various models across different periods.
+The presence of asterisks (*) denotes statistically significant decreases in performance during specific periods (p < 0.05 for *, p < 0.01 for ††, and p < 0.001 for ††*), benchmarked against their accuracy before release.
+Grey cells indicate the models have not been released in this period and represent the average scores of the models prior to their release.
+Model 23/1 -23/2 23/3 -23/4 23/5 -23/6 23/7 -23/8 23/9 -23/10 23/11 -23/12 24/1 -24/2 24/3 -24/4
+OPT-13B 0.55 0.48 0.31** 0.39 0.41 0.52 0.37* 0.27**
+OPT-2.7B 0.17 0.20 0.17* 0.22 0.25 0.22 0.26 0.26
+LLaMA-7B 0.21* 0.17** 0.16** 0.22* 0.25 0.22** 0.22* 0.24*
+Pythia-12B 0.48 0.47 0.32** 0.43 0.48 0.39* 0.43 0.31**
+Falcon-rw-1B 0.22 0.16 0.14* 0.28 0.27 0.21 0.20 0.24
+Baichuan-7B-Chat 0.32 0.24* 0.30 0.23 0.19** 0.20* 0.34
+InternLM-Chat-7B 0.35 0.23** 0.33 0.27 0.29 0.30 0.31
+GPT-3.5-turbo-230613 0.49 0.29** 0.37* 0.45 0.44 0.37* 0.30**
+GPT-4-230613 0.6 0.39** 0.43* 0.41** 0.37** 0.48 0.40**
+LLaMA-2-13B 0.52 0.43 0.48 0.36** 0.39* 0.40*
+LLaMA-2-7B 0.27 0.26 0.27 0.22 0.26 0.34
+LLaMA-2-7B-Chat 0.45 0.30* 0.46 0.35** 0.37 0.38
+Baichuan2-7B-Chat 0.37 0.35 0.43 0.31 0.37 0.41
+Baichuan2-13B-Base 0.49 0.30** 0.46 0.42 0.28** 0.40*
+Baichuan2-13B-Chat 0.6 0.33** 0.43 0.41* 0.26** 0.27**
+Zhongjing-Base 0.38 0.30 0.33 0.26* 0.39
+Mistral-7B-v0.1 0.44 0.41 0.34* 0.28* 0.34*
+Phi-1.5 0.41 0.38 0.39 0.30 0.25**
+Colossal-LLaMA-2-7B-Base 0.42 0.43 0.37 0.28* 0.33*
+Qwen-14B-Chat 0.39 0.25** 0.23** 0.33 0.27**
+Qwen-7B 0.36 0.36 0.30 0.33 0.25**
+Skywork-13B-Base 0.37 0.39 0.29* 0.30 0.27*
+ChatGLM3-6B 0.42 0.38 0.33* 0.26** 0.34
+Zephyr-7B-beta 0.4 0.32 0.31* 0.28* 0.28*
+Yi-6B 0.29 0.27 0.16** 0.22 0.34
+Yi-6B-Chat 0.38 0.25** 0.39 0.35
+Qwen-1.8B 0.35 0.22** 0.37 0.30
+RWKV-v5-Eagle-7B 0.29 0.29 0.33 0.28
+GPT-4-231106 0.66 0.44** 0.39** 0.36**
+Phi-2 0.31 0.23* 0.22 0.30
+Command R+ 0.54 0.38**
+Mixtral-8x22B-Instruct-v0.1 0.56 0.40**
+Phi-3-mini-4k-instruct 0.38 0.26**
+Qwen1.5-110B-Chat 0.56 0.40**
+4.2.2 Observation
+Earlier models perform worse.
+As the latest scores in Table 3 shows, models released earlier perform worse.
+The poor performance may be due to two factors.
+One is the natural decay over time, as evidenced by the clear downturn in LLaMA-7B’s performance.
+Over time, earlier models fail to adapt to new data.
+The other is the initial lack of advanced training techniques, which may have resulted in models like OPT showing near-random performance from the outset, without significant changes.
+This is mainly because earlier models were trained with less advanced techniques, leading to poorer performance.
+Most models exhibit weak temporal generalization ability.
+Experimental results show that many models have significant fluctuations in accuracy percentages after release, indicating weak temporal generalization ability.
+Models with strong predictive capabilities fail to maintain their scores, even if only for a short period.
+For instance, GPT-4-230613 has demonstrated a significant decline in four out of six periods after release, while GPT-4-231106 has seen a significant decrease across all three observed periods.
+This variability and instability suggest that models have insufficient adaptive capabilities when dealing with new, unseen data.
+Some models exhibit near-random performance after the cutoff event, highlighting their limited understanding of the world and inability to generalize effectively from past information to future contexts.
+This shows that current models’ temporal generalization ability is significantly lacking and needs further improvement.
+More advanced models perform better in future predictions.
+.
+For instance, Qwen1.5-110B-Chat, Command R+(104B), Mixtral-8x22B-Instruct-v0.1, along with GPT-4, have achieved the highest scores during the period from March to April 2024.
+It is worth noting that certain versions of LLaMA-2 and Baichuan2 have also performed commendably.
+This observation may stem from a bias, as models like Qwen1.5-110B-Chat , Command R+(104B), Mixtral-8x22B-Instruct-v0.1 are the most recently released model, and a more robust evaluation awaits further data updates.
+The advanced training techniques employed by these models likely contribute to their enhanced accuracy and stability.
+As time advances and an increasing amount of data is accumulated, the models’ assessments are expected to become more precise and valuable.
+5 Conclusion
+We rigorously define and qualify temporal generalization and bias within LLMs, uncovering significant challenges in predicting and adapting to future contexts.
+Our experiments reveal that LLMs often struggle with temporal biases, affecting reliability in dynamic scenarios requiring accurate forecasting.
+The evaluation framework and datasets we introduce allow for an in-depth assessment of model performance over time.
+The findings highlight shortcomings in existing benchmarks related to temporal shifts and underscore the urgent need for improvements in LLM training and updating processes to enhance adaptability and mitigate biases.
+Limitations
+While we strive to collect textual data from a variety of sources such as news articles, scientific papers, and forum discussions, the proportion and diversity of the texts may still fall short of fully reflecting the diversity and complexity of language.
+Moreover, such evaluations on text loss require access to the model’s logits, which is challenging for closed-source models or models accessed via APIs.
+Future Work
+Moving forward, it is crucial to refine LLM development and evaluation techniques to better handle evolving informational environments, ensuring their continued relevance and utility in various applications.
+Looking ahead, our agenda for future work includes several ambitious goals aimed at enhancing the versatility and depth of our evaluation framework. Key among these objectives are:
+  • Integrating a wider array of models, particularly focusing on those based on the Transformer architecture, such as Griffin and Mamba models.
+  • Expanding the number of checkpoints accessed for a more comprehensive temporal analysis of model performance.
+  • Delving into the relationship between token-level loss metrics and overall scores, which could offer more granular insights into model capabilities.
+  • Developing methodologies to assess the obsolescence of benchmarks and models, potentially offering a dynamic approach to evaluating the relevance and efficiency of language models over time.
+These initiatives represent our commitment to advancing the state of the art in language model evaluation, ensuring that our tools remain at the forefront of technological progress and innovation.
+Community Call to Action
+We urge the community to take note of the above phenomenon: models tend to have a better comprehension of earlier events.
+This observation supports the findings in  [26], which suggest that models are most familiar with knowledge up to 2020.
+This observation is contrary to our initial hypothesis that models may overfit current events, which is significantly more concerning.
+It suggests that caution is needed, especially for those hoping to use model judgments directly.
+Ethics Statement
+We fully acknowledge the need to adhere to ethical standards in handling and generating textual data. Hence, we commit to:
+  • We utilize various datasets in our experiments, including publicly available sources like news articles and academic papers, which do not contain personally identifiable information or offensive content.
+However, some datasets created or used during our study contain sensitive information or are bound by specific usage agreements that prevent public sharing.
+Therefore, these datasets are utilized strictly within our experiments and are not released as part of our open-source materials, adhering to ethical guidelines and agreements.
+  • Responsibility for Content Generated by Open-Source Models: We recognize that content generated by our system or through the use of open-source models involved in the evaluation might reflect or amplify biases.
+While efforts are made to minimize these biases and improve the fairness of the content, we cannot take responsibility for the generated content.
+We encourage users and developers to critically engage with these models and remain vigilant of their outputs.
+  • Addressing Data Leak Issues: Our system aims to tackle data leak issues by dynamically generating benchmarks and ensuring models generalize to new or future texts.
+We commit to continuously monitoring and updating our evaluation methods to avoid potential data leakage, ensuring the validity and reliability of our assessment results.
+  • Transparency: We are committed to maintaining the highest level of transparency in our work, including our evaluation methods, data sources used, and assessment outcomes.
+We encourage open critique and suggestions for improvement of our methodology, aiming to foster knowledge sharing and technological advancement.
+We believe that through this approach, we can offer a valuable resource to the research and development community, while fostering deep reflection and discussion on ethical issues in the artificial intelligence field.
+Licence
+5.1 Good Judgment Inc.
+For personal use of its services and content, a limited, revocable, non-exclusive license is granted.
+Without explicit written permission from Good Judgment Inc., this license restricts the collection, aggregation, copying, duplication, display, or derivative use of its content.
+5.2 arXiv
+The arXiv dataset is licensed under the Creative Commons Zero (CC0 1.0) license.
+This means it is in the public domain, and users can freely use, modify, and distribute the data without any restrictions.
+5.3 Wikipedia
+Wikipedia content is licensed under the Creative Commons Attribution-ShareAlike (CC BY-SA) license.
+This allows users to share and adapt the content as long as appropriate credit is given, and any derived works are distributed under the same license.
+Other text will not be released.
+Acknowledgement
+This work was supported by the Shenzhen Science and Technology Program (JCYJ20220818103001002), Shenzhen Doctoral Startup Funding (RCBS20221008093330065), Tianyuan Fund for Mathematics of National Natural Science Foundation of China (NSFC) (12326608), Shenzhen Key Laboratory of Cross-Modal Cognitive Computing (grant number ZDSYS20230626091302006), and Shenzhen Stability Science Program 2023, Shenzhen Key Lab of Multi-Modal Cognitive Computing.
+Appendix A Testing Statistical Hypothesis
+Due to the inherent randomness in data, directly comparing outcomes of two experiments is insufficient to justify the superiority of one outcome over the other.
+To address this issue, practitioners often resort to hypothesis testing to make convincing decisions, like those demonstrated in [27].
+A board application of hypothesis testing lies in the A/B testing literature, which assesses the performance of various features of a product and performs post-experiment inferences, as introduced in [28].
+Appendix B Details of Data Collection
+The data for our system is sourced from a diverse set of online platforms, each selected for its relevance to different data types and domains.
+We utilize a Python-based crawling framework, which is detailed in our publicly available repository.
+This framework is designed to be adaptable, allowing for the inclusion of additional sources as they become relevant.
+The primary data sources include:
+  • Financial News: Crawled from Yahoo, providing up-to-date information on global financial trends.
+  • Political News: BBC News is used to gather the latest political developments worldwide.
+  • Discussion Forums: Reddit is scraped to capture current discussions across a variety of tofigures.
+  • Online Literature: Wattpad offers a rich source of contemporary fiction and non-fiction.
+  • Encyclopedia: Wikipedia’s latest changes feed is monitored for updates across all domains.
+  • Academic Papers: arXiv is used to access the newest research across multiple scientific fields.
+  • Code Repositories: GitHub Trending provides insights into the latest developments in software engineering.
+  • Question and Answer Forums: Quora, focusing on popular topics such as Technology, Mathematics, Health, and Movies.
+To facilitate efficient and effective data collection, our system’s environment is configured with a specific Python package index URL and utilizes Playwright for web navigation and content extraction.
+This setup addresses potential challenges, such as page refresh requirements on Yahoo and hostname verification issues encountered with certain websites.
+Our system employs a combination of tools and libraries, including Requests, Playwright, and various PDF processing utilities, to collect the latest textual data from the internet.
+This approach enables us to dynamically update our benchmark datasets with fresh information, ensuring that our evaluation reflects current language use and information trends.
+Data Pre-processing For some raw content clean the text by removing these tags to ensure that only the content text is analyzed.
+Filtering out short texts: Extremely short texts may not provide enough contextual information for analysis or model training.
+Setting a minimum text length threshold of 100 chars to filter out such instances.
+External Dataset For the analysis of text likelihood and bias, we employed not only our own collection of historical arXiv data but also sampled entries from two additional datasets.
+We incorporated data from the BBC News and Wikitext datasets, which are publicly available through the Hugging Face datasets repository.
+Specifically, the datasets used were bbc_news_alltime and wikitext_alltime, accessible via the following URLs:
+Sampling from these datasets was carried out every three months, with each instance involving the collection of 50 entries per dataset.
+This structured sampling approach was designed to ensure a consistent and representative analysis of content over time, enabling an effective assessment of textual bias and likelihood trends.
+Appendix C Data Overview of One Crawl
+Tab.
+4 presented below offers a comprehensive overview of the diverse sources and categories of texts that were analyzed in one data crawl.
+Classified into distinct groups such as Academic STEM, Academic Non-STEM, Internet QA, and Internet articles, the table delineates the number of texts extracted from each source along with their average length.
+This includes detailed counts from different academic archives like arXiv, various categories from Internet question and answer platforms like Reddit and Quora, and articles from well-known online platforms including Wiki and BBC.
+Table 4: Overview of text categories with their sources counts, and average lengths in one sample crawl.
+Category Source Count Length
+Academic STEM arXiv Mathematics 50 42850
+arXiv Computer Science 50 43336
+arXiv Physics 50 42580
+arXiv Statistics 50 42619
+arXiv Electrical Engineering and Systems Science 50 43027
+Academic Non-STEM arXiv Economics 50 42592
+arXiv Quantitative Finance 50 43969
+arXiv Quantitative Biology 50 42974
+Code GitHub 30 21849
+Internet QA Reddit 583 643
+Quora Health 20 1150
+Quora Movies 69 1349
+Quora Mathematics 70 1365
+Quora Technology 80 1674
+Internet article Wiki 220 4091
+Wattpad 38 6615
+BBC 45 4792
+Yahoo 41 5373
+Appendix D Detailed results of BPC generalization on Wiki and BBC dataset
+Table 5: Percentage Change in BPC Following Release, measured by BBC news data from 2020 to 2024.3.
+Dash ("-") symbols in the table indicate missing data points due to the later release dates of some models, which means there is no available data for those specific time intervals.
+Model Base BPC BPC Change at 3 months (%) BPC Change at 6 months (%) BPC Change at 9 months (%) BPC Change at 12 months (%) mean
+OPT-13B 0.450 2.273 -2.087 1.265 0.839 0.572
+OPT-2.7B 0.475 1.970 -2.228 1.192 1.143 0.519
+LLaMA-7B 0.450 1.270 0.177 -7.776 3.266 -0.766
+Pythia-12B 0.480 -0.233 -3.888 -2.612 -1.012 -1.936
+Falcon-rw-1B 0.510 0.459 -7.814 1.201 -1.474 -1.907
+Baichuan-13B-Chat 0.467 -3.645 -1.078 0.413 - -1.436
+LLaMA-2-13B 0.416 -6.928 3.502 1.683 - -0.581
+LLaMA-2-7B 0.434 -7.314 3.092 1.279 - -0.981
+LLaMA-2-7B-Chat 0.505 -8.308 2.695 0.664 - -1.650
+Zhongjing-Base 0.465 -8.488 2.662 -1.290 - -2.372
+InternLM-Chat-7B 0.515 -2.476 -1.344 0.236 - -1.195
+Baichuan2-7B-Base 0.451 -3.059 0.595 -4.649 - -2.371
+Baichuan2-7B-Chat 0.517 -3.571 -0.038 -5.315 - -2.975
+Mistral-7B-v0.1 0.398 1.110 5.347 -0.661 - 1.932
+Phi-1.5 0.602 -1.772 1.131 -1.837 - -0.826
+Baichuan2-13B-Base 0.439 -2.943 0.882 -4.390 - -2.150
+Baichuan2-13B-Chat 0.495 -3.541 -0.995 0.232 - -1.435
+Colossal-LLaMA-2-7B-Base 0.618 2.825 3.715 - - 3.270
+Qwen-14B-Chat 0.445 2.682 4.847 - - 3.764
+Qwen-7B 0.441 1.892 3.689 - - 2.791
+Qwen-7B-Chat 0.466 1.809 3.606 - - 2.708
+Skywork-13B-Base 0.429 7.163 3.261 - - 5.212
+ChatGLM3-6B 0.832 5.567 1.684 - - 3.626
+Zephyr-7B-beta 0.435 7.340 3.928 - - 5.634
+Yi-6B 0.427 6.478 3.144 - - 4.811
+Yi-6B-Chat 0.446 3.030 -2.798 - - 0.116
+Qwen-1.8B 0.520 1.899 -2.082 - - -0.091
+Qwen-1.8B-Chat 0.599 1.635 -2.033 - - -0.199
+RWKV-v5-Eagle-7B 0.440 4.956 2.278 - - 3.617
+TinyLLaMA-1.1B-Chat-v0.6 0.513 2.539 -3.678 - - -0.569
+Phi-2 0.487 1.882 - - - 1.882
+Table 6: Percentage Change in Language Likelihood Following Release, measured by wiki data from 2020 to 2024.3.
+Dash ("-") symbols in the table indicate missing data points due to the later release dates of some models, which means there is no available data for those specific time intervals.
+Model Base BPC BPC Change at 3 months (%) BPC Change at 6 months (%) BPC Change at 9 months (%) BPC Change at 12 months (%) mean
+OPT-13B 0.075 -0.623 0.105 -0.413 -0.818 -0.437
+OPT-2.7B 0.082 -0.658 0.024 -0.528 -0.915 -0.519
+LLaMA-7B 0.052 0.214 -0.425 0.379 -0.067 0.025
+Pythia-12B 0.071 -0.578 -0.933 -0.859 -0.903 -0.818
+Falcon-rw-1B 0.083 -1.272 -1.113 -1.131 -1.207 -1.181
+Baichuan-13B-Chat 0.058 0.019 0.124 -0.125 - 0.006
+LLaMA-2-13B 0.041 1.700 1.446 1.796 - 1.647
+LLaMA-2-7B 0.050 0.774 0.349 0.517 - 0.547
+LLaMA-2-7B-Chat 0.060 0.132 -0.186 -0.093 - -0.049
+Zhongjing-Base 0.050 0.756 0.269 0.569 - 0.531
+InternLM-Chat-7B 0.080 -0.482 -0.301 -0.250 - -0.344
+Baichuan2-7B-Base 0.059 0.523 0.157 0.185 - 0.289
+Baichuan2-7B-Chat 0.068 0.343 -0.035 -0.073 - 0.078
+Mistral-7B-v0.1 0.053 0.381 -0.028 0.018 - 0.124
+Phi-1.5 0.098 0.066 -0.125 -0.248 - -0.102
+Baichuan2-13B-Base 0.053 1.019 0.797 0.900 - 0.905
+Baichuan2-13B-Chat 0.062 0.299 0.628 0.512 - 0.480
+Colossal-LLaMA-2-7B-Base 0.065 -0.042 -0.199 - - -0.121
+Qwen-14B-Chat 0.049 1.301 1.939 - - 1.620
+Qwen-7B 0.061 0.428 0.430 - - 0.429
+Qwen-7B-Chat 0.065 0.302 0.226 - - 0.264
+Skywork-13B-Base 0.050 0.075 0.276 - - 0.175
+ChatGLM3-6B 0.105 -0.187 -0.245 - - -0.216
+Zephyr-7B-beta 0.059 -0.049 -0.033 - - -0.041
+Yi-6B 0.058 -0.071 0.022 - - -0.024
+Yi-6B-Chat 0.060 -0.028 0.023 - - -0.003
+Qwen-1.8B 0.080 -0.131 -0.251 - - -0.191
+Qwen-1.8B-Chat 0.091 -0.331 -0.496 - - -0.414
+RWKV-v5-Eagle-7B 0.065 0.149 0.220 - - 0.184
+TinyLLaMA-1.1B-Chat-v0.6 0.070 -0.332 -0.473 - - -0.403
+Phi-2 0.072 -0.037 - - - -0.037
+Appendix E TBI of Models across the arXiv, BBC, and Wiki datasets
+Table 7: Comprehensive list of models with their release dates and metrics for TBI (multiplied by 1000) across the arXiv, BBC, and Wiki datasets.
+Negative TBI values indicate a trend of increasing BPC, which refers to decreasing performance on MMLU.
 Model Name Released Date TBI*1000 (arXiv) TBI*1000 (BBC) TBI*1000 (Wiki)
@@
 Phi-2 Dec 2023 -21.2 669.3 -44.7
-Table 1: Comprehensive list of models with their release dates and metrics for TBI (multiplied by 1000) across the arXiv, BBC, and Wiki datasets.
-Negative slope values indicate a trend of increasing BPC, which refers to decreasing performance.
-3.3 Temporal Generalization
-To quantify model temporal generalization, we computed the BPC for various models on the Wiki and BBC datasets.
-We established a base BPC using data from the six months prior to each model’s release.
-Subsequently, we measured the changes in BPC at 3, 6, 9 and 12 months post-release relative to the base BPC.
-The specific results are detailed in Appendix C
-The generalization of BPC in Wikipedia dataset shows that the models exhibit a relatively stable performance on Wikipedia data, which suggests a consistent ability to handle well-structured and formal text types like those found in Wikipedia.
-The generalization of BPC in BBC dataset shows that, as opposed to Wikipedia data, providing a more significant differentiation between models.
-This distinction is crucial as it reveals the models’ ability to generalize to future, unseen data.
-Notably, although there is an overall trend of increasing BPC for BBC text from 2020 to 2024, the immediate post-cutoff performance of individual models varies.
-For instance, Baichuan2 series appears to adapt better to new texts, showcasing an improved familiarity with them.
-On the other hand, models like Zephyr exhibit poorer generalization in this context.
-In our analysis across both the Wiki and BBC datasets, we have observed significant correlations between model attributes and performance decay, detailed in Table 2.
-The data shows a consistent pattern where higher benchmark scores correlate with faster performance decay, indicating a trade-off between capability and flexibility.
-This suggests that models with higher capabilities and sophistication, as reflected in their task scores, may compromise adaptability, becoming so specialized that they struggle to adjust to new or evolving data scenarios without a noticeable decrease in performance.
-Specifially, in the Wiki dataset, larger models, which generally have higher capabilities, tend to decay quicker (with a correlation of 0.504), likely due to overfitting on training data.
-Conversely, in the BBC dataset, where the information is more dynamic and rapidly changing, the correlation between model size and performance decay is almost negligible, suggesting that the impact of model size on decay is significantly less in environments that continuously present new challenges.
-Additionally, across both datasets, newer models show larger correlations, hinting that despite their advanced capabilities, they may decay faster, potentially as a consequence of these enhanced capabilities.
-These insights emphasize the importance of designing models that not only perform well on benchmarks but also maintain robustness over time and across varying data environments.
-As machine learning applications continue to grow in complexity and scale, the ability to generalize well temporally will become increasingly critical.
-This analysis serves as a cornerstone for future research aimed at enhancing the longevity and adaptability of machine learning models in dynamic settings.
-Task or model feature Correlation with BPC change rate average (Wiki) Correlation with BPC change rate average (BBC)
-Model Size 0.504 -0.044
-Release Date(month since 2015) 0.267 0.313
-MMLU 0.599 0.384
-MMLU:Humanities 0.500 0.317
-MMLU:SocialSciences 0.594 0.311
-MMLU:STEM 0.393 0.186
-MMLU:Other 0.528 0.153
-GSM8 0.473 0.188
-Ar20c 0.380 -0.071
-Hellawasg 0.581 0.248
-TrustfulQA 0.200 0.485
-Winogrande 0.632 0.204
-Longbench 0.404 0.549
-Table 2: Combined correlation between task scores and performance decay on Wikipedia and BBC
-3.4 Correlation with Existing Benchmarks
+Appendix F Most common tags of the Gjo questions
+Table 8: Frequency of Topics
+Topic Count
+Business 648
+Non-US Politics 453
+Security and Conflict 446
+Society 425
+US Politics 363
+Technology 359
+Finance 299
+Elections and Referenda 288
+Economic Indicators 273
+Foreign Policy 271
+US Policy 247
+Health 245
+Leader Entry/Exit 219
+Economic Policy 160
+Sports 148
+Entertainment 148
+Environment 80
+Open 16
+Table 9: Most common tags of the Gjo questions
+Appendix G Correlation with Existing Benchmarks
+Following our previous chapter I, we continue to explore the correlation between prediction performance and existing benchmarks, using the same benchmarks as in I.
+We have categorized the tags of questions into several groups: Societal Issues Items, Political and Government Affairs, Economic and Business Items, and Science and Technology Items.
+Each group represents a significant area of interest.
+In our analysis, we focus on how the correlation between these categories and various benchmarks evolves over time.
+We use correlation coefficients to measure the strength of association between the performance scores of questions within each category and the benchmarks across different times.
+Notably, our analysis revealed that while the Economic and Business Items and Science and Technology Items on base models displayed relatively high initial correlations, there appears to be little to no relationship in Societal issues Items and Political and Government Affairs, we think such topics are influenced by a multitude of unpredictable factors such as social changes, policy adjustments, and public sentiment.
+These elements make accurate predictions particularly challenging, leading to the observed low correlation with performance metrics.
+Detailed figures are in Appendix N.
+Despite relatively high initial correlations in 2 question groups, there is a general lack of correlation between model predictions and existing benchmarks (mean correlation coefficient around zero).
+In addition, significant deviations occur in nearly half of the cases from early stages to subsequent periods.
+This suggests that the models’ understanding and alignment with benchmarks can vary substantially over time.
+So temporal generalization provides novel insights and draws attention to an aspect that has been often ignored by previous benchmarks.
+Appendix H Analysis of Language Likelihood Values and Benchmark Correlation Across Text and Model Groups
+(a) STEM: base models
+(b) non-STEM: base models
+(c) Non-STEM: chat models
+Figure 4: The heatmaps display the correlations by loss for a group of models across different text chunk lengths and different benchmark scores.
+Each row represents a specific truncation length, while the columns represent different benchmark score categories.
+The values in the cells indicate the corresponding correlation coefficients.
+(a) presents the correlations by loss for all 18 base models on STEM Academia texts.
+(b) presents the correlations by loss for all 18 base models on non-STEM Academia texts.
+(c) presents the correlations by loss for all 15 chat models on non-STEM Academia texts.
+(a) STEM: base models
+(b) non-STEM: base models
+(c) Non-STEM: chat models
+Figure 5: The heatmaps display the correlations by loss for a group of models across different text chunk lengths and different benchmark scores.
+Each row represents a specific truncation length, while the columns represent different benchmark score categories.
+The values in the cells indicate the corresponding correlation coefficients.
+(a) presents the correlations by loss for all 18 base models on STEM Academia texts.
+(b) presents the correlations by loss for all 18 base models on non-STEM Academia texts.
+(c) presents the correlations by loss for all 15 chat models on non-STEM Academia
+(a) STEM: all chat models
+(b) STEM: 7B level chat models
+Figure 6: (a) presents the correlations by loss for all 15 chat models on STEM Academia texts. (b) presents the correlations by loss for all 9 7B level chat models on STEM Academia texts.
+From Fig.
+6 we can conclude that the scale of models being observed does not significantly alter outcomes, suggesting that performance correlations across benchmarks are generally consistent regardless of model size.
+Appendix I Correlation with Existing Benchmarks
 The Selected Existing Benchmarks
@@
 The selected benchmarks encompass common sense reasoning, depth and breadth of knowledge across diverse subjects, truthfulness, natural language understanding, and mathematical reasoning, ensuring a well-rounded assessment of the model’s performance.
+Table 10: Overview of AI Benchmarks and Their Evaluation Focus
 Benchmark Description
-HellaSwag Tests an AI’s common sense reasoning by requiring it to complete sentences or narratives that reflect everyday logic, highlighting the model’s ability to predict logical continuations.
-(Zellers et al., 2019)
-MMLU Evaluates an AI’s breadth and depth of understanding across a diverse range of subjects, from science to literature, showcasing its versatility and comprehensive grasp of human language.
-(Hendrycks et al., 2021)
+HellaSwag Tests an AI’s common sense reasoning by requiring it to complete sentences or narratives that reflect everyday logic, highlighting the model’s ability to predict logical continuations. [29]
+MMLU Evaluates an AI’s breadth and depth of understanding across a diverse range of subjects, from science to literature, showcasing its versatility and comprehensive grasp of human language. [30]
 TruthfulQA Focuses on the model’s capacity to provide honest and accurate responses, particularly to questions where the truth may not be intuitive, testing its commitment to truthfulness and its proficiency in avoiding misinformation.
-(Lin et al., 2022)
-Winogrande Assesses common sense reasoning and natural language understanding through sentence completion tasks, measuring an AI’s ability to apply common sense knowledge in language processing.
-(Sakaguchi et al., 2019)
-GSM8K Challenges AI models with elementary-level math problems, testing their mathematical reasoning capabilities and their understanding and application of basic math concepts. (Cobbe et al., 2021)
-Table 3: Overview of AI Benchmarks and Their Evaluation Focus
+[31]
+Winogrande Assesses common sense reasoning and natural language understanding through sentence completion tasks, measuring an AI’s ability to apply common sense knowledge in language processing. [32]
+GSM8K Challenges AI models with elementary-level math problems, testing their mathematical reasoning capabilities and their understanding and application of basic math concepts. [33]
 Base vs. Chat Models
-Table 4 shows the comparative analysis of base and tuned models.
+Tab.
+11 shows the comparative analysis of base and tuned models.
 Our findings suggest that: 1) Base models exhibit a stronger correlation with established benchmarks, implying that these models, in their fundamental form, possess a robust capacity for modeling language.
@@
 2) The exploration of tuned models reveals potential deeper patterns in model tuning, suggesting that specific tuning strategies may unlock further enhancements in model performance across different tasks and domains.
+Table 11: Correlation with Different Benchmarks, Base vs. Chat Models
 Benchmarks Loss of Base Model Loss of Chat Model Delta
@@
 TrustfulQA 0.05 0.39 0.34
-Table 4: Correlation with Different Benchmarks, Base vs. Chat Models
 Correlation alone Text Length
-Table 5 is about a long text understanding benchmark Longbench.
+Tab. 12 is about a long text understanding benchmark Longbench.
 For chat models in this task, the correlation between BPC and scores is logical—higher probabilities of generating the text correlate with higher scores, and this correlation strengthens with longer text lengths.
@@
 Thus, relying solely on language likelihood as an indicator of overall performance might overlook crucial aspects of cognitive and interpretative abilities that are essential for more complex applications.
+Table 12: Correlation with Longbench values across different lengths
 Length 10 30 50 100 200 300 400 500 800 1000 1200 1500 2000
@@
 Chat Models 0.01 -0.07 -0.16 -0.39 -0.53 -0.58 -0.61 -0.63 -0.65 -0.65 -0.67 -0.67 -0.68
-Table 5: Correlation with Longbench values across different lengths
 On the Genres of Corpora
-As can be seen in Appendix  D, whether all models are observed or only those of a specific scale does not make a big difference.
+As can be seen in Appendix  H, whether all models are observed or only those of a specific scale does not make a big difference.
 Comparisons across different model sizes show that the differences are not substantial, indicating consistency in performance correlations across benchmarks.
-Further analysis from Table 6 within the MMLU benchmark sections reveals interesting patterns.
+Further analysis from Tab.
+13 within the MMLU benchmark sections reveals interesting patterns.
 While we assume models to be generally excel in domains closely aligned with their training data, this correlation is not strong.
@@
 Similarly, Internet data, including forums and news, also shows comparative closeness in performance, indicating nuanced relationships across different types of content.
+Table 13: Correlation Coefficients between BPC in specific areas and Domain-Specific Performance
 Benchmark Academia STEM Academia non-STEM Code Internet Article Internet Chat
@@
 MMLU:Social Sciences -0.49 -0.42 -0.46 -0.25 -0.31
-Table 6: Correlation Coefficients between BPC in specific areas and Domain-Specific Performance
 Additionally, compared to chat forum texts, more formal academic texts demonstrate a stronger correlation with benchmarks.
 The genre of text has a minor impact on the sign of the correlation coefficient, which is relatively small compared to the differences across model types.
-3.5 The Drawbacks of Evaluating Language Likelihood
-Inaccessibility to API-based Models
-A fundamental limitation of using language likelihood as a metric lies in its inapplicability to closed-source models, where key indicators such as perplexity, logits, or losses are inaccessible.
-This restriction excludes a broad spectrum of proprietary models from evaluation and significantly narrows the scope of comparative analysis.
-The consequent inability to assess these models hampers our ability to gain a comprehensive understanding of the language technology landscape and the capabilities of different models within it.
-Susceptibility to Training Focus
-The effectiveness of language likelihood as a metric can be compromised by models that are finely tuned or trained with a particular emphasis on certain formats or domains.
-In such cases, the likelihood measure might disproportionately favor these specialized areas, skewing the overall evaluation and presenting a distorted view of a model’s general performance.
-4 Temporal Generalization in terms of Prognostication Prediction
-The above Drawbacks in Sec.
-3.5 motivated us to research temporal generalization and bias from the perspective of future prediction.
-In the context of the model release date, events that occurred before this date are considered ”past events”, while those that took place after are termed ”future events”.
-In this chapter, we have divided the experiments into two categories based on this criterion: Bias and Temporal Generation.
-4.1 Experimental Settings
-4.1.1 Data Source
-We have sourced all questions from  Good Judgment Open and all binary questions from  Metaculus through web scraping and API downloads.
-The collected data includes initiation and conclusion dates, domains, questions, descriptions, and answers, which are ordered appropriately for further processing.
-4.1.2 Extraction of Predictive Responses
-Using the lm-eval-harness  (Gao et al., 2023) tool, we predict model choices based on probability associated with each option.
-This probability-driven method serves as a reliable metric for evaluating the predictive performance of language models across a variety of topics and question formats.
-4.2 Temporal Bias
-As can be seen in Table 7, most analyzed models exhibit a negative slope.
-This is an uncommon scenario in a typical statistical analysis, indicating a general tendency across models to perform better or be more aligned with historical data than recent data.
-We define the following classification thresholds to discern subtle differences in how models handle temporal information, despite an overall inclination towards nostalgia.
-  • Positive TBI (Neophilia Bias): TBI values greater than 0.001 suggest the model has a neophilia bias.
-This implies the model’s performance is better on more recent data, potentially favoring newer over older training data.
-  • Near-Zero TBI (Balanced): TBI values between -0.001 and 0.001 indicate a balanced performance over time.
-A TBI of 0, in particular, suggests that the model’s performance is evenly distributed around the mean, showing no temporal bias.
-  • Negative TBI (Nostalgia Bias): TBI values less than -0.001 indicate a nostalgia bias, suggesting the model may favor historical data, possibly overlooking important recent information.
-The choice of ±0.001 as the threshold for these categories is designed to fine-tune the sensitivity of our analysis to subtle temporal biases, ensuring that the classification is both meaningful and practically significant.
-model TBI y-axis intercept bias category
-OPT-13B -1.37 0.70 Nostalgia
-OPT-2.7B -2.85 0.97 Nostalgia
-LLaMA-7B -3.71 0.96 Nostalgia
-Pythia-12B -3.90 0.95 Nostalgia
-Falcon-rw-1B -2.47 0.84 Nostalgia
-Baichuan-13B-Chat -2.79 0.98 Nostalgia
-Baichuan-7B-Chat -1.17 0.80 Nostalgia
-GPT-3.5-turbo-0613 -0.81 0.86 Symmetrical
-LLaMA-2-7B -1.66 0.75 Nostalgia
-LLaMA-2-13B -0.80 0.73 Symmetrical
-LLaMA-2-7B-Chat -3.92 0.93 Nostalgia
-Zhongjing-Base 0.08 0.63 Symmetrical
-InternLM-Chat-7B -1.76 0.94 Nostalgia
-Baichuan2-7B-Base -3.50 0.84 Nostalgia
-Baichuan2-7B-Chat -3.71 0.94 Nostalgia
-Mistral-7B-v0.1 -0.47 0.75 Symmetrical
-Phi-1.5 0.24 0.38 Symmetrical
-Baichuan2-13B-Base -3.06 0.72 Nostalgia
-Baichuan2-13B-Chat -1.53 0.70 Nostalgia
-Colossal-LLaMA-2-7B-Base -3.38 0.91 Nostalgia
-Qwen-14B-Chat -2.31 0.93 Nostalgia
-Qwen-7B -0.35 0.69 Symmetrical
-Qwen-7B-Chat -2.98 0.62 Nostalgia
-Skywork-13B-Base -1.46 0.65 Nostalgia
-ChatGLM3-6B -4.46 0.92 Nostalgia
-Zephyr-7B-beta -2.73 0.99 Nostalgia
-Yi-6B -5.18 0.98 Nostalgia
-Yi-6B-Chat 0.34 0.71 Symmetrical
-Qwen-1.8B -3.59 0.80 Nostalgia
-Qwen-1.8B-Chat -4.32 0.96 Nostalgia
-GPT-4-1106 -0.08 0.96 Symmetrical
-TinyLLaMA-1.1B-Chat-v0.6 -1.78 0.53 Nostalgia
-Gemini-pro-1213 -0.43 0.76 Symmetrical
-Phi-2 -1.54 0.77 Nostalgia
-Table 7: Predict performance of models: TBI(slope) and y-axis intercept are linear fit parameters, bias category classifies the nature of the bias based on TBI
-Table 8 shows an analysis covering the period from 2015 to 2024, with the year 2019 serving as a pivotal midpoint, a distinct pattern emerges from the performance data of various language models: all models reached their peak accuracy before 2020.
-This observation points to a notable trend toward ”nostalgia,” where models demonstrate optimal performance on data from the earlier part of the timeline analyzed.
-This tendency highlights the models’ alignment with the informational characteristics prevalent prior to the latter half of the observed period, suggesting their training data may disproportionately reflect earlier times.
-The peak accuracy periods analyzed are based on the average accuracy of all questions within three-month intervals, to provide a stable and robust assessment of each model’s performance.
-Peak Period Model Peak Time of Accuracy Cutoff Time Interpolation
-Post-2019 TinyLLaMA1.1B 2019-07 2023-11 52
-2017 to 2018 GPT4-1106 2018-07 2023-06 59
-LLaMA2-13B 2018-07 2023-02 55
-Baichuan2-13B-Base 2018-01 2023-06 65
-2015 to 2016 Baichuan2-7B-Chat 2016-01 2023-08 91
-Colossal-LLaMA-2-7B-Base 2016-01 2023-09 92
-LLaMA2-7B 2016-01 2023-02 85
-Table 8: Analysis about the time of the best-achieved accuracy and its cutoff time.
-4.3 Temporal Generalization
-It is crucial to note that for models, any data post-release represents the future, allowing us to assess their generalization capabilities on upcoming events.
-In the analysis of temporal generalization, it is important to acknowledge that the performance of models post-release primarily reflects their ability to predict future events based on past training.
-As evidenced by the data in Table 9, many models show significant fluctuations in accuracy percentages following their release.
-This variability suggests that immediately after the cutoff event, most models are essentially guessing, leading to high volatility in performance metrics.
-Given the substantial fluctuations observed post-release, the temporal generalization of current models may appear to have limited significance.
-This is largely because these models exhibit weak capabilities in adapting to new, unseen data, often resulting in what can be described as near-random performance at the cutoff point.
-Such outcomes highlight the models’ limited understanding of the world and their inability to effectively generalize from past information to future contexts.
-However, this should not detract from the potential utility of these measures in the future.
-As language models continue to evolve and become more sophisticated in their world understanding, the metric of temporal generalization will likely become more relevant and insightful.
-Future models that better comprehend and adapt to the world will provide more meaningful insights into their generalization capabilities over time, thereby enhancing the value of this assessment.
-The current limitations underscore the need for further development in language modeling, particularly in improving the models’ ability to generalize across time.
-As we progress towards models with a deeper and more accurate understanding of complex world dynamics, temporal generalization measures will play a crucial role in evaluating their long-term effectiveness and adaptability.
-Base adjusted ACC Adjusted ACC Change at 3 months (%) Adjusted ACC Change at 6 months (%) Adjusted ACC Change at 9 months (%) adjusted ACC Change at 12 months (%) Mean (%)
-Model
-OPT-2.7B 0.86 0.11 -13.63 -23.41 -18.98 -13.98
-LLaMA-7B 0.66 -0.31 -20.59 5.09 9.89 -1.48
-Pythia-12B 0.77 -27.68 -24.85 -6.22 -3.68 -15.61
-Falcon-rw-1B 0.69 -34.51 -8.30 7.15 -4.52 -10.04
-Baichuan-13B-Chat 0.64 0.69 12.79 8.00 0.13 5.40
-GPT-3.5-turbo-0613 0.65 -29.81 16.71 26.91 -39.69 -6.47
-LLaMA-2-13B 0.57 29.19 38.00 34.54 - 33.91
-LLaMA-2-7B 0.54 25.57 41.09 27.09 - 31.25
-LLaMA-2-7B-Chat 0.61 25.39 23.81 12.52 - 20.58
-Zhongjing-Base 0.59 25.16 37.70 14.65 - 25.83
-InternLM-Chat-7B 0.48 3.43 25.13 21.65 - 16.74
-Baichuan2-7B-Base 0.60 14.99 32.31 4.81 - 17.37
-Baichuan2-7B-Chat 0.60 16.12 22.67 -18.59 - 6.73
-Mistral-7B-v0.1 0.61 30.25 34.87 -22.79 - 14.11
-Phi-1.5 0.65 19.30 10.78 -17.57 - 4.17
-Baichuan2-13B-Base 0.61 10.81 17.67 6.75 - 11.75
-Colossal-LLaMA-2-7B-Base 0.58 15.99 3.33 - - 9.66
-Qwen-14B-Chat 0.59 14.68 14.30 - - 14.49
-Qwen-7B 0.62 7.45 -7.31 - - 0.07
-Qwen-7B-Chat 0.48 60.26 74.85 - - 67.55
-Skywork-13B-Base 0.61 20.20 5.04 - - 12.62
-ChatGLM3-6B 0.60 3.24 6.29 - - 4.76
-Zephyr-7B-beta 0.61 14.33 18.22 - - 16.28
-Yi-6B 0.63 8.42 -9.47 - - -0.53
-Yi-6B-Chat 0.67 13.11 -6.19 - - 3.46
-Qwen-1.8B 0.68 -6.93 -33.98 - - -20.46
-Qwen-1.8B-Chat 0.73 4.74 -8.91 - - -2.08
-RWKV-v5-Eagle-7B 0.70 -0.09 -20.38 - - -10.23
-GPT-4-1106 0.42 6.67 -7.11 - - -0.22
-TinyLLaMA-1.1B-Chat-v0.6 0.71 13.27 -16.30 - - -1.51
-Gemini-pro-1213 0.71 -15.15 -41.28 - - -28.21
-Phi-2 0.81 7.11 - - - 7.11
-Table 9: Percentage Change in Model Accuracy Following Release
-4.4 Correlation with Existing Benchmarks
-Following our previous chapter 3.4, we continue to explore the correlation between prediction performance and existing benchmarks, using the same benchmarks as in 3.4.
-We have categorized the tags of questions into several groups: Societal Issues Items, Political and Government Affairs, Economic and Business Items, and Science and Technology Items.
-Each group represents a significant area of interest.
-In our analysis, we focus on how the correlation between these categories and various benchmarks evolves over time.
-We use correlation coefficients to measure the strength of association between the performance scores of questions within each category and the benchmarks across different times.
-Notably, our analysis revealed that while the Economic and Business Items and Science and Technology Items on base models displayed relatively high initial correlations, there appears to be little to no relationship in Societal issues Items and Political and Government Affairs, we think such topics are influenced by a multitude of unpredictable factors such as social changes, policy adjustments, and public sentiment.
-These elements make accurate predictions particularly challenging, leading to the observed low correlation with performance metrics.
-Detailed figures are in Appendix H.
-The graphical representations in our analysis plot the changes in correlation coefficients over time.
-Each line in the graph represents the changing relationship between a category’s average adjusted performance and a specific benchmark, highlighting periods of stronger or weaker association.
-For a detailed breakdown of the categories and the number of entries in each, please refer to the Appendix G.
-This provides a clearer picture of the dataset’s composition and the scope of our analysis.
-4.5 Community Call to Action
-We urge the community to take note of the above phenomenon: models tend to have a better comprehension of earlier events.
-This observation supports the findings in  Cheng et al.
-(2024), which suggest that models are most familiar with knowledge up to 2020.
-This observation is contrary to our initial hypothesis that models may overfit current events, which is significantly more concerning.
-It suggests that caution is needed, especially for those hoping to use model judgments directly.
-5 Conclusion
-We rigorously define and qualify temporal generalization and bias within LLMs, uncovering significant challenges in their ability to predict and adapt to future contexts.
-The experiments reveal that LLMs often struggle with temporal biases, affecting their reliability in real-world applications, particularly in dynamic scenarios requiring accurate forecasting.
-The evaluation framework we provide allows for an in-depth assessment of these models’ performance over time.
-The findings underscore the urgent need for improvements in LLM training and updating processes to enhance their adaptability and mitigate biases.
-Moving forward, it is crucial to refine LLM development and evaluation techniques to better handle evolving informational environments, ensuring their continued relevance and utility in various applications.
-Looking ahead, our agenda for future work includes several ambitious goals aimed at enhancing the versatility and depth of our evaluation framework. Key among these objectives are:
-  • Integrating a wider array of models, particularly focusing on those based on the Transformer architecture, such as Griffin and Mamba models.
-  • Expanding the number of checkpoints accessed for a more comprehensive temporal analysis of model performance.
-  • Delving into the relationship between token-level loss metrics and overall scores, which could offer more granular insights into model capabilities.
-  • Developing methodologies to assess the obsolescence of benchmarks and models, potentially offering a dynamic approach to evaluating the relevance and efficiency of language models over time.
-These initiatives represent our commitment to advancing the state of the art in language model evaluation, ensuring that our tools remain at the forefront of technological progress and innovation.
-Limitations
-Despite the system’s innovative approach to evaluating LLMs using the latest textual data, it is important to acknowledge certain limitations.
-One significant limitation is the system’s applicability primarily to open LLMs that provide access to their logits.
-This dependency arises because our evaluation mechanism relies on analyzing these logits to compute loss metrics.
-Consequently, our system may not be directly applicable to non-open LLMs, where such detailed internal outputs are not accessible.
-This constraint highlights the need for adaptable evaluation strategies that can cater to a broader range of LLMs, including those with closed architectures.
-Ethics Statement
-We fully acknowledge the need to adhere to ethical standards in handling and generating textual data. Hence, we commit to:
-  • Handling User-Uploaded Content: We respect the privacy and copyright of content uploaded by users.
-All data provided by users will be used solely for the purpose of model evaluation and not for any other purposes.
-Reasonable measures will be taken to protect this data from unauthorized access, use, or disclosure.
-  • Responsibility for Content Generated by Open-Source Models: We recognize that content generated by our system or through the use of open-source models involved in the evaluation might reflect or amplify biases.
-While efforts are made to minimize these biases and improve the fairness of the content, we cannot take responsibility for the generated content.
-We encourage users and developers to critically engage with these models and remain vigilant of their outputs.
-  • Addressing Data Leak Issues: Our system aims to tackle data leak issues by dynamically generating benchmarks and ensuring models generalize to new or future texts.
-We commit to continuously monitoring and updating our evaluation methods to avoid potential data leakage, ensuring the validity and reliability of our assessment results.
-  • Transparency: We are committed to maintaining the highest level of transparency in our work, including our evaluation methods, data sources used, and assessment outcomes.
-We encourage open critique and suggestions for improvement of our methodology, aiming to foster knowledge sharing and technological advancement.
-We believe that through this approach, we can offer a valuable resource to the research and development community, while fostering deep reflection and discussion on ethical issues in the artificial intelligence field.
-Appendix A Details of Data Collection
-The data for our system is sourced from a diverse set of online platforms, each selected for its relevance to different data types and domains.
-We utilize a Python-based crawling framework, which is detailed in our publicly available repository.
-This framework is designed to be adaptable, allowing for the inclusion of additional sources as they become relevant.
-The primary data sources include:
-  • Financial News: Crawled from Yahoo, providing up-to-date information on global financial trends.
-  • Political News: BBC News is used to gather the latest political developments worldwide.
-  • Discussion Forums: Reddit is scraped to capture current discussions across a variety of tofigures.
-  • Online Literature: Wattpad offers a rich source of contemporary fiction and non-fiction.
-  • Encyclopedia: Wikipedia’s latest changes feed is monitored for updates across all domains.
-  • Academic Papers: arXiv is used to access the newest research across multiple scientific fields.
-  • Code Repositories: GitHub Trending provides insights into the latest developments in software engineering.
-  • Question and Answer Forums: Quora, focusing on popular topics such as Technology, Mathematics, Health, and Movies.
-To facilitate efficient and effective data collection, our system’s environment is configured with a specific Python package index URL and utilizes Playwright for web navigation and content extraction.
-This setup addresses potential challenges, such as page refresh requirements on Yahoo and hostname verification issues encountered with certain websites.
-Our system employs a combination of tools and libraries, including Requests, Playwright, and various PDF processing utilities, to collect the latest textual data from the internet.
-This approach enables us to dynamically update our benchmark datasets with fresh information, ensuring that our evaluation reflects current language use and information trends.
-Data Pre-processing For some raw content clean the text by removing these tags to ensure that only the content text is analyzed.
-Filtering out short texts: Extremely short texts may not provide enough contextual information for analysis or model training.
-Setting a minimum text length threshold of 100 chars to filter out such instances.
-External Dataset For the analysis of text likelihood and bias, we employed not only our own collection of historical arXiv data but also sampled entries from two additional datasets.
-We incorporated data from the BBC News and Wikitext datasets, which are publicly available through the Hugging Face datasets repository.
-Specifically, the datasets used were bbc_news_alltime and wikitext_alltime, accessible via the following URLs:
-Sampling from these datasets was carried out every three months, with each instance involving the collection of 50 entries per dataset.
-This structured sampling approach was designed to ensure a consistent and representative analysis of content over time, enabling an effective assessment of textual bias and likelihood trends.
-Appendix B Data Overview of One Crawl
-Table 10 presented below offers a comprehensive overview of the diverse sources and categories of texts that were analyzed in one data crawl.
-Classified into distinct groups such as Academic STEM, Academic Non-STEM, Internet QA, and Internet articles, the table delineates the number of texts extracted from each source along with their average length.
-This includes detailed counts from different academic archives like arXiv, various categories from Internet question and answer platforms like Reddit and Quora, and articles from well-known online platforms including Wiki and BBC.
-Category Source Count Length
-Academic STEM arXiv Mathematics 50 42850
-arXiv Computer Science 50 43336
-arXiv Physics 50 42580
-arXiv Statistics 50 42619
-arXiv Electrical Engineering and Systems Science 50 43027
-Academic Non-STEM arXiv Economics 50 42592
-arXiv Quantitative Finance 50 43969
-arXiv Quantitative Biology 50 42974
-Code GitHub 30 21849
-Internet QA Reddit 583 643
-Quora Health 20 1150
-Quora Movies 69 1349
-Quora Mathematics 70 1365
-Quora Technology 80 1674
-Internet article Wiki 220 4091
-Wattpad 38 6615
-BBC 45 4792
-Yahoo 41 5373
-Table 10: Overview of text categories with their sources counts, and average lengths in one sample crawl.
-Appendix C Detailed results of BPC generalization on Wiki and BBC dataset
-Model Base BPC BPC Change at 3 months (%) BPC Change at 6 months (%) BPC Change at 9 months (%) BPC Change at 12 months (%) mean
-OPT-13B 0.450 2.273 -2.087 1.265 0.839 0.572
-OPT-2.7B 0.475 1.970 -2.228 1.192 1.143 0.519
-LLaMA-7B 0.450 1.270 0.177 -7.776 3.266 -0.766
-Pythia-12B 0.480 -0.233 -3.888 -2.612 -1.012 -1.936
-Falcon-rw-1B 0.510 0.459 -7.814 1.201 -1.474 -1.907
-Baichuan-13B-Chat 0.467 -3.645 -1.078 0.413 - -1.436
-LLaMA-2-13B 0.416 -6.928 3.502 1.683 - -0.581
-LLaMA-2-7B 0.434 -7.314 3.092 1.279 - -0.981
-LLaMA-2-7B-Chat 0.505 -8.308 2.695 0.664 - -1.650
-Zhongjing-Base 0.465 -8.488 2.662 -1.290 - -2.372
-InternLM-Chat-7B 0.515 -2.476 -1.344 0.236 - -1.195
-Baichuan2-7B-Base 0.451 -3.059 0.595 -4.649 - -2.371
-Baichuan2-7B-Chat 0.517 -3.571 -0.038 -5.315 - -2.975
-Mistral-7B-v0.1 0.398 1.110 5.347 -0.661 - 1.932
-Phi-1.5 0.602 -1.772 1.131 -1.837 - -0.826
-Baichuan2-13B-Base 0.439 -2.943 0.882 -4.390 - -2.150
-Baichuan2-13B-Chat 0.495 -3.541 -0.995 0.232 - -1.435
-Colossal-LLaMA-2-7B-Base 0.618 2.825 3.715 - - 3.270
-Qwen-14B-Chat 0.445 2.682 4.847 - - 3.764
-Qwen-7B 0.441 1.892 3.689 - - 2.791
-Qwen-7B-Chat 0.466 1.809 3.606 - - 2.708
-Skywork-13B-Base 0.429 7.163 3.261 - - 5.212
-ChatGLM3-6B 0.832 5.567 1.684 - - 3.626
-Zephyr-7B-beta 0.435 7.340 3.928 - - 5.634
-Yi-6B 0.427 6.478 3.144 - - 4.811
-Yi-6B-Chat 0.446 3.030 -2.798 - - 0.116
-Qwen-1.8B 0.520 1.899 -2.082 - - -0.091
-Qwen-1.8B-Chat 0.599 1.635 -2.033 - - -0.199
-RWKV-v5-Eagle-7B 0.440 4.956 2.278 - - 3.617
-TinyLLaMA-1.1B-Chat-v0.6 0.513 2.539 -3.678 - - -0.569
-Phi-2 0.487 1.882 - - - 1.882
-Table 11: Percentage Change in BPC Following Release, measured by BBC news data from 2020 to 2024.3.
-Dash (”-”) symbols in the table indicate missing data points due to the later release dates of some models, which means there is no available data for those specific time intervals.
-Model Base BPC BPC Change at 3 months (%) BPC Change at 6 months (%) BPC Change at 9 months (%) BPC Change at 12 months (%) mean
-OPT-13B 0.075 -0.623 0.105 -0.413 -0.818 -0.437
-OPT-2.7B 0.082 -0.658 0.024 -0.528 -0.915 -0.519
-LLaMA-7B 0.052 0.214 -0.425 0.379 -0.067 0.025
-Pythia-12B 0.071 -0.578 -0.933 -0.859 -0.903 -0.818
-Falcon-rw-1B 0.083 -1.272 -1.113 -1.131 -1.207 -1.181
-Baichuan-13B-Chat 0.058 0.019 0.124 -0.125 - 0.006
-LLaMA-2-13B 0.041 1.700 1.446 1.796 - 1.647
-LLaMA-2-7B 0.050 0.774 0.349 0.517 - 0.547
-LLaMA-2-7B-Chat 0.060 0.132 -0.186 -0.093 - -0.049
-Zhongjing-Base 0.050 0.756 0.269 0.569 - 0.531
-InternLM-Chat-7B 0.080 -0.482 -0.301 -0.250 - -0.344
-Baichuan2-7B-Base 0.059 0.523 0.157 0.185 - 0.289
-Baichuan2-7B-Chat 0.068 0.343 -0.035 -0.073 - 0.078
-Mistral-7B-v0.1 0.053 0.381 -0.028 0.018 - 0.124
-Phi-1.5 0.098 0.066 -0.125 -0.248 - -0.102
-Baichuan2-13B-Base 0.053 1.019 0.797 0.900 - 0.905
-Baichuan2-13B-Chat 0.062 0.299 0.628 0.512 - 0.480
-Colossal-LLaMA-2-7B-Base 0.065 -0.042 -0.199 - - -0.121
-Qwen-14B-Chat 0.049 1.301 1.939 - - 1.620
-Qwen-7B 0.061 0.428 0.430 - - 0.429
-Qwen-7B-Chat 0.065 0.302 0.226 - - 0.264
-Skywork-13B-Base 0.050 0.075 0.276 - - 0.175
-ChatGLM3-6B 0.105 -0.187 -0.245 - - -0.216
-Zephyr-7B-beta 0.059 -0.049 -0.033 - - -0.041
-Yi-6B 0.058 -0.071 0.022 - - -0.024
-Yi-6B-Chat 0.060 -0.028 0.023 - - -0.003
-Qwen-1.8B 0.080 -0.131 -0.251 - - -0.191
-Qwen-1.8B-Chat 0.091 -0.331 -0.496 - - -0.414
-RWKV-v5-Eagle-7B 0.065 0.149 0.220 - - 0.184
-TinyLLaMA-1.1B-Chat-v0.6 0.070 -0.332 -0.473 - - -0.403
-Phi-2 0.072 -0.037 - - - -0.037
-Table 12: Percentage Change in Language Likelihood Following Release, measured by wiki data from 2020 to 2024.3.
-Dash (”-”) symbols in the table indicate missing data points due to the later release dates of some models, which means there is no available data for those specific time intervals.
-Appendix D Analysis of Language Likelihood Values and Benchmark Correlation Across Text and Model Groups
-(a) STEM: base models
-(b) non-STEM: base models
-(c) Non-STEM: chat models
-Figure 5: The heatmaps display the correlations by loss for a group of models across different text chunk lengths and different benchmark scores.
-Each row represents a specific truncation length, while the columns represent different benchmark score categories.
-The values in the cells indicate the corresponding correlation coefficients.
-(a) presents the correlations by loss for all 18 base models on STEM Academia texts.
-(b) presents the correlations by loss for all 18 base models on non-STEM Academia texts.
-(c) presents the correlations by loss for all 15 chat models on non-STEM Academia texts.
-(a) STEM: all chat models
-(b) STEM: 7B level chat models
-Figure 6: (a) presents the correlations by loss for all 15 chat models on STEM Academia texts.
-(b) presents the correlations by loss for all 9 7B level chat models on STEM Academia texts.
-We can conclude that the scale of models being observed does not significantly alter outcomes, suggesting that performance correlations across benchmarks are generally consistent regardless of model size.
-Appendix E Function Fitting Strategies
+Appendix J Function Fitting Strategies
 We considered several fitting strategies, each with its unique mathematical approach to model decay:
@@
 The variance of the residuals for each model was calculated as follows:
+Table 14: Variance of Residuals for Different Models
 Model Variance
@@
 Logarithmic 3.33e-3
-Table 13: Variance of Residuals for Different Models
-E.0.1 Findings
+J.0.1 Findings
 Our analysis found that the linear model provided the best fit, with the lowest variance among the residuals.
@@
 Below is a partial presentation of rankings and scores for selected models, showcasing their weighted accuracy scores. The complete table can be found in Table 15.
+Table 15: Model Scores
 Model Name Score
@@
 TinyLLaMA-1.1B-Chat-v0.6 366.125
-Table 14: Model Scores
-Appendix F Details of Model Accuracy Peaks
-Table15 captures the peak performance periods of various models on Freshbench. It lists specific models along with their respective dates when they achieved the best performance.
+Appendix K Details of Model Accuracy Peaks
+Tab. 16 captures the peak performance periods of various models on Freshbench. It lists specific models along with their respective dates when they achieved the best performance.
+Table 16: Detailed Record of Model Accuracy Peaks
 Date Models
@@
 2024-07
-Table 15: Detailed Record of Model Accuracy Peaks
-Appendix G Question Tag Details
+Appendix L Question Tag Details
+Table 17: Detailed breakdown and sum of entries per category within each group of benchmarks.
 Group Category Number of Entries
@@
 Total: 346
-Table 16: Detailed breakdown and sum of entries per category within each group of benchmarks.
-Appendix H Prediction Correlation over Time
+Appendix M Analysis of Best-Achieved Accuracy and Cutoff Time
+Table 18: Analysis about the time of the best-achieved accuracy and its cutoff time.
+Peak Period Model Peak Time of Accuracy Cutoff Time Interpolation
+Post-2019 TinyLLaMA1.1B 2019-07 2023-11 52
+2017 to 2018 GPT4-1106 2018-07 2023-06 59
+LLaMA2-13B 2018-07 2023-02 55
+Baichuan2-13B-Base 2018-01 2023-06 65
+2015 to 2016 Baichuan2-7B-Chat 2016-01 2023-08 91
+Colossal-LLaMA-2-7B-Base 2016-01 2023-09 92
+LLaMA2-7B 2016-01 2023-02 85
+Tab.
+18 shows an analysis covering the period from 2015 to 2024, with the year 2019 serving as a pivotal midpoint, a distinct pattern emerges from the performance data of various language models: all models reached their peak accuracy before 2020.
+This observation points to a notable trend toward "Nostalgia", where models demonstrate optimal performance on data from the earlier part of the timeline analyzed.
+This tendency highlights the models’ alignment with the informational characteristics prevalent prior to the latter half of the observed period, suggesting their training data may disproportionately reflect earlier times.
+The peak accuracy periods analyzed are based on the average accuracy of all questions within three-month intervals, to provide a stable and robust assessment of each model’s performance.
+Appendix N Prediction Correlation over Time
+The graphical representations in our analysis plot the changes in correlation coefficients over time.
+Each line in the graph represents the changing relationship between a category’s average adjusted performance and a specific benchmark, highlighting periods of stronger or weaker association.
 Figure 7: Science and Technology Items - All Model Spearman Smoothed

### v3

@@
-Is Your LLM Outdated? Evaluating LLMs at Temporal Generalization
+Is Your LLM Outdated? A Deep Look at Temporal Generalization
 Abstract
-The rapid advancement of Large Language Models (LLMs) highlights the urgent need for evolving evaluation methodologies that keep pace with improvements in language comprehension and information processing.
-However, traditional benchmarks, which are often static, fail to capture the continually changing information landscape, leading to a disparity between the perceived and actual effectiveness of LLMs in ever-changing real-world scenarios.
-Our study examines temporal generalization, which includes the ability to understand, predict, and generate text relevant to past, present, and future contexts, revealing significant temporal biases in LLMs.
-We propose an evaluation framework, for dynamically generating benchmarks from recent real-world predictions.
-Experiments demonstrate that LLMs struggle with temporal generalization, showing performance decline over time.
-These findings highlight the necessity for improved training and updating processes to enhance adaptability and reduce biases.
-Our code, dataset and benchmark are available at https://github.com/FreedomIntelligence/FreshBench.
+The rapid advancement of Large Language Models (LLMs) has led to the development of benchmarks that consider temporal dynamics, however, there remains a gap in understanding how well these models can generalize across temporal contexts due to the inherent dynamic nature of language and information.
+This paper introduces the concept of temporal generalization in LLMs, including bias in past and future generalizations.
+Then we introduce FreshBench, a new evaluation framework that employs fresh text and event prediction for assessing LLMs’ temporal adaptability, ensuring the evaluation process free from data leakage and subjective bias.
+The experiment shows significant temporal biases and a decline in performance over time.
+Our findings reveal that powerful models, while initially superior, tend to decline more rapidly in future generalization.
+Additionally, powerful open-source models demonstrate better long-term adaptability compared to their closed-source counterparts.
+Our code is available at https://github.com/FreedomIntelligence/FreshBench
 1 Introduction
-Proprietary large language models (LLMs) such as GPT-4 [1], Bard [2], Claude [3], and PaLM [4] has been paralleled by efforts in the open-source community to democratize LLMs.
-Meanwhile, the rapid push for LLM democratization [5] emphasizes the necessity for dynamic and robust benchmarks that accurately reflect their evolving capabilities.
+Proprietary large language models (LLMs)  OpenAI (2023); GoogleAI (2023); Anthropic (2023); Anil et al.
+(2023) has been paralleled by efforts in the open-source community to democratize LLMs.
+Meanwhile, the rapid advancement of LLM democratization Touvron et al.
+(2023) emphasizes the necessity for dynamic and robust benchmarks that accurately reflect their evolving capabilities.
 Traditional benchmarks, broadly classified into knowledge-based assessments and open-dialogue evaluations, each present unique challenges.
-Knowledge-based assessments, exemplified by MMLU [6] and C-Eval [7], are prone to data manipulation, raising concerns about their real-world applicability and the risk of data leakage.
-Open-dialogue evaluations like MT-Bench [8] and Alpaca-Eval [9], rely on subjective human or model-based judgments, making their outcomes susceptible to the influence of dialogue structure, and thereby, potentially compromising assessment validity.
-Given the limitations of traditional methods, researchers are exploring new ways to measure model performance more comprehensively.
-Recently, several methods have been proposed to evaluate models at different times and in various contexts[10, 11, 12].
-TimeQA[13] evaluates a model’s ability to handle temporal evolution, while Situated QA[14] focuses on a model’s performance in specific contexts.
-StreamingQA[15] introduces the concept of continuous information flow, while RealTimeQA[11] and FreshQA[12] emphasize real-time and up-to-date information processing capabilities.
-However, existing methods may be prone to data leakage during their evaluation processes and typically focus on current or past data, lacking exploration of future event prediction capabilities.
-We believe that the evaluation of LLMs should be conducted in scenarios that are resistant to hacking manipulation, with results that are objective and free from evaluator biases [16].
+Knowledge-based assessments, exemplified by MMLU Hendrycks et al.
+(2021) and C-Eval Huang et al.
+(2023), are prone to data manipulation, raising concerns about their real-world applicability and the risk of data leakage Wei et al.
+(2023); Golchin and Surdeanu (2024).
+Open-dialogue evaluations like MT-Bench Zheng et al.
+(2023) and Alpaca-Eval Li et al.
+(2023a), rely on subjective human or model-based judgments, making their outcomes susceptible to the influence of dialogue structure, and thereby, potentially compromising assessment validity.
+Given the limitations of traditional methods, researchers are increasingly developing new approaches to evaluate models in diverse and dynamic contexts.
+For instance, recent efforts have introduced methods that assess models’ abilities to handle temporal evolution, contextual understanding, and real-time information updates Fatemi et al.
+(2024); Kasai et al.
+(2024); Vu et al.
+(2023); Chen et al.
+(2021); Zhang and Choi (2021); Liška et al.
+(2022).
+While methods like RealTimeQA and FreshQA focus on continuously updating knowledge for real-time processing Kasai et al.
+(2024); Vu et al.
+(2023), others like LatestEval Li et al.
+(2023b) emphasize recent text-based comprehension.
+In contrast, our work shifts the focus to evaluating models’ world understanding through prediction, using a fact-based approach that minimizes bias and ensures objectivity in assessment.
+This method enhances robustness against data contamination and focuses on the model’s capacity for accurate, real-world event prediction.
+We believe that the evaluation of LLMs should be conducted in scenarios that are resistant to hacking manipulation, with results that are objective and free from evaluator biases Chen et al.
+(2024).
 Additionally, time as a dynamic dimension, can also ensure robust and relevant evaluations.
 This paper explores an alternative way to evaluate large language models, i.e., on temporal generalization where time is considered as an important aspect.
-By doing so, the inaccessibility of future content would largely reduce the data contamination issue.
-Moreover, a retrospective perspective on time could be eneficial for temporal bias analysis for LLMs.
-To achieve this objective, we explore two specific scenarios designed to measure model capabilities without data leakage and to assess model performance on tasks relevant to both past and future contexts.
-The first scenario involved generating text from recent sources, such as arxiv papers, news articles, or new books, and Wikipedia.
-We expected the models to generalize to these diverse, newly generated texts with high language likelihood, thereby supporting the concept of "intelligence as compression" in a manner free from data leakage.
-The second scenario focused on predicting future events, suggesting that a superior model should be able to understand the present context, incorporate world knowledge, and generalize to future scenarios.
-In summary, the evaluation at temporal generalization offers two main benefits: it eliminates the possibility of data leakage through temporal boundaries, and its results can be objectively assessed without reliance on subjective interpretations.
+This approach significantly reduces the issue of data contamination by limiting access to future content.
+Moreover, a retrospective perspective on time could be beneficial for temporal bias analysis for LLMs.
+To achieve this goal, we design two scenarios.
+The first is to have the model generate text based on recent data such as arXiv papers, news articles, new books, and Wikipedia content.
+An excellent model should be able to generalize these new texts with high language likelihood without data leakage.
+The second scenario focuses on future event prediction.
+An excellent model should be able to understand the current context, integrate world knowledge, and effectively generalize to future scenarios.
+In these two scenarios, temporal generalization evaluation has two main advantages: eliminating the possibility of data leakage through temporal boundaries and enabling objective evaluation results without relying on subjective interpretations.
 Our analysis indicates that using validation loss of new text to evaluate models is mostly effective.
 However, we observe that while performance on language likelihood is a strong indicator of a model’s linguistic capabilities, it does not always align with specific abilities measured by existing benchmarks, which may represent discrete knowledge or other capacities that are important but not directly related to language modeling.
-Our contributions are as follows: 1) We thoroughly define and quantify temporal generalization and bias, establishing a basis for understanding and assessing the capabilities of LLMs over time.
-2) We proposed an benchmark Freshbench to test the capabilities of LLMs in predicting future events.
+Our contributions are as follows: 1) We define and quantify temporal generalization and bias, establishing a basis for understanding and assessing the capabilities of LLMs over time.
+2) We proposed a benchmark Freshbench to test the capabilities of LLMs in predicting future events.
 This benchmark is dynamically updated to reflect the latest data, ensuring that our evaluations remain relevant and accurate in the face of rapidly evolving data environments.
-It experimentally assesses performance of existing LLMs across various timeframes, providing insights into their historical and ongoing performance stability and adaptability.
+It experimentally assesses the performance of existing LLMs across various timeframes, providing insights into their historical and ongoing performance stability and adaptability.
 2 Conceptualization of Temporal Generalization
-Figure 1: Schematic figure of temporal generalization.
-In this section, we will conceptualize Temporal Generalization in Sec.
-2.1 and further measure it in 2.2.
-From a retrospective perspective in the Temporal Generalization, we also conceptualize Temporal Biases in Sec.
-2.3.
-In general, prior to the cutoff time, a LLM might exhibit a Nostalgia bias that is skewed toward past dates (i.e., aligning better with historical data) or a Neophilia bias that is skewed toward future dates ( favoring more recent information).
-However, after the cutoff time, the model’s understanding of future trends declines, leading to temporal degeneration.
+In this section, we will first conceptualize Temporal Generalization, then Temporal Biases and Temporal Degeneration in Sec.
+2.1.
+We will discuss ways to measure /textitTemporal Biases and Degeneration in 2.2.
 2.1 Temporal Generalization
-LLMs face the ongoing challenge of remaining applicability as new text evolves.
+LLMs face the ongoing challenge of remaining applicable as new text evolves.
 Consequently, it is crucial for these models to not only understand and generate text based on seen (e.g., past) data but also to anticipate and adapt to future data.
 This inspires temporal generalization as below.
+Figure 1: Schematic figure of temporal generalization.
+In general, prior to the cutoff time, an LLM might exhibit a Nostalgia bias that is skewed toward past dates (i.e., aligning better with historical data) or a Neophilia bias that is skewed toward future dates ( favoring more recent information).
+However, after the cutoff time, the model’s understanding of future trends declines, leading to temporal degeneration.
 Definition 2.1.
-Temporal Generalization refers to the LLMs’ ability to accurately understand, predict, and generate text relevant to past, present, and future contexts.
+Temporal Generalization refers to the LLMs’ ability to align with the contexts of the past, present, and future.
 This requires the models to integrate and apply knowledge from historical and current data to adapt to the evolution of language and emerging trends.
@@
 LLMs must maintain a delicate balance between leveraging historical data to understand long-standing language patterns and adapting to new expressions, terminology, and topics as they emerge.
+2.1.1 For Models: Define Past, Present and Future
+Figure 2: Schematic Diagram of 3 Periods of Models
+To distinguish the past and the future, we define the cutoff time as the release date of the model.
+For instance, if an LLM is released in January 2024, then all events or information before January 2024 are considered part of the past, while everything after January 2024 is considered the future.
+Although the data after the cutoff time is not strictly future data, this approach ensures no data leakage for the model.
+As is shown in Figure 2, the period near the cutoff time can be referred to as the present.
+We evaluate the model’s prediction accuracy across these time periods, denoted as $acc_{\text{past}}$, $acc_{\text{future}}$, and $acc_{\text{present}}$, respectively.
+2.1.2 Define Temporal Bias and Temporal Degeneration
+With the three time periods defined, we can define Temporal Bias and Temporal Degeneration for models.
+Definition 2.2.
+Temporal Bias In this paper, we define temporal bias as the preferences of LLMs over time.
+[^1] Temporal bias are multifaceted in previous literature.
+For instance, can refer to the assumption of an incorrect sequence of events that misleads our reasoning about causality Dorsey (2021) or user preferences over time in point-of-interest recommendations Rahmani et al.
+(2022).
+Recently, biases have raised more challenges in pre-trained language models Yogarajan et al.
+(2023), particularly focusing on under-represented societies.
+Sarkar and Vafa (2024) examines lookahead bias in pre-trained language models, providing insights into how these biases affect model performance.
+There exist two biases, namely Nostalgia Bias and Neophilia Bias, which reflect two contrasting tendencies in how these models might predict or generate text about the past, present, or future.
+A visual representation of these biases is provided in Fig.
+1, where the trends associated with Nostalgia and Neophilia biases are plotted over time.
+Nostalgia Bias refers to the over-reliance on historical data when generating text.
+This bias causes models to reflect perspectives, styles, or information prevalent in past training data, leading to outputs that may not account for recent developments.
+Essentially, Nostalgia Bias can make a model seem ‘stuck in the past’, limiting its ability to accurately represent the present or predict the future.
+Neophilia Bias implies a model’s overemphasis on novelty and future trends, potentially at the expense of established information.
+This bias could result in predictions that are overly speculative and not well-grounded in historical data or current realities.
+Though less common due to LLM training on historical datasets, Neophilia Bias could be perceived from prompt structuring or interpretation of outputs that prioritize futuristic concepts over historical accuracy.
+Definition 2.3.
+Temporal Degeneration refers to the decline in a model’s performance in the future, that is, after it is released.
+In conclusion, the illustration for the above concepts is in Figure 1.
 2.2 Assessing Temporal Generalization
-Based on definition in Sec. 2.1, we turn to assess temporal generalization through the lens of 1) I: compression intelligence in Sec. 2.2.1 or II): prognostication prediction in Sec. 2.2.2.
-2.2.1 Case I: Compression Intelligence over Time
+Based on definition in Sec. 2.1, we turn to assess temporal generalization through the lens of 1) I: compression intelligence in Sec. 2.2.1 and II): future event prediction in Sec. 2.2.2.
+2.2.1 Case I: Compression Intelligence
 In the assessment of compression intelligence, metrics such as perplexity (PPL) and bits per character (BPC) offer insights into model performance.
@@
 Unlike traditional BPC which is specific to character-level language models, our adapted Bits Per Character (BPC) metric offers a comparison between models by normalizing the log likelihood based on character count, irrespective of the tokenization method, allowing for a more equitable assessment of model performance across different languages and tokenization schemes, the BPC is calculated as follows:
-$\small BPC(T)=\frac{-\sum_{i=1}^{N}\log p(w_{i}|w_{1},\ldots,w_{i-1})}{\text{len-utf-8}(T)}$ (1)
+$\!BPC(T)\!=\!\frac{-\sum_{i=1}^{N}\log p(w_{i}|w_{1},\ldots,w_{i-1})}{\text{len-utf-8}(T)}$ (1)
 […]
-2.2.2 Case II: Prognostication Prediction
-LLMs could performs prognostication prediction, as seen in [18] which employed GPT-3 in a few-shot setting to tackle binary questions from the Metaculus platform.
-Further, [19] enriched the discourse by introducing a comprehensive dataset that includes forecasting questions alongside a news corpus.
+2.2.2 Case II: Future Event Prediction
+LLMs could perform future event prediction, as seen in Bonde (2022) which employed GPT-3 in a few-shot setting to tackle binary questions from the Metaculus platform.
+Further, Zou et al.
+(2022) enriched the discourse by introducing a comprehensive dataset that includes forecasting questions alongside a news corpus.
 This approach enhances model training by incorporating varied and relevant data contexts.
-Moreover, [20] showcased how integrating news content can significantly improve the accuracy of predictions in real-time scenarios.
-Building on this foundation, our research seeks to explore the temporal generalization capabilities of LLMs: how well they adapt and remain accurate as new data emerges and contexts evolve?.
+Moreover, Halawi et al.
+(2024) showcased how integrating news content can significantly improve the accuracy of predictions in real-time scenarios.
+Building on this foundation, our research seeks to explore the temporal generalization capabilities of LLMs: how well they adapt and remain accurate as new data emerges and contexts evolve?
 As an example of using LLMs to predict political outcomes, we gather new factual data about future events, such as "What will President Biden’s approval rating be as of 7 June 2024, according to FiveThirtyEight?".
@@
 We define prediction accuracy, denoted by $Acc$, as follows:
-$\small Acc=\frac{N_{\text{correct}}}{N_{\text{total}}}$ (2)
+$Acc=\frac{N_{\text{correct}}}{N_{\text{total}}}$ (2)
 where $N_{\text{correct}}$ is the number of correct predictions among $N_{\text{total}}$ future prediction questions.
@@
 This approach also allows continuous evaluation with the latest data, ensuring the assessment reflects the model’s current capabilities and the evolving information landscape.
-2.3 Temporal Bias in a Retrospective View
-Temporal biases are multifaceted in previous literature.
-For instance, temporal bias can refer to the assumption of an incorrect sequence of events that misleads our reasoning about causality [21] or user preferences over time in point-of-interest recommendations [22].
-Recently, biases have raised more challenges in pre-trained language models [23], particularly focusing on under-represented societies.
-[24] examines lookahead bias in pre-trained language models, providing insights into how these biases affect model performance.
-In the context of LLMs, temporal bias refers to the preferences of LLMs over time.
-There exists two biases, namely Nostalgia Bias (see Definition 2.2) and Neophilia Bias (see Definition 2.3), which reflect two contrasting tendencies in how these models might predict or generate text about the past, present, or future.
-A visual representation of these biases is provided in Fig.
-1, where the trends associated with Nostalgia and Neophilia biases are plotted over time.
-Definition 2.2.
-Nostalgia Bias refers to the over-reliance on historical data when generating text.
-This bias causes models to reflect perspectives, styles, or information prevalent in past training data, leading to outputs that may not account for recent developments.
-Essentially, Nostalgia Bias can make a model seem ‘stuck in the past’, limiting its ability to accurately represent the present or predict the future.
-Definition 2.3.
-Neophilia Bias implies a model’s overemphasis on novelty and future trends, potentially at the expense of established information.
-This bias could result in predictions that are overly speculative and not well-grounded in historical data or current realities.
-Though less common due to LLM training on historical datasets, Neophilia Bias could be perceived from prompt structuring or interpretation of outputs that prioritize futuristic concepts over historical accuracy.
-Challenges in Temporal Generalization Predicting future text with LLMs while navigating Nostalgia Bias and Neophilia Bias presents distinct challenges.
-Nostalgia Bias can lead to outdated predictions, while overcorrection towards Neophilia Bias might result in speculative predictions disconnected from past trajectories.
-Balancing these biases is crucial to leverage historical and contemporary data effectively, ensuring future predictions are informed by past patterns and open to new possibilities.
-This requires careful adjustment of the model’s training data and integration of up-to-date information.
 3 Temporal Generalization through the Lens of Compression Intelligence
-Temporal generalization in language models evaluates their ability to maintain performance over time, even as data evolves.
 This section discusses how various models handle data from different periods, showcasing their performance stability or volatility.
+Key Findings
+  • In general, while the ranking of models varied slightly between datasets, the performance of individual models remained relatively stable over time within each dataset type, as can be seen in Figure 5 and 5.
+Here we made a discovery regarding the distinct characteristics of the Wikipedia and BBC datasets at 3.1.2.
+  • Another interesting observation was made in the arXiv dataset shown in Figure  5, where a sharp decrease in BPC occurred around March 2024.
+This timing coincides with the peak in ChatGPT usage, as reported by Exploding Topics [^2] https://explodingtopics.com/blog/chatgpt-users, which saw over 1.8 billion visits in that month.
+In essence, it appears that humans may be adapting their writing to more closely resemble LLM outputs, rather than LLMs significantly enhancing their ability to model human language.
+Figure 3: BPC on Wiki text
+Figure 4: BPC on BBC news text
+Figure 5: BPC on arXiv text
 3.1 Temporal Bias
 3.1.1 Protocol for Quantitative Measurement
-Our system aggregates data from various platforms, including Yahoo for financial news, BBC for political insights, Reddit for discussions, Wattpad for literature, Wikipedia for encyclopedic updates, arXiv for academic research, GitHub for software trends, and Quora for diverse questions and answers.
-Utilizing a Python-based crawling framework, detailed in our public repository, it’s designed for adaptability and efficient data collection, with Playwright addressing navigation and specific site challenges.
-We also did some basic data pre-processing, due to space limitation, details in Appendix B.
-An overview of the data volume and the average entry length from a single crawl is provided in Appendix C.
-Data sources of Fresh Corpora We selected a set of models due to their diversity and significance in advancing the capabilities of NLP systems.
-They are not only benchmarks of their respective sizes, ranging from 3 billion to 13 billion parameters, but also exemplify the rapid progression in NLP technology.
-The models originate from diverse sources, showcasing the collaborative and competitive spirit driving innovation in the industry.
-Experimental Settings To further delve into the model’s performance dynamics over varying text lengths, we truncate the input text to the first 5000 characters and compute the sum of loss.
-The loss values are then accumulated to derive their sum, providing insight into the model’s efficiency at different stages of text comprehension.
-Throughout this experiment, approximately 300 GPU hours on NVIDIA A800 were consumed.
+Data sources Our system aggregates data from various platforms, including Yahoo for financial news, BBC for political insights, Reddit for discussions, Wikipedia for encyclopedic updates, arXiv for academic research, GitHub for software trends, and Quora for diverse questions and answers.
+Utilizing a Python-based crawling framework, it’s designed for adaptability and efficient data collection, with Playwright addressing navigation and specific site challenges.
+We also did some basic data pre-processing, due to space limitation, details in Appendix F.
+An overview of the data volume and the average entry length from a single crawl is provided in Appendix G.
+Experiment settings are listed in Appendix A.
 3.1.2 Observation of Temporal Bias in Compression Intelligence
-To comprehensively explore the temporal generalization capabilities of models, we plot the BPC curve of fresh text over time.
+To comprehensively explore the temporal generalization capabilities of models, we plot the BPC curve over time.
 In text data before release, a negative slope indicates the presence of Neophilia Bias.
 Conversely, a positive slope would suggest that the model’s performance is decreasing, which shows Nostalgia Bias.
-We sampled arXiv papers from 2015 to 2024 and applied smoothing techniques to the data because the number of samples was not sufficient to completely stabilize the average language likelihood.
-Nonetheless, a slight downward trend in model BPC can be observed, which is noteworthy.
-What is more, models with initially higher BPC showed a larger percentage decrease in BPC (as seen in models like OPT-2.7B and PHI-2 in Figure  3).
-For models with lower initial BPC, the decrease trend is less pronounced.
-Taking arXiv text as an example, BPC decreases over time, LLMs exhibit a Neophilia bias rather than a Nostalgia bias, capturing the present more effectively than the past.
-Turning to BBC news text, the observed trends in BPC from 2020 to 2023, depicted in Fig.
-3, provide insight into how well these models process and understand news content over a span of several years.
-We note a pronounced increase in BPC across the observed period, which suggests a diminishing adaptability of the models to newer data, indicative of a Nostalgia bias.
-This trend may stem from the dynamic and evolving nature of news narratives.
-Detailed result for various sources are in E
-Figure 2: BPC on arXiv text from 2015.9 to 2024.1 (smoothed)
-Figure 3: BPC on BBC news text from 2020 to 2023, with three-month intervals (smoothed)
-3.2 Temporal Generalization
+ArXiv BPC has a huge decline after GPT is widely used, as mentioned in Key Findings.
+Models show stronger Nostalgia Bias when evaluated with BBC text than Wiki text.
+From the BPC charts, it’s not easy to spot differences between models.
+However, our calculations in Appendix I reveal variations in how well each model can adapt specific text over time.
+Difference of Text Feature An insightful reviewer helped us found the reason of the different patterns of Wiki an BBC text.
+New Wiki texts with recent temporal markers do not necessarily imply that the content is new in a temporal context.
+For instance, in 2017-01, there was a Wiki article titled "History of India".
+However, the information within it was not new and could potentially have been sourced from other materials.
+Conversely, BBC texts are news articles and are highly time-sensitive.
+To quantify the difference, we uniformly sampled approximately 500 data pieces from both the Wiki and BBC datasets (taking into account that news may also cover historical events) and asked gpt-4o to classify the content.
+The results are as follows:
+  • Wiki: Historical: 97.83%, Others 333”Others” refers to content not associated with specific historical events, focusing instead on general, educational, and ever-relevant information. :2.17%.
+  • BBC: Recent: 91.86%, Historical: 8.14%.
+So, it is reasonable that Wiki text has a flat curve in BPC.
+Claim the Value of Wiki Dataset Since the Wikipedia dataset is continuously updated, it is not easily subject to direct contamination.
+Although a substantial portion of Wiki pertains to old information, this content still represents valuable knowledge that we need.
+An ideal model should consistently perform well on such knowledge.
+(If a model shows a decline in performance on Wikipedia, it is likely that the model is overfitting.)
+3.2 Temporal Degeneration
 3.2.1 Protocol for Quantitative Measurement
 Definition 3.1.
-Temporal Bias Index (TBI) For each model in a certain period, we record $BPC_{i}$ at $n$ time points $t_{i}$, forming $n$ data points $(t_{i},p_{i})$, where $i=1,\ldots,n$. We then fit a linear model:
-$\small BPC_{i}=a*t_{i}+b$ (3)
-where Temporal Bias Index (TBI) is defined as the coefficient $a$. TBI is crucial for understanding the direction and rate of change in model performance over time.
+Temporal Bias Index (TBI) For each model in a certain period, we record $BPC_{i}$ at $n$ time points, forming $n$ data points $(i,BPC_{i})$, where $i=1,\ldots,n$.
+Here the time interval between $BPC_{i}$ and $BPC_{i+1}$ is set as two month.
+We then fit a linear model:
+$BPC_{i}=a*i+b$ (3)
+then Temporal Bias Index (TBI) is defined as the coefficient $a$. TBI is crucial for understanding the direction and rate of change in model performance over time.
 To quantify model temporal generalization, we computed the BPC for various models on the Wiki and BBC datasets and evaluated their performance using this index.
@@
 After that, we measured the changes in BPC at 3, 6, 9 and 12 months post-release relative to the base BPC; and assessed the overall trend in model performance using TBI.
-The specific results are detailed in Appendix D.
-Table 1: Correlation between Model Attributes and Their Performance Decay Across Various Tasks on Wikipedia and BBC Benchmarks
+The specific results are detailed in Appendix H.
+Table 1: Correlation between Model Attributes and Their TBI on Wiki and BBC text
 Corre Model Release MMLU Human- Social- STEM MMLU GSM8 Ar20c Hella- Trust- Wino- Long-
@@
 3.2.2 Observation
-Models perform stably on Wikipedia but show more variation on BBC.
-In the Wikipedia dataset, models exhibit stable performance, demonstrating their ability to handle well-structured and formal text.
-However, larger models in the Wikipedia dataset deteriorate faster (correlation coefficient of 0.504), likely due to overfitting.
-In contrast, the BBC dataset shows significant variations in model performance, indicating insufficient adaptability to dynamic and informal text.
-From 2020 to 2024, the overall BPC for BBC texts shows an increasing trend, indicating gradual performance deterioration.
-Individual models show considerable differences in performance on new data; for instance, the Yi, LLaMA2, and Baichuan2 series adapt relatively well to new texts, while Phi-1.5 and OPT series exhibit poorer generalization (see Appendix E).
+Temporal generation pattern is mostly determined by the nature of data Models perform stably on Wikipedia but show more variation on BBC.
+This reveals their ability to handle well-structured and formal text like Wiki and insufficient adaptability to dynamic text like BBC text.
+The overall BPC for BBC texts shows an increasing trend, indicating gradual performance degeneration.
+Individual models show considerable differences in performance on new data; for instance, Yi, and Baichuan2 series adapt relatively well to new texts, while Phi-1.5 and OPT series exhibit poorer generalization (see Appendix I).
 These findings indicate that in dynamic environments, different models have significantly varying adaptation capabilities.
-Positive correlation between model attributes and performance decay.
+Model performance positively correlates with performance decay.
 (see Tab.
@@
 There is a positive correlation between model attributes (such as model size and release date) and performance decay.
-A positive BPC slope indicates an increase in BPC over time, reflecting performance deterioration.
-Larger model sizes may lead to rapid performance decay, suggesting a trade-off between capability and flexibility.
-In dynamic environments (such as the BBC dataset), the impact of model size on performance decay is smaller.
+A positive BPC slope indicates an increase in BPC over time, reflecting performance degeneration.
+In the Wiki dataset, larger models deteriorate faster (correlation coefficient of 0.504), suggesting a trade-off between capability and flexibility.
+In more dynamic environments (such as the BBC dataset), the impact of model size on performance decay is smaller.
 Additionally, newer models show higher correlations, meaning that despite their advanced capabilities, they may deteriorate faster.
@@
 As machine learning applications grow in complexity and scale, temporal generalization will become increasingly critical.
-4 Temporal Generalization through the Lens of Prognostication Prediction
-Advantages of Prognostication Prediction over Compression Intelligence Compression Intelligence is limited by their inapplicability to closed-source models, where key indicators like logits or losses are inaccessible.
+4 Temporal Generalization through the Lens of Future Event Prediction
+Advantages of Future Event Prediction over Compression Intelligence Compression Intelligence is limited by its inapplicability to closed-source models, where key indicators like logits or losses are inaccessible.
 This exclusion narrows comparative analysis and hinders a comprehensive understanding of the language technology landscape.
@@
 This motivated us to research temporal generalization and bias from the perspective of future prediction.
-In this section, we investigate tempera generalization (in Sec.
-4.1) and biases (Sec.
-4.2) through the Lens of Prognostication Prediction.
+In this section, we investigate temporal generalization (in Sec. 4.1) and biases (Sec. 4.2) through the Lens of Future Event Prediction.
+Key Findings
+  • Models prefer older knowledge, with non-uniform knowledge distribution skewed towards the past. This poses a risk as users expect up-to-date knowledge close to the model’s release date.
+  • Within model series such as Claude,Gemini, and GPT, newer versions are becoming more powerful but also experience more rapid performance decline.
+Miniature models significantly underperform compared to their full -sized counterparts in temporal tasks.
+  • Newer models generally exhibit stronger performance. Closed-source models tend to perform slightly better at the beginning, yet open-source models demonstrate superior generalization capabilities.
 4.1 Temporal Biases
 4.1.1 Protocol for Quantitative Measurement: a Hypothesis Testing Method
-To distinguish the past and the future, we take the border as the release date of the model.
-For example, if a LLM is released at Jan.
-2024 (also denoted as ‘cutoff time’ in this work), the time after Jan.
-2024 would be the future.
-We could collect question prediction tasks in different periods.
-Based on different periods, We compute the prediction accuracy in these different periods, denoted as $acc_{\text{past}}$ and $acc_{\text{future}}$, where accuracy of the present (somewhere between the past and future) the is denoted as $acc_{\text{present}}$.
+We discussed generally what is past, present, and future for a model in Sec 2.1.1.
+Now, let’s clarify the context of the prediction problem.
+For each prediction task, there is an initiation time (when the question is posed) and a close time (when the answer is revealed).
+If a question’s closure time precedes the model’s release date (for instance, by more than three months), then that question is considered part of the past for that model.
+The present refers to questions whose close time is near the release time of the model, capturing the period around when the model was launched.
 To investigate the two mentioned temporal biases in Sec.
-2.3, we propose two hypothesis testings as follows.
+2.2, we propose two hypothesis testings as follows.
 These tests help practitioners make informed decisions despite the randomness inherent in the question datasets.
@@
 Neophilia Bias:
-$\small T_{1}:\ H_{0}:\mathbf{E}[acc_{\text{past}}]=\mathbf{E}[acc_{\text{present}}]\ \text{versus}\ H_{1}:\mathbf{E}[acc_{\text{past}}]<\mathbf{E}[acc_{\text{present}}],\$ (4)
+$T_{1}:$ $H_{0}:\mathbf{E}[acc_{\text{past}}]=\mathbf{E}[acc_{\text{present}}],\quad$ (4)
+$H_{1}:\mathbf{E}[acc_{\text{past}}]<\mathbf{E}[acc_{\text{present}}]$
 This hypothesis tests if there is a significant increase in model accuracy for current questions compared to past questions, suggesting a bias towards newer content.
@@
 Nostalgia Bias:
-$\small T_{2}:\ H_{0}^{\prime}:\mathbf{E}[acc_{\textrm{past}}]=\mathbf{E}[acc_{\text{present}}]\ \text{versus}\ H_{1}^{\prime}:\mathbf{E}[acc_{\text{past}}]>\mathbf{E}[acc_{\text{present}}],$ (5)
+$T_{2}:$ $H_{0}^{\prime}:\mathbf{E}[acc_{\textrm{past}}]=\mathbf{E}[acc_{\text{present}}],\quad$ (5)
+$H_{1}^{\prime}:\mathbf{E}[acc_{\text{past}}]>\mathbf{E}[acc_{\text{present}}]$
 Conversely, this hypothesis tests if there is a significant increase in model accuracy for past questions, indicating a preference for older or historical content.
@@
 The p-values for assessing the shifts in accuracy are computed using the standard normal cumulative distribution function $\Phi$ as follows:
-$\small\widehat{p}=1-\Phi\left(\frac{acc_{\text{1}}-acc_{\text{2}}}{\sqrt{\frac{acc_{\text{1}}(1-acc_{\text{1}})}{n_{1}}+\frac{acc_{\text{2}}(1-acc_{\text{2}})}{n_{2}}}}\right),$ (6)
+$\widehat{p}=1-\Phi\left(\frac{acc_{\text{1}}-acc_{\text{2}}}{\sqrt{\frac{acc_{\text{1}}(1-acc_{\text{1}})}{n_{1}}+\frac{acc_{\text{2}}(1-acc_{\text{2}})}{n_{2}}}}\right),$ (6)
 $acc_{\text{1}}$ and $acc_{\text{1}}$ could be either $acc_{\textrm{past}}$ or $acc_{\textrm{present}}$, which is based on which Hypothesis Testing is used.
 Note that the two Hypothesis Testings are symmetrical.
-Experimental setting We have sourced all questions from GoodJudgmentOpen [25] through web scraping and API downloads.
-The collected data includes initiation and conclusion dates, domains, questions, descriptions, and answers, which are ordered appropriately for further processing.
-Throughout this experiment, approximately 50 GPU hours on NVIDIA A800 were consumed.
-We collected 2532 questions from GoodJudgmentOpen  [25], earliest one was proposed on 2015-09-01 and the most recent one was closed on 2024-05-25.
-It is "Will the adopted EU Artificial Intelligence (AI) Act be published in the Official Journal of the European Union before 25 May 2024?" Most common tags are in Appendix F.
+Detailed experiment settings are listed in Appendix D.
 Table 2: Comparison of Model Performance in Different Pre-Release Periods (Compared to 0-20 months before Release), uses asterisks to denote significance levels: (p < 0.05), ††(p < 0.01), and ††† (p < 0.001), $\boldsymbol{\boldsymbol{\boldsymbol{\boldsymbol{\rightarrow}}}}$ refers to Neophilia Bias, $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$ refers to Nostalgia Bias.
@@
 4.1.2 Observation
-Neophilia Bias is prevalent closer to release, whereas Nostalgia Bias occurs earlier.
-As shown in Tab.
-2, among the biases, Nostalgia is the most common, followed by periods where neither bias is significant (shown as -), with Neophilia Bias being the least frequent.
-Phi-1.5 and Baichuan2-13B series show Nostalgia Bias in all three periods before release.
-Pythia-12B and Mistral-7B-v0.1 show Neophilia Bias in 20 to 40 months before release but show more significant Nostalgia in another two periods, which we attribute to random fluctuations rather than a consistent trend.
-OPT-2.7B is Neophilia overall, but this observation is likely due to its consistently near-random performance across various periods, rather than a true preference for novelty.
 Models tend to be stronger in earlier periods before release, with performance gradually declining over time.
@@
 This suggests that the models gradually degrade during the pre-release period, rather than experiencing a rapid decline around the cutoff period.
+Nostalgia Bias is prevalent.As shown in Tab.
+2, among the biases, Nostalgia is the most common, followed by periods where neither bias is significant (shown as -), with Neophilia Bias being the least frequent.
+Phi-1.5 and Baichuan2-13B series show Nostalgia Bias in all three periods before release.
+Mistral-7B-v0.1 show Neophilia Bias in 20 to 40 months before release but show more significant Nostalgia in another two periods, which we attribute to random fluctuations rather than a consistent trend.
 This observation reveals that models generally favor older knowledge and prefer information from more distant past periods.
@@
 This poses a potential risk in our utilization of these models because there is an expectation that the knowledge utilized should be as up-to-date as possible, ideally close to the model’s release date.
-4.2 Temporal Generalization
+4.2 Temporal Degeneration
 4.2.1 Protocol
-In the prospective view when especially using LLMs for prognostication prediction, there might be a temporal degeneration issue defined as below:
+We discussed generally Temporal Degeneration in 2.3, here we need to define specifically.
 Definition 4.1.
-Temporal Degeneration The capability of prognostication prediction might be degraded over time due to the larger generalization gap for longer time.
-To qualitatively measure Temporal Degeneration, we propose a new hypothesis testing.
+Temporal Degeneration The capability of future event prediction might be degraded over time due to the larger generalization gap brought by time.
+To qualitatively measure Temporal Degeneration, we propose a hypothesis testing like 4.1.1.
 Hypothesis Testing 3.
 Temporal Degeneration
-$\small T_{3}:$ $\ H_{0}:\mathbf{E}[acc_{\text{future}}]=\mathbf{E}[acc_{\text{present}}]\ \text{versus}\ H_{1}:\mathbf{E}[acc_{\text{future}}]<\mathbf{E}[acc_{\text{present}}],$ (7)
+$T_{3}:$ $H_{0}:\mathbf{E}[acc_{\text{future}}]=\mathbf{E}[acc_{\text{present}}],$ (7)
+$H_{1}:\mathbf{E}[acc_{\text{future}}]<\mathbf{E}[acc_{\text{present}}]$
 Similar to Sec 4.1.1, $acc_{\text{future}}$ represents the model accuracy on questions proposed after the model’s release.
@@
 Table 3: Accuracy performance of various models across different periods.
-The presence of asterisks (*) denotes statistically significant decreases in performance during specific periods (p < 0.05 for *, p < 0.01 for ††, and p < 0.001 for ††*), benchmarked against their accuracy before release.
+The presence of asterisks (*) denotes statistically significant decreases in performance during specific periods (p < 0.05 for *, p < 0.01 for **, and p < 0.001 for ***), benchmarked against their accuracy before release.
 Grey cells indicate the models have not been released in this period and represent the average scores of the models prior to their release.
-Model 23/1 -23/2 23/3 -23/4 23/5 -23/6 23/7 -23/8 23/9 -23/10 23/11 -23/12 24/1 -24/2 24/3 -24/4
-OPT-13B 0.55 0.48 0.31** 0.39 0.41 0.52 0.37* 0.27**
-OPT-2.7B 0.17 0.20 0.17* 0.22 0.25 0.22 0.26 0.26
-LLaMA-7B 0.21* 0.17** 0.16** 0.22* 0.25 0.22** 0.22* 0.24*
-Pythia-12B 0.48 0.47 0.32** 0.43 0.48 0.39* 0.43 0.31**
-Falcon-rw-1B 0.22 0.16 0.14* 0.28 0.27 0.21 0.20 0.24
-Baichuan-7B-Chat 0.32 0.24* 0.30 0.23 0.19** 0.20* 0.34
-InternLM-Chat-7B 0.35 0.23** 0.33 0.27 0.29 0.30 0.31
-GPT-3.5-turbo-230613 0.49 0.29** 0.37* 0.45 0.44 0.37* 0.30**
-GPT-4-230613 0.6 0.39** 0.43* 0.41** 0.37** 0.48 0.40**
-LLaMA-2-13B 0.52 0.43 0.48 0.36** 0.39* 0.40*
-LLaMA-2-7B 0.27 0.26 0.27 0.22 0.26 0.34
-LLaMA-2-7B-Chat 0.45 0.30* 0.46 0.35** 0.37 0.38
-Baichuan2-7B-Chat 0.37 0.35 0.43 0.31 0.37 0.41
-Baichuan2-13B-Base 0.49 0.30** 0.46 0.42 0.28** 0.40*
-Baichuan2-13B-Chat 0.6 0.33** 0.43 0.41* 0.26** 0.27**
-Zhongjing-Base 0.38 0.30 0.33 0.26* 0.39
-Mistral-7B-v0.1 0.44 0.41 0.34* 0.28* 0.34*
-Phi-1.5 0.41 0.38 0.39 0.30 0.25**
-Colossal-LLaMA-2-7B-Base 0.42 0.43 0.37 0.28* 0.33*
-Qwen-14B-Chat 0.39 0.25** 0.23** 0.33 0.27**
-Qwen-7B 0.36 0.36 0.30 0.33 0.25**
-Skywork-13B-Base 0.37 0.39 0.29* 0.30 0.27*
-ChatGLM3-6B 0.42 0.38 0.33* 0.26** 0.34
-Zephyr-7B-beta 0.4 0.32 0.31* 0.28* 0.28*
-Yi-6B 0.29 0.27 0.16** 0.22 0.34
-Yi-6B-Chat 0.38 0.25** 0.39 0.35
-Qwen-1.8B 0.35 0.22** 0.37 0.30
-RWKV-v5-Eagle-7B 0.29 0.29 0.33 0.28
-GPT-4-231106 0.66 0.44** 0.39** 0.36**
-Phi-2 0.31 0.23* 0.22 0.30
-Command R+ 0.54 0.38**
-Mixtral-8x22B-Instruct-v0.1 0.56 0.40**
-Phi-3-mini-4k-instruct 0.38 0.26**
-Qwen1.5-110B-Chat 0.56 0.40**
+Model 23/1 -23/2 23/3 -23/4 23/5 -23/6 23/7 -23/8 23/9 -23/10 23/11 -23/12 24/1 -24/2 24/3 -24/4 24/5 -24/6 24/7 -24/8
+Baichuan-13B-Chat 0.52 0.26** 0.20** 0.36** 0.32** 0.33** 0.20** 0.23** 0.17**
+GPT-3.5-turbo-230613 0.49 0.29** 0.37* 0.45 0.44 0.37* 0.29** 0.28** 0.26**
+GPT-4-230613 0.6 0.39** 0.43* 0.41** 0.38** 0.48 0.40** 0.35** 0.36**
+Llama-2-13B 0.52 0.43 0.48 0.36** 0.39* 0.40* 0.28** 0.34**
+Llama-2-7B 0.27 0.26 0.27 0.22 0.26 0.33 0.18** 0.24
+LLaMA2-7B-Chat 0.45 0.30* 0.46 0.35** 0.37 0.38 0.27** 0.28**
+Baichuan2-13B-Base 0.49 0.30** 0.46 0.42* 0.28** 0.39* 0.27** 0.27**
+Baichuan2-13B-Chat 0.5 0.33** 0.43 0.41* 0.26** 0.29** 0.29** 0.33**
+Baichuan2-7B-Base 0.42 0.26** 0.30* 0.24** 0.33 0.33* 0.25** 0.22**
+Baichuan2-7B-Chat 0.37 0.35 0.43 0.31 0.37 0.40 0.23** 0.27*
+Colossal-LLaMA-2-7B-Base 0.42 0.43 0.37 0.28* 0.33* 0.25** 0.26**
+Mistral-7B-v0.1 0.44 0.41 0.35* 0.28* 0.34* 0.25** 0.24**
+Phi-1.5 0.41 0.38 0.38 0.30 0.26** 0.32* 0.27**
+Qwen-14B-Chat 0.39 0.25** 0.24** 0.33 0.27** 0.28** 0.19**
+Zephyr-7B-beta 0.4 0.32 0.32* 0.28* 0.28** 0.31* 0.19**
+Yi-6B 0.29 0.27 0.16** 0.22 0.33 0.21* 0.21*
+Gemini 0.43 0.20** 0.37 0.24** 0.21** 0.17**
+Qwen-1.8B 0.35 0.22** 0.37 0.29 0.20** 0.20**
+GPT-4-231106 0.66 0.45** 0.39** 0.36** 0.47** 0.43**
+Phi-2 0.31 0.23* 0.22 0.29 0.18** 0.15**
+Claude-3-opus-20240229 0.53 0.52 0.37** 0.47 0.36**
+Claude-3-sonnet-20240229 0.23 0.24 0.13** 0.10** 0.07**
+Command R+ 0.54 0.37** 0.43* 0.33**
+DeepSeek-V2-Chat 0.58 0.49** 0.46** 0.35**
+Mixtral-8x22B-Instruct-v0.1 0.56 0.39** 0.43** 0.35**
+Phi-3-mini-4k-Instruct 0.38 0.26** 0.24** 0.26**
+Qwen1.5-110B-Chat 0.56 0.40** 0.45* 0.36**
+Claude-3.5-Sonnet-20240620 0.63 0.48** 0.23**
+Gemini-1.5-Pro 0.5 0.47 0.35**
+Gemini-1.5-flash 0.39 0.23** 0.22*
+Qwen2-72B-Instruct 0.6 0.45** 0.40**
+LLaMA-3.1-405B-Instruct 0.5 0.41
+GPT-4o 0.62 0.39**
+GPT-4o-mini-2024-07-18 0.48 0.28**
 4.2.2 Observation
-Earlier models perform worse.
-As the latest scores in Table 3 shows, models released earlier perform worse.
-The poor performance may be due to two factors.
-One is the natural decay over time, as evidenced by the clear downturn in LLaMA-7B’s performance.
-Over time, earlier models fail to adapt to new data.
-The other is the initial lack of advanced training techniques, which may have resulted in models like OPT showing near-random performance from the outset, without significant changes.
-This is mainly because earlier models were trained with less advanced techniques, leading to poorer performance.
-Most models exhibit weak temporal generalization ability.
-Experimental results show that many models have significant fluctuations in accuracy percentages after release, indicating weak temporal generalization ability.
-Models with strong predictive capabilities fail to maintain their scores, even if only for a short period.
-For instance, GPT-4-230613 has demonstrated a significant decline in four out of six periods after release, while GPT-4-231106 has seen a significant decrease across all three observed periods.
-This variability and instability suggest that models have insufficient adaptive capabilities when dealing with new, unseen data.
-Some models exhibit near-random performance after the cutoff event, highlighting their limited understanding of the world and inability to generalize effectively from past information to future contexts.
-This shows that current models’ temporal generalization ability is significantly lacking and needs further improvement.
-More advanced models perform better in future predictions.
-.
-For instance, Qwen1.5-110B-Chat, Command R+(104B), Mixtral-8x22B-Instruct-v0.1, along with GPT-4, have achieved the highest scores during the period from March to April 2024.
-It is worth noting that certain versions of LLaMA-2 and Baichuan2 have also performed commendably.
-This observation may stem from a bias, as models like Qwen1.5-110B-Chat , Command R+(104B), Mixtral-8x22B-Instruct-v0.1 are the most recently released model, and a more robust evaluation awaits further data updates.
-The advanced training techniques employed by these models likely contribute to their enhanced accuracy and stability.
-As time advances and an increasing amount of data is accumulated, the models’ assessments are expected to become more precise and valuable.
-5 Conclusion
-We rigorously define and qualify temporal generalization and bias within LLMs, uncovering significant challenges in predicting and adapting to future contexts.
+Earlier models perform worse, while newer advanced models perform better in future predictions.
+As the latest scores in Table LABEL:grey_table show, earlier models, show a significant performance decline across multiple time periods.
+For example, LLaMA-7B’s accuracy dropped from 0.11 in March 2023 to 0.08 by March 2024, with no improvement in later evaluation periods.
+In contrast, newer models exhibit much stronger future prediction capabilities.
+Close-source models perform well initially but struggle with temporal generalization, while open-source models show better stability.
+As is shown in Appendix C, in both pre- and post-release scenarios, proprietary models such as GPT-4 and Claude-3.5 slightly outperform their open-source counterparts.
+However, when predicting outcomes for events known only post-release, all models exhibited substantial declines in performance.
+Our analysis highlights models with less than 31% decline in post-release accuracy in bold, illustrating relatively stable performance.
+Conversely, models that experienced declines greater than 39% are marked by underline, indicating substantial performance degeneration.
+Notably, models such as Command R+, Mixtral-8x22B-Instruct-v0.1, DeepSeek-V2-Chat, and LLaMA-3.1-405B-Instruct demonstrate slower declines in accuracy post-release.
+Conversely, models like Claude-3.5-Sonnet-20240620, GPT-4-231106, and GPT-4o show rapid declines in performance,
+The case studies of model iterations within the Claude series and the GPT models illustrate the dynamics of model evolution and generalization capabilities over time.
+Notably, the Claude-3-opus-20240229 model shows the least decline, indicating a potentially better tuning for broader applications or general datasets.
+This contrasts with the more significant declines seen in Claude-3-sonnet-20240229 and Claude-3.5-Sonnet-20240620, suggesting that these iterations may have undergone specific architectural tweaks, utilized different training data scopes, or implemented optimization strategies that did not generalize as effectively.
+Similarly, the progression from GPT-3.5-turbo-230613 to GPT-4-231106 shows an increasing trend in both pre-release accuracy and the magnitude of decline post-release, with the GPT-4-231106 model experiencing the largest drop (39.39%).
+This trend could indicate a push towards optimizing these models for higher initial accuracy, possibly at the expense of their ability to generalize effectively over time.
+Miniature versions of models show significant declines in temporal performance.
+The data reveals that miniature versions of models underperform compared to their full-sized counterparts in temporal tasks.
+For instance, Gemini.pro had an accuracy of 0.37 in March 2024, but this dropped sharply to 0.17 by July 2024, indicating a significant loss in precision for temporal tasks.
+Similarly, the mini version of GPT-4 showed considerably lower performance across multiple time periods, particularly in April 2024, where its accuracy was just 0.21, around 0.1 lower than the full-sized GPT-4 in the same period.
+This illustrates that while mini models improve inference efficiency, they exhibit poor generalization over time, with performance rapidly declining in long-term prediction tasks.
+5 Conclusion and Community Call to Action
+We rigorously define and quantify temporal generalization and bias within LLMs, uncovering significant challenges in predicting and adapting to future contexts.
 Our experiments reveal that LLMs often struggle with temporal biases, affecting reliability in dynamic scenarios requiring accurate forecasting.
@@
 The findings highlight shortcomings in existing benchmarks related to temporal shifts and underscore the urgent need for improvements in LLM training and updating processes to enhance adaptability and mitigate biases.
-Limitations
-While we strive to collect textual data from a variety of sources such as news articles, scientific papers, and forum discussions, the proportion and diversity of the texts may still fall short of fully reflecting the diversity and complexity of language.
-Moreover, such evaluations on text loss require access to the model’s logits, which is challenging for closed-source models or models accessed via APIs.
+Specifically, we urge the community to take note of this phenomenon: models tend to better comprehend earlier events.
+Cheng et al.
+(2024) suggests that models are most familiar with knowledge up to the year 2020.
+However, our findings indicate that the understanding of world events (even in cases where models can remember) lingers around 2015, with a trend towards even earlier years (beyond the time range of our test data).
+This observation is contrary to our initial hypothesis that models may overfit current events, which is significantly more concerning.
+It suggests that caution is needed, especially for those hoping to use model judgments directly.
+Acknowledgements
+This work was supported by the Shenzhen Science and Technology Program (JCYJ20220818103001002), Shenzhen Doctoral Startup Funding (RCBS20221008093330065), Tianyuan Fund for Mathematics of National Natural Science Foundation of China (NSFC) (12326608), Shenzhen Key Laboratory of Cross-Modal Cognitive Computing (grant number ZDSYS20230626091302006), and Shenzhen Stability Science Program 2023.
 Future Work
 Moving forward, it is crucial to refine LLM development and evaluation techniques to better handle evolving informational environments, ensuring their continued relevance and utility in various applications.
-Looking ahead, our agenda for future work includes several ambitious goals aimed at enhancing the versatility and depth of our evaluation framework. Key among these objectives are:
+Looking ahead, we suggest future work includes several ambitious goals aimed at enhancing the versatility and depth of our evaluation framework. Key among these objectives are:
+  • Exploring mitigation strategies, through we conducted preliminary investigations into prompt engineering strategies that could potentially mitigate temporal performance degradation, see Abstract B.
   • Integrating a wider array of models, particularly focusing on those based on the Transformer architecture, such as Griffin and Mamba models.
@@
 These initiatives represent our commitment to advancing the state of the art in language model evaluation, ensuring that our tools remain at the forefront of technological progress and innovation.
-Community Call to Action
-We urge the community to take note of the above phenomenon: models tend to have a better comprehension of earlier events.
-This observation supports the findings in  [26], which suggest that models are most familiar with knowledge up to 2020.
-This observation is contrary to our initial hypothesis that models may overfit current events, which is significantly more concerning.
-It suggests that caution is needed, especially for those hoping to use model judgments directly.
+Limitations
+While we strive to collect textual data from a variety of sources such as news articles, scientific papers, and forum discussions, the proportion and diversity of the texts may still fall short of fully reflecting the diversity and complexity of language.
+Also, the English-centric nature of the questions and answers limits the benchmark’s applicability to non-English speaking regions and global events.
+Moreover, such evaluations on text loss require access to the model’s logits, which is challenging for closed-source models or models accessed via APIs.
+Potential Risks
+Bias and Fairness Risks: When used to make predictions about future events, language models can inadvertently perpetuate or amplify biases present in their training data.
+This could lead to biased forecasts that reflect and potentially exacerbate existing societal prejudices.
+Such biased outputs may result in models that are unfair or discriminatory, which could harm individuals or groups and damage the reputation of the organizations involved.
 Ethics Statement
@@
 Licence
-5.1 Good Judgment Inc.
+Good Judgment Inc.
 For personal use of its services and content, a limited, revocable, non-exclusive license is granted.
 Without explicit written permission from Good Judgment Inc., this license restricts the collection, aggregation, copying, duplication, display, or derivative use of its content.
-5.2 arXiv
+arXiv
 The arXiv dataset is licensed under the Creative Commons Zero (CC0 1.0) license.
 This means it is in the public domain, and users can freely use, modify, and distribute the data without any restrictions.
-5.3 Wikipedia
+Wikipedia
 Wikipedia content is licensed under the Creative Commons Attribution-ShareAlike (CC BY-SA) license.
@@
 Other text will not be released.
-Acknowledgement
-This work was supported by the Shenzhen Science and Technology Program (JCYJ20220818103001002), Shenzhen Doctoral Startup Funding (RCBS20221008093330065), Tianyuan Fund for Mathematics of National Natural Science Foundation of China (NSFC) (12326608), Shenzhen Key Laboratory of Cross-Modal Cognitive Computing (grant number ZDSYS20230626091302006), and Shenzhen Stability Science Program 2023, Shenzhen Key Lab of Multi-Modal Cognitive Computing.
-Appendix A Testing Statistical Hypothesis
+Appendix A Related Work
+Given the limitations of traditional methods, researchers are exploring new ways to measure model performance more comprehensively.
+Recently, several methods have been proposed to evaluate models at different times and in various contexts Fatemi et al.
+(2024); Kasai et al.
+(2024); Vu et al.
+(2023).
+TimeQA Chen et al.
+(2021) evaluates a model’s ability to handle temporal evolution, while Situated QA Zhang and Choi (2021) focuses on a model’s performance in specific contexts.
+StreamingQA Liška et al.
+(2022) introduces the concept of continuous information flow, while RealTimeQA Kasai et al.
+(2024) and FreshQA Vu et al.
+(2023) emphasize real-time and up-to-date information processing capabilities.
+LatestEval  Li et al.
+(2023b) utilizes the most recent text to construct questions and answers, covering six most frequently queried categories including explanation, summary, reason, demonstrations, existence, and possible usage.
+However, it’s essential to clarify that RealTimeQA Kasai et al.
+(2024) and FreshQA Vu et al.
+(2023), which continuously update topics, are inclined towards knowledge.
+For example, RealTimeQA Kasai et al.
+(2024) has such a sample in their paper, "Which wildly popular show was recently greenlit for a new season?" (Answer: Squid Game).
+These methods are typically used for evaluating retrieval-augmented generation (rag) series models.
+On the other hand, methods like LatestEval  Li et al.
+(2023b), which construct questions using the most recent information, lean towards enhancing reading comprehension.
+In contrast to these approaches, our focus is on testing the model’s understanding of the world, as demonstrated through prediction.
+Our method of evaluation is based on facts and computation loss.
+This ensures that our evaluation is resistant to data contamination, objective, and free from biases in question construction and judgment.
+This approach allows us to examine the model’s ability to process and predict real-world events accurately and efficiently.
+Future-knowledge Prediction
+Previous research has focused on how to automatically predict real-world events.
+For instance, Zou et al.
+(2022) constructed a dataset of questions obtained from tournament predictions as an evaluation benchmark for automatic forecasting.
+However, this dataset may already be outdated because its knowledge is likely already embedded within current LLMs.
+Recently, studies have shown that LLMs can match human performance in the task of forecasting the future.
+For example, Schoenegger et al.
+(2024) demonstrates that the predictive performance of an ensemble of multiple LLMs is superior to that of a single LLM, and it even surpasses the performance of a group of human forecasters.
+Model selection and Experimental Setting We selected a set of models due to their diversity and significance in advancing the capabilities of NLP systems.
+They are not only benchmarks of their respective sizes, ranging from 1 billion to 40 billion parameters, but also exemplify the rapid progression in NLP technology.
+We set the context size to 2048, then compute the sum of the negative logarithms of each segment.
+We then divide the computed sum by the length of the text under UTF-8 encoding.
+Throughout this experiment, approximately 300 GPU hours on NVIDIA A800 were consumed.
+Appendix B Exploring Mitigation Strategies
+While our primary focus is temporal evaluation rather than solution development, we conducted preliminary investigations into prompt engineering strategies that could potentially mitigate temporal performance degradation.
+B.1 Try Different Prompt to Mitigate
+Motivated by observed safety constraints and refusal patterns in base models and the result of Pham and Cunningham (2024), we designed five specialized prompts to investigate how different framing strategies affect temporal prediction:
+Figure 6: Accuracy of LlaMA3.1-8B-Instruct with different prompts
+Figure 7: Accuracy of Qwen2.5-7B with different prompts
+Figure 8: Accuracy of Baichuan2-7B-Chat with different prompts
+Figure 9: Accuracy of Yi-6B with different prompts
+  • Base: Used direct incentivization ("tip of \$200") and explicit guessing encouragement to counter safety-driven refusals.
+  • Dream Framework (v1): Encourages speculative thinking by positioning the model as a visionary oracle.
+  • Time Travel (v2): Frames predictions as historical memories from a future perspective.
+  • Multiverse (v3): Leverages parallel universe observations to bypass temporal constraints.
+  • Precise Thinking (v4): Emphasizes structured temporal reasoning chains.
+  • Future Recall (v5): Simulates retrospective analysis from a future vantage point.
+See Figure 9, 9, 9 and 9.
+Specifically, different models respond variably to the designed prompts.
+For Yi-6B, there were scarcely any discernible differences in performance across the several prompts we employed.
+In contrast, for the Qwen2.5-7B model, all modified prompts generally outperformed the base prompt in most scenarios.
+Analyzing the results from Llama3.1-8B-Instruct, Qwen2.5-7B and Baichuan2-7B-Chat, we found that prompts v5(Future Recall) and v4(Precise Thinking) demonstrated relatively better effectiveness.
+The specific impacts vary from model to model, which is closely related to the model’s instruction-following capabilities.
+In addition, our exploratory analysis revealed significant performance variations across question categories, with the model achieving lowest accuracy on "Finance" (58.1%), "Business" (55.4%).
+and "Economic Indicators" (62.3%) categories, compared to stronger performance on "Foreign Policy", (81.2%), "Security and Conflict" (79.6%) and "Leader Entry/Exit" (83.7%) categories.
+It is plausible that fine-tuning models with analysis and prediction data (especially in these weak areas) could fully unlock their inner predictive potential.
+Appendix C Full Table of Decline in Prediction Accuracy
+See Table 4.
+Table 4: Comparison of Powerful Open-source and Closed-source LLMs
+Model Pre-release Accuracy First Post-release Accuracy Decline Percentage Decline
+Baichuan-13B-Chat 0.52 0.23 0.29 55.77%
+GPT-3.5-turbo-230613 0.49 0.33 0.16 32.65%
+GPT-4-230613 0.60 0.41 0.19 31.67%
+Llama-2-13B 0.52 0.45 0.07 12.50%
+Llama-2-7B 0.27 0.27 0.01 1.85%
+LLaMA2-7B-Chat 0.45 0.38 0.07 15.56%
+Baichuan2-13B-Base 0.49 0.38 0.11 22.45%
+Baichuan2-13B-Chat 0.50 0.38 0.12 24.00%
+Baichuan2-7B-Base 0.42 0.28 0.14 33.33%
+Baichuan2-7B-Chat 0.37 0.39 -0.02 -5.41%
+Colossal-LLaMA-2-7B-Base 0.42 0.40 0.02 4.76%
+Mistral-7B-v0.1 0.44 0.38 0.06 13.64%
+Phi-1.5 0.41 0.38 0.03 7.32%
+Qwen-14B-Chat 0.39 0.24 0.15 37.18%
+Zephyr-7B-beta 0.40 0.32 0.08 20.00%
+Gemini 0.43 0.29 0.14 33.72%
+Yi-6B 0.29 0.22 0.07 25.86%
+Qwen-1.8B 0.35 0.29 0.05 15.71%
+GPT-4-231106 0.66 0.42 0.24 36.36%
+Phi-2 0.31 0.23 0.08 27.42%
+Command R+ 0.54 0.40 0.14 25.93%
+Claude-3-opus-20240229 0.53 0.45 0.09 16.04%
+Claude-3-sonnet-20240229 0.23 0.18 0.05 19.57%
+DeepSeek-V2-Chat 0.58 0.47 0.10 18.10%
+Mixtral-8x22B-Instruct-v0.1 0.56 0.41 0.15 26.79%
+Phi-3-mini-4k-Instruct 0.38 0.25 0.13 34.21%
+Qwen1.5-110B-Chat 0.56 0.43 0.14 24.11%
+Claude-3.5-Sonnet-20240620 0.63 0.35 0.28 43.65%
+Gemini-1.5-Pro 0.50 0.41 0.09 18.00%
+Gemini-1.5-flash 0.39 0.23 0.17 42.31%
+Qwen2-72B-Instruct 0.60 0.43 0.17 29.17%
+LLaMA-3.1-405B-Instruct 0.50 0.41 0.09 18.00%
+GPT-4o 0.62 0.39 0.23 37.10%
+GPT-4o-mini-2024-07-18 0.48 0.28 0.20 41.67%
+Appendix D Experimental setting
+For this experiment, we sourced all questions from Good Judgment Open using web scraping.
+The data collected includes initiation and conclusion dates, domains, questions, descriptions, and answers, which were appropriately ordered for further processing.
+Approximately 50 GPU hours on NVIDIA A800 were consumed during this experiment.
+Up to 2024.10.1, we collected a total of 2769 questions from Good Judgment Open .
+The earliest question was proposed on 2015-09-01: "Will Congress pass a resolution disapproving the Joint Comprehensive Plan of Action?" The most recent question was closed on 2024-10-01: "What will be the 12-month percentage change in the US Consumer Price Index (CPI) for September 2024?" Question samples and most common tags are in Appendix J.
+Appendix E Testing Statistical Hypothesis
 Due to the inherent randomness in data, directly comparing outcomes of two experiments is insufficient to justify the superiority of one outcome over the other.
-To address this issue, practitioners often resort to hypothesis testing to make convincing decisions, like those demonstrated in [27].
-A board application of hypothesis testing lies in the A/B testing literature, which assesses the performance of various features of a product and performs post-experiment inferences, as introduced in [28].
-Appendix B Details of Data Collection
+To address this issue, practitioners often resort to hypothesis testing to make convincing decisions, like those demonstrated in Zhou et al.
+(2016).
+A board application of hypothesis testing lies in the A/B testing literature, which assesses the performance of various features of a product and performs post-experiment inferences, as introduced in Russac et al.
+(2021).
+Appendix F Details of Data Collection
 The data for our system is sourced from a diverse set of online platforms, each selected for its relevance to different data types and domains.
@@
 This structured sampling approach was designed to ensure a consistent and representative analysis of content over time, enabling an effective assessment of textual bias and likelihood trends.
-Appendix C Data Overview of One Crawl
+Table 5: Most common tags of the Gjo questions
+Topic Count
+Business 648
+Non-US Politics 453
+Security and Conflict 446
+Society 425
+US Politics 363
+Technology 359
+Finance 299
+Elections and Referenda 288
+Economic Indicators 273
+Foreign Policy 271
+US Policy 247
+Health 245
+Leader Entry/Exit 219
+Economic Policy 160
+Sports 148
+Entertainment 148
+Environment 80
+Open 16
+Appendix G Data Overview of One Crawl
 Tab.
@@
 This includes detailed counts from different academic archives like arXiv, various categories from Internet question and answer platforms like Reddit and Quora, and articles from well-known online platforms including Wiki and BBC.
-Table 4: Overview of text categories with their sources counts, and average lengths in one sample crawl.
-Category Source Count Length
-Academic STEM arXiv Mathematics 50 42850
-arXiv Computer Science 50 43336
-arXiv Physics 50 42580
-arXiv Statistics 50 42619
-arXiv Electrical Engineering and Systems Science 50 43027
-Academic Non-STEM arXiv Economics 50 42592
-arXiv Quantitative Finance 50 43969
-arXiv Quantitative Biology 50 42974
-Code GitHub 30 21849
-Internet QA Reddit 583 643
-Quora Health 20 1150
-Quora Movies 69 1349
-Quora Mathematics 70 1365
-Quora Technology 80 1674
-Internet article Wiki 220 4091
-Wattpad 38 6615
-BBC 45 4792
-Yahoo 41 5373
-Appendix D Detailed results of BPC generalization on Wiki and BBC dataset
+Appendix H Detailed results of BPC generalization on Wiki and BBC dataset
+See Table 6.
 Table 6: Percentage Change in BPC Following Release, measured by BBC news data from 2020 to 2024.3.
@@
 Phi-2 0.072 -0.037 - - - -0.037
-Appendix E TBI of Models across the arXiv, BBC, and Wiki datasets
+Appendix I TBI of Models across the arXiv, BBC, and Wiki datasets
+See Table 8.
 Table 8: Comprehensive list of models with their release dates and metrics for TBI (multiplied by 1000) across the arXiv, BBC, and Wiki datasets.
@@
 Phi-2 Dec 2023 -21.2 669.3 -44.7
-Appendix F Most common tags of the Gjo questions
-Table 8: Frequency of Topics
-Topic Count
-Business 648
-Non-US Politics 453
-Security and Conflict 446
-Society 425
-US Politics 363
-Technology 359
-Finance 299
-Elections and Referenda 288
-Economic Indicators 273
-Foreign Policy 271
-US Policy 247
-Health 245
-Leader Entry/Exit 219
-Economic Policy 160
-Sports 148
-Entertainment 148
-Environment 80
-Open 16
-Table 9: Most common tags of the Gjo questions
-Appendix G Correlation with Existing Benchmarks
-Following our previous chapter I, we continue to explore the correlation between prediction performance and existing benchmarks, using the same benchmarks as in I.
-We have categorized the tags of questions into several groups: Societal Issues Items, Political and Government Affairs, Economic and Business Items, and Science and Technology Items.
-Each group represents a significant area of interest.
-In our analysis, we focus on how the correlation between these categories and various benchmarks evolves over time.
-We use correlation coefficients to measure the strength of association between the performance scores of questions within each category and the benchmarks across different times.
-Notably, our analysis revealed that while the Economic and Business Items and Science and Technology Items on base models displayed relatively high initial correlations, there appears to be little to no relationship in Societal issues Items and Political and Government Affairs, we think such topics are influenced by a multitude of unpredictable factors such as social changes, policy adjustments, and public sentiment.
-These elements make accurate predictions particularly challenging, leading to the observed low correlation with performance metrics.
-Detailed figures are in Appendix N.
-Despite relatively high initial correlations in 2 question groups, there is a general lack of correlation between model predictions and existing benchmarks (mean correlation coefficient around zero).
-In addition, significant deviations occur in nearly half of the cases from early stages to subsequent periods.
-This suggests that the models’ understanding and alignment with benchmarks can vary substantially over time.
-So temporal generalization provides novel insights and draws attention to an aspect that has been often ignored by previous benchmarks.
-Appendix H Analysis of Language Likelihood Values and Benchmark Correlation Across Text and Model Groups
+Table 9: Overview of text categories with their sources counts, and average lengths in one sample crawl.
+Category Source Count Length
+Academic STEM arXiv Mathematics 50 42850
+arXiv Computer Science 50 43336
+arXiv Physics 50 42580
+arXiv Statistics 50 42619
+arXiv Electrical Engineering and Systems Science 50 43027
+Academic Non-STEM arXiv Economics 50 42592
+arXiv Quantitative Finance 50 43969
+arXiv Quantitative Biology 50 42974
+Code GitHub 30 21849
+Internet QA Reddit 583 643
+Quora Health 20 1150
+Quora Movies 69 1349
+Quora Mathematics 70 1365
+Quora Technology 80 1674
+Internet article Wiki 220 4091
+Wattpad 38 6615
+BBC 45 4792
+Yahoo 41 5373
+Appendix J Gjo data samples and most common tags of the Gjo questions
+Gjo data samples see Table 5 and Gjo questions see Table 10.
+Table 10: Sample of Questions and Their Respective Answers (Correct Answer in Bold): The percentages represent the proportion of forecasters’ votes on a prediction website.
+Question Start Time End Time Possible Answers
+What will be the 12-month percentage change in the US Consumer Price Index (CPI) for September 2024?
+May 31, 2024 05:00PM UTC Oct 01, 2024 07:01AM UTC Up by less than 1.200% or down: 0% Up by at least 1.200%, but less than 1.800%: 0% Up by at least 1.800%, but less than 2.400%: 53% Up by at least 2.400%, but less than 3.000%: 44% Up by at least 3.000%, but less than 3.600%: 3% Up at least 3.600%, but less than 4.200%: 0% Up by 4.200% or more: 0%
+If Israeli Defense Forces (IDF) ground forces invade the Gaza Strip before 7 November 2023, when will Israel publicly announce or acknowledge that IDF ground forces have left the Gaza Strip?
+Oct 17, 2023 05:00PM UTC Oct 10, 2024 07:01AM UTC Before 30 November 2023: 0% Between 30 November 2023 and 13 January 2024: 0% Between 14 January 2024 and 13 March 2024: 0% Between 14 March 2024 and 11 June 2024: 0% Between 12 June 2024 and 9 October 2024: 2% Not before 10 October 2024: 96% IDF ground forces will not invade the Gaza Strip before 7 November 2023: 2%
+What will Kamala Harris’ favorability rating be as of 3 October 2024, according to FiveThirtyEight?
+Sep 17, 2024 08:00AM UTC Oct 03, 2024 07:01AM UTC Lower than 42.0%: 1% At least 42.0%, but less than 44.0%: 4% At least 44.0%, but less than 46.0%: 7% At least 46.0%, but less than 48.0%: 70% At least 48.0%, but less than 50.0%: 15% At least 50.0%, but less than 52.0%: 2% At least 52.0%, but less than 54.0%: 1% 54.0% or higher: 0%
+Appendix K Analysis of Language Likelihood Values and Benchmark Correlation Across Text and Model Groups
 (a) STEM: base models
@@
 12 we can conclude that the scale of models being observed does not significantly alter outcomes, suggesting that performance correlations across benchmarks are generally consistent regardless of model size.
-Appendix I Correlation with Existing Benchmarks
+Appendix L Correlation with Existing Benchmarks
 The Selected Existing Benchmarks
@@
 Benchmark Description
-HellaSwag Tests an AI’s common sense reasoning by requiring it to complete sentences or narratives that reflect everyday logic, highlighting the model’s ability to predict logical continuations. [29]
-MMLU Evaluates an AI’s breadth and depth of understanding across a diverse range of subjects, from science to literature, showcasing its versatility and comprehensive grasp of human language. [30]
+HellaSwag Tests an AI’s common sense reasoning by requiring it to complete sentences or narratives that reflect everyday logic, highlighting the model’s ability to predict logical continuations.
+Zellers et al.
+(2019)
+MMLU Evaluates an AI’s breadth and depth of understanding across a diverse range of subjects, from science to literature, showcasing its versatility and comprehensive grasp of human language.
+Hendrycks et al.
+(2021)
 TruthfulQA Focuses on the model’s capacity to provide honest and accurate responses, particularly to questions where the truth may not be intuitive, testing its commitment to truthfulness and its proficiency in avoiding misinformation.
-[31]
-Winogrande Assesses common sense reasoning and natural language understanding through sentence completion tasks, measuring an AI’s ability to apply common sense knowledge in language processing. [32]
-GSM8K Challenges AI models with elementary-level math problems, testing their mathematical reasoning capabilities and their understanding and application of basic math concepts. [33]
+Lin et al.
+(2022)
+Winogrande Assesses common sense reasoning and natural language understanding through sentence completion tasks, measuring an AI’s ability to apply common sense knowledge in language processing.
+Sakaguchi et al.
+(2019)
+GSM8K Challenges AI models with elementary-level math problems, testing their mathematical reasoning capabilities and their understanding and application of basic math concepts. Cobbe et al. (2021)
 Base vs. Chat Models
@@
 Chat Models 0.01 -0.07 -0.16 -0.39 -0.53 -0.58 -0.61 -0.63 -0.65 -0.65 -0.67 -0.67 -0.68
-On the Genres of Corpora
-As can be seen in Appendix  H, whether all models are observed or only those of a specific scale does not make a big difference.
-Comparisons across different model sizes show that the differences are not substantial, indicating consistency in performance correlations across benchmarks.
-Further analysis from Tab.
-13 within the MMLU benchmark sections reveals interesting patterns.
-While we assume models to be generally excel in domains closely aligned with their training data, this correlation is not strong.
-Specifically, STEM scores do not show strong correlation with STEM-related texts.
-Moreover, academic papers, even in Social Sciences and Humanities, exhibit stronger correlations than Internet data, which includes forums and news.
-In this context, code-related data show similarities in performance correlation with academic texts, suggesting a closer alignment than previously anticipated.
-Similarly, Internet data, including forums and news, also shows comparative closeness in performance, indicating nuanced relationships across different types of content.
-Table 13: Correlation Coefficients between BPC in specific areas and Domain-Specific Performance
-Benchmark Academia STEM Academia non-STEM Code Internet Article Internet Chat
-MMLU:STEM -0.17 -0.19 -0.11 -0.01 -0.01
-MMLU:Humanities -0.54 -0.53 -0.51 -0.29 -0.39
-MMLU:Social Sciences -0.49 -0.42 -0.46 -0.25 -0.31
-Additionally, compared to chat forum texts, more formal academic texts demonstrate a stronger correlation with benchmarks.
-The genre of text has a minor impact on the sign of the correlation coefficient, which is relatively small compared to the differences across model types.
-Appendix J Function Fitting Strategies
-We considered several fitting strategies, each with its unique mathematical approach to model decay:
-Linear Decay
-Linear decay is represented by a straight line, suggesting a constant rate of decay over time. The model is given by the equation $y=ax+b$.
-Exponential Decay
-Exponential decay is modeled by an exponential function, expressed as $y=y_{0}\cdot e^{-kt}$, where $y_{0}$ is the initial value, $t$ is time or another decaying factor, and $k$ is the decay constant that denotes the rate of decay.
-Power Law Decay
-Power law decay implies a polynomial relationship and is expressed as $y=at^{-k}$, where $a$ is a positive constant, $t$ is time or another related factor, and $k$ represents the exponent characterizing the rate of decay.
-Logarithmic Decay
-Logarithmic decay suggests a decrease that becomes more gradual over time, typically modeled as $y=a+b\log(x)$, where $a$ and $b$ are constants, and $x$ represents time or another related factor.
-The variance of the residuals for each model was calculated as follows:
-Table 14: Variance of Residuals for Different Models
-Model Variance
-Linear 2.346e-3
-Exponential 1.71e3
-Power Law 3.78e-3
-Logarithmic 3.33e-3
-J.0.1 Findings
-Our analysis found that the linear model provided the best fit, with the lowest variance among the residuals.
-This indicates a steady, unchanging rate of knowledge decay over time, as opposed to the more complex models that assume varying rates.
-We aimed to evaluate the understanding and recency of knowledge across various language models, including GPT-4 and others.
-To accomplish this, we devised a methodology to calculate a weighted accuracy score for each model, reflecting not only its performance but also the contemporaneity of its knowledge base.
-We began by assigning weights to different quarters, starting from 0 for the first quarter of 2015, up to 1 for the most recent quarter.
-This linear interpolation of weights was chosen to emphasize the value of more current information, underlining our hypothesis that accuracy concerning more recent knowledge is indicative of a model’s modernity and relevance.
-The formula applied was straightforward: for each model, we multiplied its accuracy (Acc) by the corresponding weight based on the time difference since the first quarter of 2015.
-The resulting weighted accuracies were then summed up to yield an overall score for each model, which we used to rank them accordingly.
-GPT-4 and other leading models demonstrated a notably later peak in knowledge, suggesting that their databases include more recent information.
-This is particularly valuable in fast-evolving fields where the timeliness of information is crucial.
-Below is a partial presentation of rankings and scores for selected models, showcasing their weighted accuracy scores. The complete table can be found in Table 15.
-Table 15: Model Scores
-Model Name Score
-GPT4-1106 794.275
-GPT4-0613 699.8
-Zephyr-7B-beta 601.35
-InternLM-Chat-7B 592.15
-GPT3.5-turbo-0613 578.025
-Baichuan-13B-Chat 571.8
-Qwen-14B-Chat 565.45
-OPT-2.7B 547.375
-Yi-6B-Chat 541.05
-Mistral-7B-v0.1 528.725
-Baichuan-7B-Chat 527.825
-LLaMA-7B 524.075
-Baichuan2-7B-Chat 517.725
-Pythia-12B 513.375
-Colossal-LLaMA2-7B-Base 513.35
-LLaMA2-7B-Chat 508.4
-LLaMA2-13B 508.15
-Qwen-7B 506.7
-Qwen-1.8B-Chat 505.225
-Phi-2 497.4
-Yi-6B 496.15
-Zhongjing-Base 490.1
-LLaMA2-7B 486.2
-Falcon-rw-1B 481.15
-ChatGLM3-6B 479.8
-Baichuan2-7B-Base 479.175
-Baichuan2-13B-Chat 463.75
-OPT-13B 453.075
-Baichuan2-13B-Base 446.225
-Skywork-13B-Base 445.8
-Qwen-1 8B 435.775
-RWKV5-Eagle-7B 408.425
-Qwen-7B-Chat 394.9
-Phi-1.5 385.975
-TinyLLaMA-1.1B-Chat-v0.6 366.125
-Appendix K Details of Model Accuracy Peaks
+Appendix M Details of Model Accuracy Peaks
 Tab. 14 captures the peak performance periods of various models on Freshbench. It lists specific models along with their respective dates when they achieved the best performance.
@@
 2024-07
-Appendix L Question Tag Details
-Table 17: Detailed breakdown and sum of entries per category within each group of benchmarks.
-Group Category Number of Entries
-Political and Government Affairs Leader Entry/Exit 199
-US Politics 338
-Elections and Referenda 265
-US Policy 223
-Foreign Policy 242
-Non-US Politics 415
-Security and Conflict 413
-Total: 2095
-Economic and Business Items Business 575
-Economic Indicators 253
-Finance 284
-Economic Policy 147
-Total: 1259
-Societal issues Items Society 379
-Health 237
-Environment 75
-Sports 120
-Entertainment 132
-Total: 943
-Science and Technology Items Technology 346
-Total: 346
-Appendix M Analysis of Best-Achieved Accuracy and Cutoff Time
+Appendix N Analysis of Best-Achieved Accuracy and Cutoff Time
 Table 15: Analysis about the time of the best-achieved accuracy and its cutoff time.
@@
 The peak accuracy periods analyzed are based on the average accuracy of all questions within three-month intervals, to provide a stable and robust assessment of each model’s performance.
-Appendix N Prediction Correlation over Time
-The graphical representations in our analysis plot the changes in correlation coefficients over time.
-Each line in the graph represents the changing relationship between a category’s average adjusted performance and a specific benchmark, highlighting periods of stronger or weaker association.
-Figure 7: Science and Technology Items - All Model Spearman Smoothed
-Figure 8: Science and Technology Items - Base Model Spearman Smoothed
-Figure 9: Societal Issues Items - All Model Spearman Smoothed
-Figure 10: Political and Government Affairs - All Model Spearman Smoothed
-Figure 11: Economic and Business Items - All Model Spearman Smoothed